### v1

# Harmonic covers of skeleta

Abstract.
The geometry of a toroidal scheme over a DVR is encoded in a $\mathbb{Z}$-PL space known as the dual polyhedral complex. Any such dual complex is a skeleton, i.e. a nonarchimedean analytic retract, and admits a combinatorial divisor theory via specialisation. These structures on the dual complex interact via a Poincaré-Lelong slope formula, which interprets specialisations of divisors as Laplacians of PL functions. The main result presented here shows that finite covers of toroidal schemes give harmonic morphisms of dual complexes, i.e. morphisms that preserve the tropical Laplace equation. A crucial ingredient is a balancing condition which is a variant of the tropical multiplicity formula for dual complexes.
We apply these results to obtain a Riemann-Hurwitz formula for covers of skeleta in any dimension: the Laplacian of the different function is the tropical relative canonical divisor.
Key words and phrases:
Tropical complexes, Berkovich skeleta, harmonic morphisms, toroidal schemes, Poincaré-Lelong formula, logarithmic differentials, Riemann-Hurwitz
2010 Mathematics Subject Classification:
Primary 14G22, 31C20; secondary 14D10, 32P05, 14T20, 31C05
The author was supported by the Fund for Scientific Research Flanders (PhD Fellowship 11F0123N and Research Project G0B1721N)
1. Introduction
1.1Motivation.
Harmonic functions play an important role in latest developments in nonarchimedean geometry. See for instance the recent work [GJR24], where it is shown (§7.6 in loc. cit.) that general morphisms of Berkovich analytic spaces are harmonic morphisms in the sense that harmonic functions pull back to harmonic functions. Since it is well-known that Berkovich spaces retract onto piecewise linear subspaces called skeleta, one is naturally led to contemplate harmonic morphisms of PL spaces in order to study Berkovich spaces from a polyhedral viewpoint, leading to a combinatorial or “tropical” approach to analytic geometry.
Skeleta are commonly constructed as dual complexes of toroidal schemes. We will focus on finite covers of such skeleta, working over discretely valued ground fields of arbitrary residue characteristic.
1.2Summary of the results..
We present four Theorems on skeleta and finite covers.
  • First is a balancing condition (Theorem 1.12) which gives a combinatorial notion of local degree.
  • Second is a variant of the Poincaré-Lelong formula for skeleta (Theorem 1.19).
  • Thirdly we show that that finite covers of skeleta are harmonic morphisms (Theorem 1.22).
  • Fourth is a Riemann-Hurwitz formula for skeleta (Theorem 1.26).
Let us now give a more detailed overview of this work.
1.3Toroidal schemes.
Throughout we work over a discretely valued ground field $k$ whose ring of integers will be denoted by $k^{\circ}$. For simplicity the residue field $\tilde{k}$ is assumed algebraically closed. Log-structures are defined with respect to the Zariski site, and we equip $S=\mathrm{Spec}\,k^{\circ}$ with the “standard” log-structure $S^{\dagger}$ defined via $k^{\circ}\setminus 0\to k^{\circ}$.
In this article, a toroidal scheme over $S$ is defined as an integral separated finite type $S$-scheme ${X}$ equipped with a Zariski log-structure
${X}^{\dagger}=({X},\mathscr{M}_{X})$
over $S^{\dagger}$ which is log-regular [Kat94], the last condition means ${X}^{\dagger}$ has some mild “toric” singularities that are adequately controlled by the log-structure – basic properties will be recalled in Section 2. Then the log-structure of ${X}^{\dagger}$ is divisorial with respect to the support
$D_{{X}}=\{x\in{X}:\mathscr{M}_{{X},x}\neq\mathscr{O}_{{X},x}^{\times}\},$
so that ${X}^{\dagger}$ is equivalently determined by the pair $({X},D_{{X}})$ for some “boundary” divisor $D_{{X}}$. Toroidal pairs generalise both:
  • snc pairs: ${X}$ is regular and $D_{{X}}$ is strict normal crossings (containing ${X}_{s}$)
  • toric pairs: ${X}$ is a toric $S$-scheme, $D_{{X}}$ is torus-invariant (containing ${X}_{s}$).
We emphasise that a toroidal $S$-scheme ${X}^{\dagger}$ is not necessarily (strictly) semistable, meaning that the special fiber ${X}_{s}$ is possibly not a reduced scheme. It is possible to study $S$-schemes which are toroidal with respect to the étale topology, but we do not pursue this here.
1.4Dual complexes.
A toroidal $S$-scheme ${X}^{\dagger}$ is stratified by the intersections of components of $D_{{X}}$, also called the log-stratification. The dual complex $\mathcal{D}({X}^{\dagger})$ is obtained as the dual intersection complex of the log-stratification and is a polyhedral complex which may possibly have some unbounded polyhedra corresponding to horizontal log-strata. An important fact is that the dual complex $\mathcal{D}({X}^{\dagger})$ admits a natural topological embedding
$\mathcal{D}({X}^{\dagger})\to{X}^{\mathrm{an}}$
whose image is called the skeleton $\mathrm{sk}({X}^{\dagger})$ of ${X}^{\dagger}$, see [BM19] and precursors [Ber99, KS06, MN15, GRW16]. The closure $\overline{\mathrm{sk}}({X}^{\dagger})$ of $\mathrm{sk}({X}^{\dagger})$ is a retract of ${X}^{\mathrm{an}}$ (5.31).
1.5Toroidal covers.
Throughout the paper, we denote by
$f:X^{\prime\dagger}\longrightarrow{X}^{\dagger}$
a finite morphism of proper toroidal $S^{\dagger}$-schemes or short a toroidal cover, and the associated cover of dual complexes is denoted by
$\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})\quad\text{ orsometimes }\quad\phi:\Sigma^{\prime}\to\Sigma.$
We are primarily interested in such covers because they provide an adequate “monomialised” setting to study analytifications of covers of $k$-varieties via degeneration techniques. There is an advantage of working with toroidal schemes as opposed to snc pairs: finite covers of snc pairs hardly arise in nature – the so-called “impossibility of simultaneous resolution” according to Abhyankar [Kat94, §12]. In contrast, if in the above only $f_{\eta}$ and ${X}^{\dagger}$ are given then the normalisation $f:X^{\prime\dagger}\to{X}^{\dagger}$ is a toroidal cover assuming $f$ is tamely ramified along $D_{X}$, and Kato raised the question whether in wildly ramified situations toroidal covers also exist (see §3.5).
$\mathbb{Z}$-PL complexes.
To simplify terminology we define a $\mathbb{Z}$-PL complex as a strictly convex rational polyhedral complex equipped with a $(\mathbb{Z},|k^{\times}|)$-piecewise linear structure (also known as an integral $k$-affine structure). In particular each polyhedron $\sigma$ of a $\mathbb{Z}$-PL complex $\Sigma$ is equipped, in a compatible fashion, with a lattice $M_{\sigma}$ of $\mathbb{Z}$-affine functions. For toroidal $S$-schemes ${X}^{\dagger}$ the dual complex $\mathcal{D}({X}^{\dagger})$ comes canonically equipped with a $\mathbb{Z}$-PL structure where the lattices $M_{\sigma}$ consist of toric coordinates on $\sigma$. More precisely every polyhedron $\sigma$ of $\mathcal{D}({X}^{\dagger})$ corresponds to a unique generic point $x$ of a log-stratum of ${X}^{\dagger}$ and the lattice of $\mathbb{Z}$-affine functions $M_{\sigma}$ is then given by
$M_{\sigma}=\mathscr{C}^{\mathrm{gp}}_{{X},x},$
this is the groupification of the characteristic monoid $\mathscr{C}_{x}=\mathscr{M}_{{X},x}^{\sharp}$ at $x$; details in Section 3. Geometrically $M_{\sigma}$ is the lattice of Cartier divisors on $\mathrm{Spec}\,\mathscr{O}_{{X},x}$ that are supported on $D_{{X},x}$, the inverse image of $D_{{X}}$ in $\mathscr{O}_{{X},x}$. This leads to an isomorphism (Lemma 5.4) between the group of Cartier divisors supported on $D_{{X}}$ and the group of $\mathbb{Z}$-affine functions on $\Sigma$, defined as
(1.7) $M_{\Sigma}\coloneqq\lim_{\sigma\in\Sigma}M_{\sigma}$
1.8Local degrees and the lattice index.
Let $\Sigma$ be a polyhedral complex of dimension $n$. By a facet-ridge pair $(\sigma,\tau)$ of $\Sigma$ we mean a pair of a facet $\sigma$ (a $n$-dimensional polyhedron) of $\Sigma$ and a ridge $\tau$ (a $(n-1)$-dimensional polyhedron) of $\Sigma$ such that $\tau$ is a face of $\sigma$.
Now let
$\phi:\Sigma^{\prime}\to\Sigma$
be a finite cover of $n$-dimensional $\mathbb{Z}$-PL complexes. Consider maps of a facet-ridge pair $(\tau^{\prime},\sigma^{\prime})$ of $\Sigma^{\prime}$ to a given facet-ridge pair $(\tau,\sigma)$ of $\Sigma$ – see Figure 1.11 for a picture. Then, fixing $\tau^{\prime}$, we call $\phi:\Sigma^{\prime}\to\Sigma$ balanced at $\tau^{\prime}$ if the following sum of lattice indices, called the local degree at $\tau^{\prime}$, is is independent of the choice of $\sigma$ and thus well-defined :
(1.9) $\deg_{\tau^{\prime}}\phi\coloneqq\sum_{\begin{subarray}{c}\text{facets }\,\sigma^{\prime}\\ (\sigma^{\prime},\tau^{\prime})\mapsto(\sigma,\tau)\end{subarray}}[M_{\sigma^{\prime}}:M_{\sigma}].$
The sum is over all facets $\sigma^{\prime}$ such that $(\sigma^{\prime},\tau^{\prime})$ is a facet-ridge pair above a fixed pair $(\sigma,\tau)$. We call $\phi$ a balanced cover if it is balanced at every ridge $\tau^{\prime}$, and then assuming $\Sigma$ is connected we can define the degree of $\phi$ as
(1.10) $\deg\phi\coloneqq\sum_{\tau^{\prime}\mapsto\tau}\deg_{\tau^{\prime}}\phi,$
independently of the choice of ridge $\tau$ of $\Sigma$.
Figure 1.11. A finite cover of $\mathbb{Z}$-PL complexes is called balanced if the local degree $\deg_{\tau^{\prime}}\varphi\coloneqq\sum_{(\sigma^{\prime},\tau^{\prime})\mapsto(\sigma,\tau)}[M_{\sigma^{\prime}}:M_{\sigma}]$ is independent of $\sigma$ (Definition 3.20).
Theorem 1.12 (Balancing formula, See Theorem 3.24).
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a finite morphism of toroidal $S^{\dagger}$-schemes. Then
$\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})$
is a balanced cover of $\mathbb{Z}$-PL complexes and
$\deg\phi=\deg f.$
1.13Metrisation of $\mathbb{Z}$-PL complexes.
Any $\mathbb{Z}$-PL complex $\Sigma$ can be equipped with a canonical measure $\mu_{\Sigma}$ that we call the conformal measure, such that isometries are isomorphisms (i.e. unimodular bijections). In essence, it is the unique piecewise Lebesgue measure which satisfies the following conformality property (Lemma 4.7): for a finite cover of bounded $\mathbb{Z}$-PL polyhedra $\sigma^{\prime}\to\sigma$, we have
(1.14) $\mathrm{vol}(\sigma)/\mathrm{vol}(\sigma^{\prime})=[M_{\sigma^{\prime}}:M_{\sigma}].$
1.15Specialisation of Cartier divisors.
Given a $\mathbb{Q}$-Cartier divisor $\mathscr{D}$ on ${X}$ we have intersection numbers $\mathscr{D}\cdot C\in\mathbb{Q}$ for any proper regular curve $C\subset{X}$. We define the specialisation $\mathrm{sp}_{*}\mathscr{D}$ as the combinatorial divisor
(1.16) $\mathrm{sp}_{*}\mathscr{D}\coloneqq\sum_{\tau}\frac{\mathscr{D}\cdot E_{\tau}}{\mathrm{vol}(\tau)}[\tau],$
summing over all vertical ridges $\tau$ of $\Sigma$; here $E_{\tau}$ is the $1$-dimensional log-stratum associated to $\tau$. In fact $\mathrm{sp}_{*}\mathscr{D}$ only depends on the linear equivalence class of $\mathscr{D}$.
1.17.
From the remainder of the introduction we assume that the dual complex $\mathcal{D}({X}^{\dagger})$ is simplicial, equivalently Weil divisors supported on $D_{X}$ are $\mathbb{Q}$-Cartier (5.8).
1.18Laplacians of PL functions.
For the definition of slopes of PL functions, we need to take into account additional structure next to the $\mathbb{Z}$-PL structure of the dual complex $\Sigma=\mathcal{D}({X}^{\dagger})$. In the spirit of the works [GRW16, Car20] we additionally decorate $\Sigma$ with some intersection numbers called $\alpha$-constants – this gives rise to the notion of a $\mathbb{Z}$-PL tropical complex (Definition 6.12). Informally speaking the slope $\partial_{\sigma/\tau}F$ of a $\mathbb{Z}$-affine PL function $F:\Sigma\to\mathbb{R}$ on a $\mathbb{Z}$-PL tropical complex $(\Sigma,\alpha)$ along a facet-ridge pair $(\sigma,\tau)$ should be thought of as an appropriate directional derivative of $F$ in a direction normal to $\tau$ that is distinguished by the $\alpha$-constants. The precise definition is given in 6.15.
A combinatorial divisor on $\Sigma$ is a formal sum of ridges with rational coefficients, and the Laplacian of $F$ is defined as the combinatorial divisor
$\Delta(F)=\sum_{(\sigma,\tau)}\partial_{\sigma/\tau}F[\tau],$
summing over facet-ridge pairs $(\sigma,\tau)$ with $\tau$ bounded; and we call $F$ harmonic if $\Delta(F)=0$.
Theorem 1.19 (Poincaré-Lelong slope formula, see Theorem 6.20).
Let $\mathscr{D}$ be a $\mathbb{Q}$-Cartier divisor on ${X}$ supported on $D_{{X}}$, and let $F_{\mathscr{D}}\in M_{\mathcal{D}({X}^{\dagger})}$ be the associated $\mathbb{Q}$-PL function, as in 1.6. Then
(1.20) $\Delta(F_{\mathscr{D}})={\mathrm{sp}}_{*}\mathscr{D}.$
Similarly for $\mathbb{Q}$-Cartier divisors.
1.21.
By the slope formula (1.20) it follows that $F_{\mathscr{D}}$ is harmonic if and only if $\mathscr{D}$ is numerically trivial on the vertical $1$-dimensional log-strata. This is for instance the case if $\mathscr{D}$ is a principal divisor, and then the function $F_{\mathscr{D}}$ is usually called smooth.
Given a balanced cover $\phi:\Sigma^{\prime}\to\Sigma$ of $\mathbb{Z}$-PL tropical complexes (Definition 6.12) we define the pullback of a combinatorial divisor via the rule $\phi^{*}[\tau]=\sum_{\tau^{\prime}\mapsto\tau}\deg_{\tau^{\prime}}\phi[\tau^{\prime}]$ and extending linearly.
Theorem 1.22 (Harmonic morphisms, see Theorem 6.22).
Let $\phi:\Sigma^{\prime}\to\Sigma$ be a balanced cover of $\mathbb{Z}$-PL tropical complexes. Then for any $\mathbb{Q}$-PL function $F$ we have
(1.23) ${\Delta}(\phi^{*}F)=\phi^{*}{\Delta}(F).$
in particular if $F$ is harmonic then so is $\phi^{*}F$, that is $\phi$ is a harmonic morphism.
1.24An extension to compactified skeleta.
Let $\overline{\Sigma}=\overline{\mathrm{sk}}({X}^{\dagger})$ denote the closure of $\mathrm{sk}({X}^{\dagger})$ in ${X}^{\mathrm{an}}$ – this is an analog of the compactified skeleton introduced in [GRW16]. Then there exists a continuous retraction $\overline{\rho}:{X}^{\mathrm{an}}\to\overline{\Sigma}$. One can extend the definition of specialisation to combinatorial divisors on $\overline{\Sigma}$, as well as an extended laplacian, and Theorem 1.19 extends to an equality $\overline{\Delta}(F_{\mathscr{D}})=\overline{\mathrm{sp}}_{*}\mathscr{D}$, see 6.25.
1.25Reminders on Riemann-Hurwitz formula.
The classical Riemann-Hurwitz formula can be reformulated as follows. If $g:X\to Y$ is a finite seperable morphism of proper smooth $k$-curves, then the pullback map
$g^{*}\Omega_{Y/k}\to\Omega_{X/k}$
yields a global section $\tau_{X/Y}$, called the trace section, of the relative canonical bundle
$\omega_{X/Y}\coloneqq\det\Omega_{X/Y}\cong\underline{\mathrm{Hom}}(g^{*}\Omega_{Y/k},\Omega_{X/k}).$
The Riemann-Hurwitz formula is then obtained by computing the degree of the relative canonical divisor $K_{X/Y}\coloneqq\mathrm{div}\,\tau_{X/Y}$ in two ways:
$\chi(X)-(\deg g)\chi(Y)=\sum_{x\in X}\mathrm{length}_{\mathscr{O}_{X,x}}{\Omega_{X/Y,x}}.$
The summand of the right-hand side is also known as the different $\delta_{g}(x)\coloneqq\mathrm{length}_{\mathscr{O}_{X,x}}{\Omega_{X/Y,x}}$ of $g$ at $x$.
Theorem 1.26 (Riemann-Hurwitz formula for skeleta, see Theorem 8.7).
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a finite seperable cover of proper toroidal $S$-schemes. Let $\phi:\Sigma^{\prime}\to\Sigma$ be the assocated cover of skeleta. Then there exists a $\mathbb{Z}$-PL function $\delta:\Sigma^{\prime}\to\mathbb{R}_{\geq 0}$, called the different function, such that for each divisorial point $x^{\prime}\in\Sigma^{\prime}$ and $x=\phi(x^{\prime})$ we have
$\delta(x)=\frac{1}{e(\mathscr{H}(x^{\prime})/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}{\Omega^{\mathrm{log}}_{\mathscr{H}(x^{\prime})^{\circ}/\mathscr{H}(x)^{\circ}}}$
and moreover
(1.27) $\Delta(\delta)=K_{\Sigma^{\prime}/\Sigma}$
where $K_{\Sigma^{\prime}/\Sigma}=\sum_{\tau^{\prime}}-\chi_{\Sigma^{\prime}}(\tau^{\prime}/\tau)[\tau^{\prime}]$ is the tropical relative canonical divisor (Definition 8.3).
1.28Weight functions.
The method of the proof of Theorem 8.7 also gives the following results, further detailed in Section 8.
  • Suppose $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k})$ is a nonzero holomorphic canonical form and let $\mathrm{wt}_{\omega}:{X}^{\mathrm{mon}}\to\mathbb{R}_{\geq 0}$ denote the associated weight function of [MN15]. Then
    $\Delta(\mathrm{wt}_{\omega}|_{\Sigma})=K_{\Sigma}-\rho_{*}\mathrm{div}\,\omega;$
    this extends a Theorem of Baker-Nicaise [BN16] to arbitrary dimension, see 8.12.
  • There exists a function $\lambda:X^{\mathrm{an}}\to\mathbb{R}_{\geq 0}$ which is $\mathbb{Z}$-affine on the faces of dual complexes, such that if $x\in X^{\mathrm{an}}$ is divisorial we have $\lambda(x)=\frac{1}{e(\mathscr{H}(x)/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}\Omega^{\mathrm{log}}_{\mathscr{H}(x)^{\circ}/k^{\circ}}$; this gives some evidence for an expectation of Temkin on the piecewise linearity of Kähler norms (Remark 8.11).
1.29.
(Companion paper) The companion paper [Wae25] contains some complements and applications in the case of curves. There we study the metrisation of Berkovich curves over discretely valued fields in more detail and show that the different function controls simultaneous curve skeleta (shown earlier in [CTT16] with different hypotheses) and give applications to reduction types of arithmetic curves and wild surface quotient singularities.
1.30Related work.
The balancing formula of 1.12 is strongly analogous to tropical multiplicity formulas [ST08, BPR16, GRW16, GJR24], where similar sums of lattice indices play the role of the degree. Our result is different since we consider covers of a skeleton by a skeleton and not covers of a tropicalisation by a skeleton. Roughly speaking our case corresponds to the setting of tropical multiplicity one. A similar notion of a balanced cover of complexes in all dimensions appears in [Var24].
The notion of a harmonic morphism presented in 1.22 closely mimics the eponymous notion in Riemannian geometry [Fug78, Ish79, BW03, EF01], but is not directly related. In dimension one discrete analogues were well-studied, in particular see the works [Ura00, BN09, ABBR15].
Lebesgue measures on dual complexes were studied in [JN20], where the conformal behaviour with respect to tamely ramified base change was already observed. Our construction of the conformal measures was also influenced by the available theory of metrisation of Berkovich curves [BN16, BPR13, CTT16, Duc14].
The main inspiration for the slope formula 1.19 is the prior work [GRW16], which deals with the strictly semistable case over algebraically closed ground fields and builds on earlier work in [Car20, BPR16].
The different function for covers of analytic curves was studied systematically in a series of works [Tem16, Tem17, CTT16, BT20] over algebraically closed ground fields, and has its historical origin in Gabber’s work on the $p$-adic Riemann existence problem (80s unpublished), see [Lüt93]. The Riemann-Hurwitz formula 1.26 for possibly wildly ramified covers of skeleta was pioneered in [CTT16] over curves over algebraically ground fields, see the references there for earlier work on wild ramification of analytic curves. In upcoming work, Hübner-Temkin [HT25] study analytic properties of the different function over any nonarchimedean field, but not the Riemann-Hurwitz formula.
After a preliminary version of this paper was written, Yanbo Fang kindly pointed the author to the earlier work of A. Gross [Gro18], where tropical intersection theory is developed on the dual complex of a toroidal variety over a field. There is some overlap, but his results do not adress mixed characteristic situations.
1.31Overview of the paper.
Section 2 covers preliminaries on logarithmic regularity. For convenience we give a self-contained exposition of the tools from logarithmic geometry that we use. We have put all necessary background on polyhedral geometry and fans in a separate Appendix A. In Section 3 we review dual complexes of toroidal $S$-schemes, describe their $\mathbb{Z}$-PL structure and prove Theorem 1.12. In Section 4 we construct the conformal measures on $\mathbb{Z}$-PL complexes and give explicit descriptions for dual complexes. Section 5 studies specialisation of Cartier divisors and some intersection numbers are computed. Section 6 introduces slopes and the Laplacian of PL functions and contains the proofs of Theorems 1.19 and  1.22. Section 7 contains some preparations on logarithmic differentials, our comparison of canonical sheaves and log-canonical sheaves may be of independent interest. In Section 8 we discuss potential theory of differential forms and prove Theorem 1.26.
1.32Conventions and notation.
We admit a modest familiarity with logarithmic structures [Kat89, Kat94], geometry of toric fans [Oda88], and Berkovich analytifications [Ber90, BPR13].
Throughout we denote by $k$ a discretely valued field. Its ring of integers is denoted by $k^{\circ}$ and its residue field, assumed algebraically closed, is denoted by $\tilde{k}$. We let $p\geq 1$ denote the characteristic exponent of $\tilde{k}$. We write $\varpi$ for a choice of uniformiser of $k^{\circ}$. We write $S=\mathrm{Spec}\,k^{\circ}$. The generic and closed point of $S$ are denoted by $\eta$ and $s$ respectively, so $S=\{\eta,s\}$ as a set.
All rings and monoids are assumed to be commutative. All schemes are assumed to be locally Noetherian. Logarithmic structures are denoted by $(\cdot)^{\dagger}$ and defined with respect to the Zariski site. If $X$ is a geometric object, then $|X|$ denotes the underlying topological space. Underlined objects like $\underline{\mathrm{Hom}}$ denote sheaves.
For the reader’s convenience, we discuss conventions on monoids, cones, fans, polyhedra, and $\mathbb{Z}$-PL spaces at length in appendix A, in summary:
  • All cones and polyhedra are assumed convex rational polyhedral and $\mathbb{Z}$-PL (i.e. integral $k$-affine, see A.33), they are denoted by $\sigma=(|\sigma|,N_{\sigma})$ where $N_{\sigma}$ is the lattice of integral points of the real linear span $(N_{\sigma})_{\mathbb{R}}$ of $|\sigma|$. We write $M_{\sigma}=N_{\sigma}^{*}=\mathrm{Hom}(N_{\sigma},\mathbb{Z})$ for the dual lattice of $\mathbb{Z}$-affine functions on $\sigma$.
  • Given a polyhedral complex $\Sigma$ we write $\Sigma(n)$ for the set of $n$-dimensional polyhedrons of $\Sigma$.
  • The letter $M$ is used frequently, so we use the following convention: straight $M$ refers to dual lattice of $\mathbb{Z}$-affine functions and curly $\mathscr{M}$ refers to log-structure.
  • If $\mathscr{M}$ is a monoid, we write $\mathscr{M}^{\times}$ for its group of invertible elements, $\mathscr{M}^{\sharp}\coloneqq\mathscr{M}/\mathscr{M}^{\times}$ for its sharpification and $\mathscr{M}^{\mathrm{gp}}$ for its groupification. We denote by $\mathscr{M}^{\mathrm{int}}$ and $\mathscr{M}^{\mathrm{sat}}$ the integralisation and saturation respectively. Most monoids we encounter will be fs, i.e. fine and saturated.
Berkovich $k$-analytification $(\cdot)^{\mathrm{an}}$ will only be used in Section 8. Completed residue fields are denoted $\mathscr{H}(\cdot)$.
Shorthand Meaning
fs monoid fine saturated monoid
snc divisor strict normal crossings divisor
log-structure logarithmic structure with respect to Zariski site
$\mathbb{Z}$-PL $(\mathbb{Z},|k^{\times}|)$-piecewise linear
$S^{\dagger}$ $S=\mathrm{Spec}\,k^{\circ}$ equipped with standard log-structure $k^{\circ}\setminus\{0\}\to k^{\circ}$
${X}^{\dagger}$ a toroidal $S$-scheme
$D_{{X}}$ support of the log-structure of ${X}^{\dagger}$, viewed as a reduced divisor.
$F({X}^{\dagger})$ Kato fan associated to ${X}^{\dagger}$
$\mathscr{C}_{x}=\mathscr{C}_{{X},x}$ characteristic monoid at a point $x\in F({X}^{\dagger})$.
$\Sigma({X}^{\dagger})$ cone complex associated to $F({X}^{\dagger})$
$C(x)=\mathbb{R}_{\geq 0}\mathscr{C}_{x}^{\vee}$ cone of $\Sigma({X}^{\dagger})$ associated to $x\in F({X}^{\dagger})$, i.e. dual cone of $\mathscr{C}_{{X},x}$
$M_{x}=M_{C(x)}=\mathscr{C}_{{X},x}^{\mathrm{gp}}$ lattice of $\mathbb{Z}$-linear functions on $C(x)$,
$N_{x}=N_{C(x)}=(\mathscr{C}_{{X},x}^{\vee})^{\mathrm{gp}}$ lattice of integral points on $C(x)$
$\langle\cdot,\cdot\rangle:M_{x}\times N_{x}\to\mathbb{Z}$ canonical pairing
$\mathcal{D}({X}^{\dagger})$ dual polyhedral complex associated to ${X}^{\dagger}$, i.e. “$\left[\varpi\right]\mapsto 1$”-slice of $\Sigma({X}^{\dagger})$.
$\sigma_{x}$ face of $\mathcal{D}({X}^{\dagger})$ associated to $x\in F({X}^{\dagger})$, i.e. “$\left[\varpi\right]\mapsto 1$”-slice of $C(x)$.
$M_{\sigma_{x}}=M_{x}$ lattice of $\mathbb{Z}$-affine functions on $\sigma_{x}$
$E_{\sigma}$ closed stratum associated to a polyhedron $\sigma$ of $\mathcal{D}({X}^{\dagger})$.
$\mathrm{sk}({X}^{\dagger})$ skeleton of ${X}^{\dagger}$, i.e. homeomorphic image of $\mathcal{D}({X}^{\dagger})$ in ${X}_{\eta}^{\mathrm{an}}$.
$M_{\mathrm{sk}({X}^{\dagger})}$ group of $\mathbb{Z}$-affine functions on $\mathrm{sk}(\mathrm{sk}({X}^{\dagger})$
$\mu=\mu_{\Sigma}$ conformal measure on a $\mathbb{Z}$-PL complex $\Sigma$
$(\sigma,\tau)$ a facet-ridge pair of a PL-space
$f:X^{\prime\dagger}\to{X}^{\dagger}$ A toroidal cover, i.e. a finite morphism of toroidal $S$-schemes
$\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})$ cover of dual complexes associated to $f$
Table 1.32. List of abbreviations and notations used throughout the paper
1.33Acknowledgements.
The results of this article are part of the author’s PhD thesis, supervised by Johannes Nicaise; I am very grateful for his guidance and advice over the years. It is also my pleasure to thank Stefan Wewers, Lars Halvard Halle, Michaël Maex, Michael Tëmkin, Yanbo Fang, Antoine Ducros, Alejandro Vargas and Soham Karwa for valuable discussions related to this work.
2. Background on logarithmic regularity
2.1.
In this section we recall some well-known properties of log-regular log-schemes [Kat94]. Some properties of monoids, fans and polyhedral geometry are recalled in Appendix A. Experienced readers may wish to skip to Section 3 and refer back as needed. Comprehensive references on log-geometry include [Ogu18, Kat89, Kat94, Niz06, GR18, Tem24].
2.2Log-schemes.
Let $X^{\dagger}=(X,\mathscr{O}_{X},u:\mathscr{M}_{X}\to\mathscr{O}_{X})$ be a log-scheme, that is a scheme $(X,\mathscr{O}_{X})$ equipped with a log-structure $u:\mathscr{M}_{X}\to\mathscr{O}_{X}$. If $X$ is a scheme and $D\subset X$ is a closed subset we write $X(\mathrm{log}D)$ for the log-scheme $(X,\mathscr{O}_{X},\mathscr{O}_{X}\cap i_{*}\mathscr{O}^{\times}_{X\setminus D}\to\mathscr{O}_{X})$ where $i:X\setminus D\to X$ is the open immersion. If $u:\mathscr{M}_{X}\to\mathscr{O}_{X}$ is a pre-log-structure we denote by $u^{a}:\mathscr{M}_{X}^{a}\coloneqq\mathscr{M}_{X}\oplus_{u^{-1}\mathscr{O}_{X}^{\times}}\mathscr{O}_{X}^{\times}\to\mathscr{O}_{X}$ the associated log-structure. Recall that a log-scheme $X^{\dagger}$ is called coherent if $X$ is covered by charts $P_{U}\longrightarrow\Gamma(U,\mathscr{M}_{U})$ for some finitely generated monoids $P_{U}$. Recall that this means that if $\underline{P_{U}}$ denotes the constant sheaf on $U$, we have an isomorphism $\underline{P_{U}}^{a}\overset{\sim}{\longrightarrow}\mathscr{M}_{U}$, in other words there is a strict map $U^{\dagger}\to\mathbb{A}_{P}\coloneqq(\mathrm{Spec}\,(\mathbb{Z}[P]),\underline{P}^{a})$. We call $X^{\dagger}$ fine (resp. fs) if it is coherent and $\mathscr{M}_{X}$ has integral (resp. saturated) stalks. Convention: if we call $X^{\dagger}$ integral, then we mean the underlying scheme $X$ is integral. The forgetful functor from the category of fine (resp. fs) log-schemes to coherent log-schemes admits a right adjoint namely integralisation $(\cdot)^{\mathrm{int}}$ (resp. saturation $(\cdot)^{\mathrm{fs}}$). Beware that integralisation or saturation may change the underlying scheme. Base change in the category of fine (resp. fs) log-schemes is obtained by applying $(\cdot)^{\mathrm{int}}$ (resp. $(\cdot)^{\mathrm{fs}}$) to the naive base change. All fs log-schemes $X^{\dagger}=(X,\mathscr{M}_{X})$ admit a chart near $x$ by the characteristic monoid $\mathscr{C}_{X,x}\coloneqq\mathscr{M}_{X,x}^{\sharp}$, also known as a neat chart, see for instance [Ogu18, §II.2.3]. If $X^{\dagger}$ is coherent then by [GR18, §12.2.21] $\mathscr{C}_{X}$ is constructible and the rank function $x\mapsto\mathrm{rank}(x)\coloneqq\mathrm{rank}_{\mathbb{Z}}\mathscr{C}_{X,x}^{\mathrm{gp}}$ is upper-semicontinuous; so we obtain a rank stratification of $X=\sqcup_{r\geq 0}X_{r}$ by locally closed subsets of equal rank.
2.3log-parameters.
Let $X^{\dagger}$ be a log-scheme, and let $\mathscr{M}^{+}_{X}\coloneqq\mathscr{M}_{X}\setminus\mathscr{M}_{X}^{\times}$ be the maximal monoid ideal of $\mathscr{M}_{X}$. Let $\mathcal{I}_{X}$ be the (ring) ideal sheaf of $\mathscr{O}_{X}$ generated by $\mathscr{M}^{+}_{X}$ – we also call $\mathcal{I}_{X}$ the sheaf of logarithmic parameters. We write
$D_{X}=\mathrm{supp}\,\mathcal{I}_{X}$
for the support of $\mathcal{I}_{X}$. Note $D_{X}$ consists of the points of $X$ where the log-structure is non-trivial, so we will also call $D_{X}$ the support of the log-structure.
2.4Log-regularity.
A log-scheme $X^{\dagger}$ is defined to be log-regular at a point $x\in X$ if $X^{\dagger}$ is fs at $x$ and $V(\mathcal{I}_{X,x})$ defines a regular closed subscheme of $\mathrm{Spec}\,\mathscr{O}_{X,x}$ of codimension equal to $\mathrm{rank}(x)$.
Any log-regular schemes is normal [Kat94, §4.1] and a fs log-scheme log-smooth111 Recall that a morphism $f^{\dagger}:X^{\dagger}\longrightarrow Y^{\dagger}$ of coherent log-schemes is called log-smooth if it is locally of finite type and formally log-smooth (defined via infinitesimal liftings of log-thickenings in [Ogu18, §IV.3.1]). Log-smooth morphisms are preserved by composition and fs base change [Ogu18, IV.3.2]. In practice one uses Kato’s criterion [Kat89, §3.5]or [Ogu18, §3.1.13]: if $\phi:P\to Q$ is a coherent chart at $x\in X$ such that #$\#\ker\phi^{\mathrm{gp}}$ and #$\#\left(\mathrm{coker}\phi^{\mathrm{gp}}\right)_{\mathrm{tor}}$ are invertible in $k(x)$, and $X\to Y\times_{\mathrm{Spec}\,\mathbb{Z}[P]}\mathrm{Spec}\,\mathbb{Z}[Q]$ is smooth (in the usual sense), then $f$ is log-smooth at $x$; and analogously for log-étale morphisms. over a log-regular scheme is log-regular [Kat94, §8.2].
2.5Properties of log-regular log-schemes.
Suppose $X^{\dagger}$ is a log-regular log-scheme. It follows from [Kat94, §11.6] that the support $D_{X}$ is of pure codimension one and $X^{\dagger}=X(\mathrm{log}D_{X})$. Also $V(\mathcal{I}_{X,x})\subset\mathrm{Spec}\,\mathscr{O}_{X,x}$ is the intersection of the $\mathrm{rank}(x)$ components of $D_{X,x}$ that contain $x$, where we write $D_{X,x}$ for the inverse image of $D_{X}$ along $\mathrm{Spec}\,\mathscr{O}_{X,x}\to X$. Observe that $\mathscr{C}_{X,x}$ is isomorphic to the monoid of effective Cartier divisors on $\mathrm{Spec}\,\mathscr{O}_{X,x}$ supported on $D_{X,x}$.
Example 2.6 (snc divisors).
For example suppose that $X$ is a regular integral scheme of dimension $d$ and $D_{X}$ defines a (reduced) snc divisor on $X$ [Sta, 0CBN]; recall this means that at every point $x\in X$ there exists a system of regular parameters $t_{1},\dots,t_{d}\in\mathscr{O}_{X,x}$ such that $D_{X,x}$ is defined by $\prod_{i=1}^{r}t_{i}=0$ in $\mathscr{O}_{X,x}$ for some $r\leq d$. Then $X^{\dagger}=X(\mathrm{log}D_{X})$ is log-regular: indeed, near $x$ we have a sharp fs chart $\mathbb{N}^{r}\to\mathscr{O}_{X}:(n_{1},\dots,n_{r})\mapsto t_{1}^{n_{1}}\cdots t_{r}^{n_{r}}$ so $\mathscr{C}_{X,x}\cong\mathbb{N}^{r}$ and $\mathrm{rank}(x)=r$, and $\mathcal{I}_{X,x}$ is generated by $t_{1},\dots,t_{r}$ with $\mathrm{codim}V(\mathcal{I}_{X,x})=r$.
In fact we have a converse: if $X^{\dagger}$ is a log-regular log-scheme then by 2.5 we have $X^{\dagger}=X(\mathrm{log}D_{X}))$ and if the underlying scheme $X$ is regular then $D_{X}$ is a snc divisor: indeed, at each $x\in X$, the quotient of the regular ring $\mathscr{O}_{X,x}$ by $\mathcal{I}_{X,x}$ is regular, which implies that $\mathcal{I}_{X,x}$ is generated by regular parameters by [Sta, 00NR].
2.7Fan of a log-regular log-scheme.
Let $X^{\dagger}$ be any log-regular log-scheme. Then we construct the fan $F(X^{\dagger})$ associated to $X^{\dagger}$ as follows. As a topological space, $F(X^{\dagger})$ is defined as the subspace $\{x\in X:\,\mathcal{I}_{X,x}=\mathfrak{m}_{X,x}\}$, with $\mathcal{I}_{X}$ as defined in 2.3. We denote by $\mathscr{C}_{F(X^{\dagger})}$ the restriction of $\mathscr{C}_{X}$ along the canonical inclusion $F(X^{\dagger})\to X$. Then $(F(X^{\dagger}),\mathscr{C}_{F(X^{\dagger})})$ is a fan in the sense of Kato (see Definition A.21) – indeed: suppose $U\to X$ is an affine open neighbourhood of $x$ such that $P\to\Gamma(U,\mathscr{M}_{U})$ is an fs chart of $X^{\dagger}$ near $x$, then [GR18, 12.6.9] shows that after possibly further shrinking $U$ we have
$F(U^{\dagger})\cong(\mathrm{Spec}\,P)^{\sharp}.$
There is a canonical strict morphism of sharp monoidal spaces
$\pi_{X^{\dagger}}:(X,\mathscr{C}_{X})\to F(X^{\dagger}),$
called the characteristic morphism of $X^{\dagger}$, which sends a point $x$ to the prime ideal defined by $I_{X^{\dagger},x}$ in $\mathrm{Spec}\,\mathscr{O}_{X,x}$, it is open and in fact initial among all open maps of $(X,\mathscr{M}_{X}^{\sharp})$ to a sharp monoidal space (see op.cit. or [Kat94, 10.2]), this gives a categorical222informally, one can think of $\pi$ as a categorical quotient, though unfortunately this does not make sense since fans and log-schemes live in different categories. It is plausible one can remedy this by using Artin stacks instead of Kato fans – see [ACM+15] and the work of Ulirsch [Uli18]. definition of $F(X^{\dagger})$. For every morphism $f:X^{\dagger}\to Y^{\dagger}$ of log-regular log-schemes we have an induced map $F(f):F(X^{\dagger})\to F(Y^{\dagger})$ of fans, obtained via the composition $F(X^{\dagger})\to(X,\mathscr{C}_{X})\to(Y,\mathscr{C}_{Y})\to F(Y^{\dagger})$.
2.8log-stratification.
Let $X^{\dagger}$ be log-regular log-scheme. Each point of $F(X^{\dagger})$ is locally closed, therefore since $\pi=\pi_{X^{\dagger}}:(X,\mathscr{M}_{X})\to F(X^{\dagger})$ is open, the fibers of $\pi$ define a stratification of $X$ into irreducible locally closed subsets, here called the log-stratification, which by [BM19, Lemma 2.2.3] coincides with the stratification defined by the support $D_{X^{\dagger}}$. Hence the points of $F(X^{\dagger})$ parametrise log-strata. For a point $x\in F(X^{\dagger})$ we denote by $\pi^{-1}{x}=E^{\circ}_{x}$ the associated log-stratum and $E_{x}=\overline{\{x\}}$ for its schematic closure. We call $E_{x}^{\circ}$ (resp. $E_{x}$) the log-stratum (resp. closed log-stratum) associated to $x$. Note that the log-stratification refines the rank stratification of 2.2.
Lemma 2.9.
Let $X^{\dagger}$ be a log-regular log-scheme and let $x\in F(X^{\dagger})$. Let $E_{x}=\overline{\{x\}}$ be the associated log-stratum. Equip $E_{x}$ with the log-structure $E_{x}^{\dagger}=E_{x}(\mathrm{log}D_{x})$ where $D_{x}=E_{x}\setminus E_{x}^{\circ}$. Then we have the following.
  1. (i)
    The points of $\mathrm{Spec}\,\mathscr{C}_{X,x}$ parametrise log-strata containing $E_{x}$.
  2. (ii)
    The log-scheme $E_{x}^{\dagger}$ is log-regular
  3. (iii)
    The natural map $\iota:F(E_{x}^{\dagger})\to F(X^{\dagger})$ is a homeomorphism onto $\overline{\{x\}}$.
  4. (iv)
    Let $y\in\overline{\{x\}}$. Then the cospecialisation map $\tau_{y,x}:\mathscr{C}_{X,y}\to\mathscr{C}_{X,x}$ is surjective
  5. (v)
    For all $y\in\overline{\{x\}}$,
    (2.10) $\ker\tau_{y,x}\cong\mathscr{C}_{E_{x}^{\dagger},\iota^{-1}(y)}.$
Proof.
Part (i) follows from the existence of neat charts (2.2) and the descriptions of $F(X^{\dagger})$ in 2.7 and 2.8.
Part (ii): For simplicity write $E=E_{x}$ and $D=D_{X}$. Let $D_{1},\dots,D_{m}$ be the irreducible components of $D$. Suppose $E=D_{1}\cap\dots\cap D_{r}$ where $r=\mathrm{rank}_{X^{\dagger}}(x)$. Then $E^{\dagger}=E(\mathrm{log}(D_{x}))$ and $D_{x}=E\cap(D_{r+1}\cup\dots\cup D_{m})$. The strata of $D_{x}$ are the strata of $D$ contained in $E$, in particular these are regular. For any $y\in E$, the rank $\mathrm{rank}_{E^{\dagger}}(y)$ is bounded by the number of components $D_{i}$ containing $E_{y}$, where $r<i\leq m$; this implies
(2.11) $\mathrm{rank}_{E^{\dagger}}(y)\leq\mathrm{rank}_{X^{\dagger}}(y)-r.$
Since $X^{\dagger}$ is log-regular, we have
$\mathrm{rank}_{X^{\dagger}}(y)=\mathrm{codim}_{X}(E_{y})=\mathrm{codim}_{E}(E_{y})+\mathrm{codim}_{X}(E)=\mathrm{codim}_{E}(E_{y})+r.$
Together, this gives $\mathrm{rank}_{E^{\dagger}}(y)\leq\mathrm{codim}_{E}(E_{y})$. By [Kat94, §2.3] it follows that equality holds, therefore $E^{\dagger}$ is log-regular.
Part (iii): The natural map $\iota:F(E^{\dagger})\to F(X^{\dagger})$ corresponds to the inclusion of the generic points of log-strata of $X^{\dagger}$ contained in $E$. Since $\iota$ preserves specialisations it is a homeomorphism onto its image.
Part (iv): Recall the notion of prime ideals and faces of a monoid from A.13. The cospecialisation $\mathscr{M}_{X,y}\to\mathscr{M}_{X,x}$ is obtained by localising $\mathscr{M}_{X,y}$ at the prime ideal corresponding to $x$, which is $\mathfrak{p}=\{s\in\mathscr{M}_{X,y}:s(x)=0\}$. So
$\mathscr{C}_{X,x}\cong\left(\left(\mathscr{M}_{X,y}\right)_{\mathfrak{p}}\right)^{\#}\cong\mathscr{M}_{X,y}/(\mathscr{M}_{X,y}\setminus\mathfrak{p})$
and this implies $\tau_{y,x}$ is surjective. In fact, $\tau_{y,x}$ is the quotient of $\mathscr{C}_{X,y}$ by the submonoid $\mathscr{C}_{X,y}\setminus\mathfrak{p}^{\#}$, i.e. the face of $\mathscr{C}_{X,y}$ associated to $x$.
Part (v): the monoid $\mathscr{C}_{X,x}$ (resp. $\mathscr{C}_{X,y}$) is naturally isomorphic with the monoid of effective Cartier divisors on $\mathrm{Spec}\,\mathscr{O}_{X,x}$ (resp. $\mathrm{Spec}\,\mathscr{O}_{X,y}$) supported on the inverse image of $D$. It follows that $\ker\tau_{y,x}$ is isomorphic to the monoid of effective Cartier divisors $\mathrm{Spec}\,\mathscr{O}_{X,y}$ supported on the inverse image of $D_{r+1}\cup\dots\cup D_{m}$, borrowing notation from (ii). On the other hand $\mathscr{C}_{E,\iota^{-1}(y)}$ is the monoid of effective Cartiers divisors on $\mathrm{Spec}\,\mathscr{O}_{E,y}$ supported on the preimage of $D_{y}$. Restriction of local equations yields a natural surjective monoid morphism $\psi:\ker\tau_{y,x}\to\mathscr{C}_{E,\iota^{-1}(y)}$. A rank count, using (iii) and (2.11) then shows that $\psi^{\mathrm{gp}}$ is injective, and hence $\psi$ is injective too.
Remark 2.12.
To bring out the analogy with snc divisors even more, one can use ideal quotients of monoids (see A.14) and rephrase (2.10) as $\mathscr{C}_{X,y}/\mathfrak{p}^{\sharp}\cong\mathscr{C}_{E,\iota^{-1}(y)}^{0}$. We found this viewpoint helpful for describing compactifications of skeleta (see 5.31), but it will not be used in an essential way.
Proposition 2.13 (Toroidal resolutions).
Let $X^{\dagger}$ be a log-regular log-scheme with fan $F=F(X^{\dagger})$, and let $F^{\prime}\to F$ be a proper subdivision of $F$ (Definition A.27) with $F^{\prime}$ locally fs. Then there exists a final object $X^{\prime\dagger}$ among all log-schemes above $X^{\dagger}$ that admit a morphism of monoidal spaces $(X^{\prime\dagger})^{\sharp}\to F^{\prime}$ such that the square of characteristic morphisms
${(X^{\prime},\mathscr{M}_{X^{\prime}}^{\sharp})}$${F^{\prime}}$${(X,\mathscr{M}_{X}^{\sharp})}$${F(X^{\dagger}).}$
commutes, we suggestively [^3] Beware that strictly speaking, this notation is incorrect: such a fibered product does not exist because fans and log-schemes live in other categories. write
$X^{\prime\dagger}=X^{\dagger}\times_{F}F^{\prime}.$
Moreover $X^{\prime\dagger}$ is log-regular and $(X^{\prime\dagger})^{\sharp}\to F^{\prime}$ is the characteristic morphism.
If additionally $F^{\prime}\to F$ is a regular subdivision as in A.28 then $X^{\prime}\to X$ is a resolution of singularities (i.e. a proper birational morphism such that $h^{-1}(X_{\mathrm{reg}})\overset{\sim}{\longrightarrow}X_{\mathrm{reg}}$); In that case we will also call $X^{\prime\dagger}\to X^{\dagger}$ a toroidal resolution.
Proof.
Omitted, this is the heart result of [Kat94, §9-10], also see [GR18, §6.6] for a more detailed proof.
For later reference we recall how to construct $(X^{\prime})^{\dagger}$: suppose that the proper subdivision $F^{\prime}\to F$ is affine-locally given by $\mathrm{Spec}\,\phi:\mathrm{Spec}\,P^{\prime}\to\mathrm{Spec}\,P$ for some morphism $\phi:P\to P^{\prime}$ of sharp fs monoids. Then $X^{\prime}$ is locally constructed as
$X\times_{\mathrm{Spec}\,\mathbb{Z}[P]}\mathrm{Spec}\,\mathbb{Z}[Q]$
where
$Q=\{a\in P^{\mathrm{gp}}:\phi^{\mathrm{gp}}(a)\in P^{\prime}\},$
and $Q$ provides a chart for the log-structure on $X^{\prime}$. An application of Kato’s criterion shows that $(X^{\prime})^{\dagger}\to X$ is log-smooth, so it follows that $(X^{\prime})^{\dagger}$ is log-regular. The morphism $X^{\prime}\to X$ is proper since $F^{\prime}\to F$ is a proper subdivision. If $F^{\prime}$ is regular then $X^{\prime}$ is regular [GR18, 12.5.35]. For more details, consult the works cited above. ∎
Remark 2.14.
The proof of ‘̀log-crepant”, see LABEL:log-crepant. in the in the sense that This will play a role in
Also see the work of Niziol [Niz06, §5.8] where log-blowups are discussed and it is shown that one can always toroidally resolve Zariski log-regular schemes via a log-blowup – we won’t use this.
2.15The dual cone complex $\Sigma({X}^{\dagger})$..
Recall that locally fs (Kato) fans correspond to cone complexes, see Section A.26. Given a log-regular log-scheme $X^{\dagger}$ we write $\Sigma(X^{\dagger})=\Sigma(F(X^{\dagger})$ for the cone complex associated to $F(X^{\dagger})$ – its construction is recalled in A.25. In summary, it is the geometrical realisation of $F({X}^{\dagger})$ defined by glueing, for all $x\in F({X}^{\dagger})$, the dual cones
$\sigma_{\mathscr{C}_{F({X}^{\dagger}),x}}\coloneqq\mathrm{Hom}(\mathscr{C}_{F({X}^{\dagger}),x},\mathbb{R}_{\geq 0})$
along the cospecialisation maps as face embeddings, see (A.24). For shortness, we will also write $C(x)\coloneqq\sigma_{\mathscr{C}_{F({X}^{\dagger}),x}}\subset\Sigma({X}^{\dagger})$ for the cone associated to $x\in F({X}^{\dagger})$. Note that
$\Sigma({X}^{\dagger})=\mathrm{colim}_{x\in F({X}^{\dagger})}C(x)$
and for any two $x_{1},x_{2}\in F({X}^{\dagger})$ either $C(x_{1})\cap C(x_{2})=\emptyset$ or we have $C(x_{1})\cap C(x_{2})=C(x)$ where $x$ is the generic point of $\overline{\{x_{1}\}}\cap\overline{\{x_{2}\}}=E_{x_{1}}\cap E_{x_{2}}$.
By construction, $\Sigma({X}^{\dagger})$ is a strictly convex rational polyhedral cone complex. In particular $\Sigma({X}^{\dagger})$ is a $\mathbb{Z}$-PL complex (see Definition A.33). Without danger of confusion we will identify $\Sigma({X}^{\dagger})$ with its underlying topological space $|\Sigma({X}^{\dagger})|=F({X}^{\dagger})(\mathbb{R}_{\geq 0})$.
Because the construction is functorial, we obtain for any morphism $X^{\prime\dagger}\to{X}^{\dagger}$ of log-regular schemes a morphism of cone complexes $\Sigma(X^{\prime\dagger})\to\Sigma({X}^{\dagger})$.
The following lemma shows that proper subdivisions can be pulled back along morphisms of log-regular schemes.
Lemma 2.16 (pullback of a subdivision).
Let $f:X^{\prime\dagger}\longrightarrow X^{\dagger}$ be a morphism of log-regular log-schemes. Then for any proper subdivision $\Sigma_{0}\to\Sigma(X^{\dagger})$ (Definition A.10), there exists a cartesian square (in the fs category) of log-regular log-schemes
[…]
such that $\Sigma(X_{0}^{\dagger})=\Sigma_{0}$ and $\Sigma(X_{0}^{\prime\dagger})\to\Sigma(X^{\prime\dagger})$ is a proper subdivision.
Proof.
By A.21 locally fs fans admit fibered products, so there is a cartesian square
[…]
where $F_{0}=F(\Sigma_{0})$ is the Kato fan associated to $\Sigma_{0}$. Let $\Sigma^{\prime}_{0}$ denote the cone complex associated to $F_{0}^{\prime}$. and let $\varphi:\Sigma(X^{\prime\dagger})\to\Sigma(X^{\dagger})$ denote the associated morphism of cone complexes. For each cone $\sigma^{\prime}$ of $\Sigma(X^{\dagger})$, the cone $\varphi(\sigma^{\prime})$ is a finite union $\cup\sigma_{i}$ of cones $\sigma_{i}$ in $\Sigma_{0}$, because $\Sigma_{0}\to\Sigma(X^{\dagger})$ is a proper subdivision. By construction $\Sigma^{\prime}_{0}$ is obtained by glueing the cones $\phi^{-1}\sigma_{i}$. So by definition $\Sigma_{0}^{\prime}\to\Sigma(X^{\prime\dagger})$ is a proper subdivision (Definition A.10). Therefore so is $F_{0}^{\prime}\to F^{\prime}$.
Now by 2.13 we have log-regular schemes $X_{0}^{\prime\dagger}=X^{\prime\dagger}\times_{F}F_{0}^{\prime}$ and $X_{0}^{\dagger}=X^{\dagger}\times_{F}F_{0}$ with fans $F_{0}^{\prime}$ and $F_{0}$ respectively. By the universal property of 2.13 we obtain a morphism $(X_{0}^{\prime})^{\dagger}\to X^{\dagger}$ such that 2.17 is cartesian. ∎
3. Balanced covers of dual complexes
3.1.
In this section, we study dual complexes of toroidal $S$-schemes. We view these as $\mathbb{Z}$-PL complexes and and prove that finite covers are balanced (Theorem 3.24).
3.2Setup.
Recall the following running assumptions, see Conventions 1.32. Throughout $k^{\circ}$ denotes a discretely valued field with a choice of uniformiser $\varpi$. The scheme $S=\mathrm{Spec}\,k^{\circ}$ is endowed with the standard log-structure $S^{\dagger}$ given by the closed point.
Definition 3.3 (Toroidal $S$-schemes).
We define a toroidal $S$-scheme as a log-scheme ${X}^{\dagger}=({X},\mathscr{M}_{{X}})$ such that ${X}$ is an integral flat separated finite type $S$-scheme.
The flatness assumption is in fact superfluous; indeed, an integral log-scheme ${X}^{\dagger}$ over $S^{\dagger}$ is log-regular at the generic point if and only if the log-structure is generically trivial, and any log-scheme $X^{\dagger}$ over $S^{\dagger}$ satisfies $D_{X}\subset{X}_{s}$, where $D_{X}$ denotes the support (see 2.3).
3.4Assumption.
Throughout, unless explicitly stated otherwise,
$f:X^{\prime\dagger}\to{X}^{\dagger}$
denotes a finite morphism of toroidal $S$-schemes, which will also call a toroidal cover. Note this implies $f^{-1}D_{{X}}\subset D_{X^{\prime}}$.
Example 3.5.
Let us show that in nature, one can often encounter morphisms $f:X^{\prime\dagger}\to{X}^{\dagger}$ as in 3.4; we give four situations below. Let ${X}^{\dagger}$ be a toroidal $S$-scheme.
  1. (1)
    Suppose ${X}^{\dagger}\to S^{\dagger}$ is log-smooth, and $S^{\prime\dagger}\to S^{\dagger}$ is a finite morphism of log-schemes, then $X^{\prime\dagger}\coloneqq{X}^{\dagger}\times^{\mathrm{fs}}S^{\prime\dagger}\to{X}^{\dagger},$ is log-smooth, and hence $X^{\prime\dagger}$ is log-regular. In particular the underlying scheme $X^{\prime}$ is normal, and so coincides with the normalisation of the base change ${X}_{S^{\prime}}={X}\times_{S}S^{\prime}$.
  2. (2)
    Let $S^{\prime\dagger}\to S^{\dagger}$ be a finite log-étale morphism of log-schemes. Then the fs base change $f^{\dagger}:X^{\prime\dagger}\coloneqq{X}^{\dagger}\times_{S^{\dagger}}^{\mathrm{fs}}S^{\prime\dagger}\to{X}^{\dagger}$ is also log-étale. Therefore, since $X^{\dagger}$ is log-regular, so is $X^{\prime\dagger}$. Similarly the underlying scheme $X^{\prime}$ is the normalisation of ${X}_{S^{\prime}}$.
  3. (3)
    Suppose $X^{\prime\dagger}$ is a fine log-scheme equipped with a finite morphism $f:X^{\prime\dagger}\to{X}^{\dagger}$ such that $f$ is étale away from $D_{{X}}$ and tamely ramified over the generic points of $D_{{X}}$. Then $f$ is log-étale by the log-purity theorem of Kato-Mochizuki [Moc99, Theorem B] (also see [GR18, §13.3.43] for a refinement), and hence $X^{\prime\dagger}$ is log-regular.
  4. (4)
    Suppose ${X}$ is regular of dimension $2$ and write $U={X}\setminus D_{{X}}$. Suppose $U^{\prime}\to U$ is a finite étale cover which is cyclic of degree $p$. Then by Abhyankar’s theorem on $p$-cyclic covers of Jungian domains – see Kato’s comments in [Kat94, §12.4] and the references cited there – it follows that there exists a modification $g:{X}_{0}\to{X}$ such that the following conditions (i)-(iii) hold:
    1. (i)
      Let $D_{{X}_{0}}=g^{-1}D_{X}$. Then the log-scheme ${X}_{0}^{\dagger}={X}_{0}(\mathrm{log}\,D_{{X}_{0}})$ is log-regular.
    2. (ii)
      The modification $g$ is an isomorphism when restricted to ${X}_{0}\setminus D_{{X}_{0}}$.
    3. (iii)
      Let $f:X^{\prime}\to{X}_{0}$ be the normalisation of ${X}_{0}$ in the function field of $U^{\prime}$. Then $X^{\prime\dagger}=X^{\prime}(\mathrm{log}\,f^{-1}D_{{X}_{0}})$ is log-regular.
    Note that if (i)-(iii) hold then $f$ is a finite morphism of $2$-dimensional toroidal $S$-schemes (the morphism $f$ is finite because ${X}_{0}$ is excellent).
Remark 3.6 (Kato’s question on simultaneous toroidal resolution).
We mention that Kato [Kat94, §12.4.1] has asked the question whether the assertion of 3.5 (4) remains true for any finite étale cover $U^{\prime}\to U$ and in arbitrary dimension. By the existence of toroidal resolutions (2.13) and Lemma 2.16, we may additionally require ${X}_{0}$ to be regular (as Kato does) without changing the outcome of this question. To the author’s knowledge Kato’s question is still wide open.
3.7Fans.
Throughout the remainder of the article $f:{X}^{\dagger}\to X^{\prime\dagger}$ is as in Assumption 3.4. The constructions of Section 2.7 yield a morphism of Kato fans $F(X^{\prime\dagger})\to F({X}^{\dagger})$. Since the morphism ${X}^{\dagger}\to S^{\dagger}$ provides structural maps $\mathbb{N}\to\mathscr{C}_{F({X}^{\dagger}),x}$, sending $1$ to the class of the uniformiser $[\varpi]\in\mathscr{C}_{F({X}^{\dagger}),x}$, we can and will view $F({X}^{\dagger})$ as a fan over $\mathrm{Spec}\,\mathbb{N}$, and $F(X^{\prime\dagger})\to F({X}^{\dagger})$ as a morphism of fans over $\mathrm{Spec}\,\mathbb{N}$. To make the notation somewhat less heavy, we will also write
$M_{x}\coloneqq\mathscr{C}_{F({X}^{\dagger}),x}^{\mathrm{gp}}.$
Note that by definition, $M_{x}$ is a lattice and $\mathrm{rank}(M_{x})=\mathrm{rank}(x)$.
Lemma 3.8.
Let $x^{\prime}\in F(X^{\prime\dagger})$ and write $x=f(x^{\prime})$. Then $M_{x}\to M_{x^{\prime}}$ is injective and $\mathrm{rank}(M_{x})=\mathrm{rank}(M_{x^{\prime}})$.
Proof.
The map $\mathscr{C}_{F({X}^{\dagger}),x}\to\mathscr{C}_{F(X^{\prime\dagger}),x^{\prime}}$ is local morphism of sharp monoids, which by sharpness is the same as injective monoid morphism (A.18). By the assumption that $x^{\prime}\in F(X^{\prime\dagger})$ and $f:X^{\prime}\to{X}$ is finite, we have a finite morphism of log-strata
$E_{x^{\prime}}=\overline{\{x^{\prime}\}}\to E_{x}=\overline{\{x\}}$
of dimension $\dim\mathscr{O}_{X^{\prime},x^{\prime}}=\dim\mathscr{O}_{{X},f(x)}$, which by log-regularity gives $\mathrm{rank}(x^{\prime})=\mathrm{rank}(x)$. ∎
3.9The dual complex $\mathcal{D}({X}^{\dagger})$.
The dual complex $\mathcal{D}({X}^{\dagger})$ of ${X}^{\dagger}$ is defined as the $\mathbb{Z}$-PL subspace of $\Sigma({X}^{\dagger})$ defined by glueing the polyhedrons
$\sigma_{x}\coloneqq\{a\in\sigma_{\mathscr{C}_{F({X}^{\dagger}),x}}:\alpha([\varpi])=1\}\subset C(x)$
by restricting the face maps of $\Sigma({X}^{\dagger})$. This defines a functor from the category of log-regular log-schemes over $S^{\dagger}$ to $\mathbb{Z}$-PL complexes. Throughout we denote by
$\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})$
the induced morphism of $\mathbb{Z}$-PL complexes. Note that the following are equivalent:
  1. (i)
    The polyhedron $\sigma_{x}$ is nonempty.
  2. (ii)
    The element $[\varpi]\in\mathscr{C}_{F({X}^{\dagger}),x}$ is nonzero.
  3. (iii)
    The point $x$ lies in the special fiber ${X}_{s}$.
Moreover, $\sigma_{x}$ is bounded if and only if the closed stratum $E_{x}$ associated to $x$ (see 2.8) is an intersection of vertical components (components contained in ${X}_{s}$) of $D=D_{{X}}$.
Remark 3.10.
Informally speaking, the points of $\sigma_{x}$ are “$k$-valuations” on the monoid $\mathscr{C}_{F({X}^{\dagger}),x}$, in the sense that they define monoid morphisms $\mathscr{C}_{F({X}^{\dagger},x}\to\mathbb{R}_{\geq 0}$ which are compatible with the normalised valuation of $k$.
Proposition 3.11 (Multiplicity formula).
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a toroidal cover. Let $y^{\prime}\in F(X^{\prime\dagger})$, write $y=f(y)$ and let $x\in\overline{\{y\}}$ be a specialisation of $y$. Then
(3.12) $\left[M_{y^{\prime}}:M_{y}\right]\cdot\left[\kappa(y^{\prime}):\kappa(y)\right]=\sum_{x^{\prime}\in\overline{\{y^{\prime}\}}\cap f^{-1}(x)}\left[M_{{x^{\prime}}}:M_{x}\right]\cdot\left[\kappa(x^{\prime}):\kappa(x)\right].$
Proof.
By induction, we can assume that $\dim_{x}{X}=\dim_{y}{X}+1$. We have induced morphisms of closed strata
[…]
and we can view $x^{\prime}$ and $x$ as codimension $1$ points of $E_{y^{\prime}}$ and $E_{y}$ respectively. By the usual multiplicity formula, see for instance [Liu02, §7.1.38], it follows that
(3.13) $\left[\kappa(y^{\prime}):\kappa(y)\right]=\sum_{x^{\prime}\in\overline{\{y^{\prime}\}}\cap f^{-1}(x)}e(\mathscr{O}_{E_{y^{\prime}},x^{\prime}}/\mathscr{O}_{E_{y},x})\cdot\left[\kappa(x^{\prime}):\kappa(x)\right].$
By Lemma 2.9 we have a diagram with exact rows
00{0}${\mathscr{C}_{E_{y},x}^{\mathrm{gp}}}$${\mathscr{C}_{{X},x}^{\mathrm{gp}}}$${\mathscr{C}_{{X},y}^{\mathrm{gp}}}$00{0}00{0}${\mathscr{C}_{E_{y^{\prime}},x^{\prime}}^{\mathrm{gp}}}$${\mathscr{C}_{X^{\prime},x^{\prime}}^{\mathrm{gp}}}$${\mathscr{C}_{X^{\prime},y^{\prime}}^{\mathrm{gp}}}$00{0}
where $E_{y}$ and $E_{y^{\prime}}$ are equipped with the log-structure obtained by restricting the log-structure of ${X}^{\dagger}$ and $X^{\prime\dagger}$ respectively. By Lemma 3.8 and the snake lemma we obtain
(3.14) $\left[M_{{x^{\prime}}}:M_{x}\right]=\left[M_{{y^{\prime}}}:M_{{y}}\right]\cdot\left[\mathscr{C}_{E_{y^{\prime}},x^{\prime}}^{\mathrm{gp}}:\mathscr{C}_{E_{y},x}^{\mathrm{gp}}\right].$
The desired formula now follows from Equations (3.13) and (3.14) and Lemma 3.15 below. ∎
Lemma 3.15.
Let $R^{\prime}/R$ be a finite extension of DVRs with ramification index $e(R^{\prime}/R)$. Equip $S^{\prime}=\mathrm{Spec}\,R^{\prime}$ and $S=\mathrm{Spec}\,R$ with the standard log-structures $S^{\prime\dagger}$ and $S^{\dagger}$. Then
$[\mathscr{C}^{\mathrm{gp}}_{S,s}:\mathscr{C}_{S,s^{\prime}}^{\mathrm{gp}}]=e(R^{\prime}/R).$
Proof.
Pick a uniformiser $\xi^{\prime}$ of $R^{\prime}$, then $\xi\coloneqq\xi^{\prime e(R^{\prime}/R)}$ uniformises $R$, and $\mathscr{C}_{s}$ (resp. $\mathscr{C}_{s^{\prime}}$) is generated by the class of $\xi$ (resp. $\xi^{\prime}$). ∎
Definition 3.16 (Finite covers of PL spaces).
Let $\varphi:\Sigma^{\prime}\to\Sigma$ be a morphism of PL spaces. We call $\varphi$ a (branched) finite cover if $\varphi$ is surjective with finite fibers and for each polyhedron $\sigma^{\prime}$ of $\Sigma^{\prime}$ we have $\dim\varphi(\sigma^{\prime})=\dim\sigma^{\prime}$.
3.17Lattice index.
Suppose $\varphi:\Sigma^{\prime}\to\Sigma$ is a finite cover of $\mathbb{Z}$-PL complexes. Let $\sigma^{\prime}$ be a polyhedron of $\Sigma^{\prime}$ and let $\sigma=\varphi(\sigma^{\prime})$. Then by assumption that $\dim\sigma=\dim\sigma^{\prime}$ it follows that
$N_{\sigma^{\prime}}\to N_{\sigma}$
is injective, where recall $N_{\sigma}$ is our notation for the lattice of integral points of $\mathbb{R}\sigma$ (as in A.31). The lattice index of $\varphi$ at $\sigma^{\prime}$ is defined as
$\left[N_{\sigma}:N_{\sigma^{\prime}}\right].$
A computation [^4] For any injective group morphism of same-rank lattices $\psi:L^{\prime}\to L$, the dual $\psi^{*}:L^{*}\to(L^{\prime})^{*}$ is also injective and $\mathrm{coker}\psi^{*}\cong\mathrm{coker}\psi$ canonically by the long exact sequence $00=\mathrm{coker}(\psi)^{*}\to L^{*}\to(L^{\prime})^{*}\to\mathrm{Ext}^{1}(\mathrm{coker}(\psi),\mathbb{Z})=\mathrm{coker}(\psi)\to\mathrm{Ext}^{1}(L,\mathbb{Z})=$ 0 . So it follows that $[L:L^{\prime}]=\#\mathrm{coker}\,\psi=\#\mathrm{coker}\,\psi^{*}=[L^{\prime*}:L^{*}]$. shows that
(3.18) $\left[M_{\sigma^{\prime}}:M_{\sigma}\right]=\left[N_{\sigma}:N_{\sigma^{\prime}}\right].$
Informally speaking, the lattice index measures the change of $\mathbb{Z}$-PL functions along $\sigma^{\prime}\mapsto\sigma$.
Note that $\sigma$ is strictly convex if and only if $\sigma^{\prime}$ is so. In that case $(M_{\sigma}^{+})^{\mathrm{gp}}=M_{\sigma}$ where $M_{\sigma}^{+}$ denotes the monoid of non-negative $\mathbb{Z}$-affine functions on $\sigma$ (see A.32).
Definition 3.19 (facet-ridge pairs $(\sigma,\tau)$).
Let $\sigma,\tau$ be polyhedrons of a PL-space $\Sigma$. Then we write $\sigma>\tau$ if and only if $\tau$ is a proper face of $\sigma$. We will also call $\sigma$ an adjacent polyhedron of $\tau$. Suppose $n=\dim\Sigma\geq 1$. A facet is defined as a $n$-dimensional polyhedron and a ridge is defined as a $(n-1)$-dimensional polyhedron. By a facet-ridge pair $(\sigma,\tau)$ we mean a pair of a facet $\sigma\in\Sigma(n)$ and a ridge $\tau\in\Sigma(n-1)$ such that $\sigma>\tau$.
Definition 3.20 (Balanced cover of $\mathbb{Z}$-PL complexes).
Suppose $\varphi:\Sigma^{\prime}\to\Sigma$ is a finite cover of $\mathbb{Z}$-PL complexes. Let $\tau^{\prime}$ be a ridge of $\Sigma^{\prime}$. Then we call $\varphi$ balanced at $\tau^{\prime}$ if the following condition holds. For each facet $\sigma$ adjacent to $\tau$, the following expression is independent of the choice of $\sigma$:
(3.21) $\deg_{\tau^{\prime}}\varphi\coloneqq\sum_{\begin{subarray}{c}\text{facets }\,\sigma^{\prime}\\ (\sigma^{\prime},\tau^{\prime})\mapsto(\sigma,\tau)\end{subarray}}[M_{\sigma^{\prime}}:M_{\sigma}],$
the sum is taken over all facets $\sigma^{\prime}$ adjacent to $\tau^{\prime}$ such that $\varphi:(\sigma^{\prime},\tau^{\prime})\mapsto(\sigma,\tau)$, see Figure 1.11 for a picture. We call $\varphi$ a balanced finite cover of $\mathbb{Z}$-PL complexes if it is balanced at every ridge $\tau^{\prime}$ of $\Sigma^{\prime}$.
If $\varphi$ is balanced at every ridge $\tau^{\prime}$ of $\Sigma^{\prime}$ and $\Sigma$ is connected then we define the degree $\deg\phi$ of $\phi$ as follows. Choose any ridge $\tau$ of $\Sigma$ and set
(3.22) $\deg\varphi\coloneqq\sum_{\tau^{\prime}\mapsto\tau}\deg_{\tau^{\prime}}\varphi,$
where we sum over ridges $\tau^{\prime}$ above $\tau$. It follows from the definitions and the assumption that $\Sigma$ is connected that $\deg\phi$ is independent of the choice of $\tau$.
3.23.
The balancing notion of 3.20 behaves well with proper subdivisions in the following sense. If $\Sigma^{\prime}\to\Sigma$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\Sigma_{0}\to\Sigma$ is a proper subdivision, then $\Sigma^{\prime}_{0}=\Sigma^{\prime}\times_{\Sigma}\Sigma_{0}\to\Sigma_{0}$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\Sigma^{\prime}_{0}\to\Sigma^{\prime}$ is a proper subdivision.
Theorem 3.24.
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a toroidal cover. Then $\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\deg\phi=\deg f$.
Proof.
The map $\phi$ is a finite cover of $\mathbb{Z}$-PL complexes by Lemma 3.8. Let $(\sigma,\tau)$ be a facet-ridge pair of $\mathcal{D}({X}^{\dagger})$, then $(\sigma,\tau)=(\sigma_{x},\sigma_{y})$ for points $x,y\in F({X}^{\dagger})$ with $x\in\overline{\{y\}}$. By Proposition 3.11 it follows that for any ridge $\tau^{\prime}$ above $\tau$ we have $\deg_{\tau^{\prime}}\phi=[L_{\tau^{\prime}}:L_{\tau}][\kappa(\tau^{\prime}):\kappa(\tau)]$, hence $\phi$ is balanced.
It remains to prove that $\deg\phi=\deg f$. For any vertex $v\in\mathcal{D}({X}^{\dagger})(0)$ (a 0-dimensional polyhedron), Lemma 3.15 and equation (3.13) (applied to the generic point of $X^{\prime}$) yield the formula
$\deg f=\sum_{v^{\prime}\in\phi^{-1}(v)}[M_{v^{\prime}}:M_{v}]\cdot[\kappa(v^{\prime}):\kappa(v)].$
Another application of Proposition 3.11 then shows that the last expression equals $\deg\phi$. ∎
Remark 3.25.
A variant of Theorem 3.24 is true without the assumption that $\tilde{k}$ is algebraically closed, but next to the $\mathbb{Z}$-PL structure on $\mathcal{D}({X}^{\dagger})$ one would have to additionally remember the degree $\deg_{\widetilde{k}}{\sigma}\coloneqq\left[\kappa(x):\widetilde{k}\right]$ on each facet $\sigma=\sigma_{x}$ of $\mathcal{D}({X}^{\dagger})$, and change the balancing condition to $\deg_{\tau^{\prime}}\phi=\sum\deg_{\sigma^{\prime}}\phi$ where $\deg_{\sigma^{\prime}}\phi\coloneqq[M_{\sigma^{\prime}}:M_{\sigma}]\deg_{\widetilde{k}}{\sigma^{\prime}}/\deg_{\widetilde{k}}{\sigma}$ (in accordance with Proposition 3.11).
4. Metrisation of $\mathbb{Z}$-PL spaces
4.1Overview.
In this section we observe that any $\mathbb{Z}$-PL space $\Sigma$ is endowed with a canonical piecewise Lebesgue measure $\mu_{\Sigma}$, that we call the conformal measure, such that isometries are isomorphisms. We give explicit descriptions of $\mu_{\Sigma}$ in the case where $\Sigma$ is a dual complex.
4.2First attempt.
Let $\sigma$ be a polyhedron; recall that by convention $\sigma$ carries a $\mathbb{Z}$-affine structure: $\sigma=(|\sigma|,N_{\sigma})$ where $N_{\sigma}$ is the lattice of $\mathbb{Z}$-affine functions. For any polyhedron $\sigma$ let $H_{0}$ be the real vector space spanned by the differences of elements of $\sigma$. Then $H_{0}$ contains the maximal lattice $N_{\sigma}\cap H_{0}$ and therefore $\sigma$ inherits by translation an integral structure. For many purposes this is already a good procedure to equip $\sigma$ with a Lebesgue measure which only depends on the $\mathbb{Z}$-affine structure, it is for instance used in the works [CD12, JN20]. However for our purposes it is slightly more natural to work with a slightly different construction that can distinguish isomorphism types of polyhedra: see Example 4.6 below.
Lemma 4.3.
Suppose $A\in\mathbb{Z}^{n\times m}$ and
  1. (i)
    $\{x\in\mathbb{Z}^{n}:Ax=0\}$ is a full-rank lattice in $\{x\in\mathbb{R}^{n}:Ax=0\}$.
  2. (ii)
    Let $b\in\mathbb{Q}^{n}\setminus\{0\}$ and suppose $\{x\in\mathbb{R}^{m}:Ax=b\}$ is nonempty. Then there exists a minimal positive rational number $\rho\in\mathbb{Q}_{>0}$ such that $Ax=\rho b$ admits an integral solution $x\in\mathbb{Z}^{n}$.
Proof.
(i) By Gaussian elimination we can assume $A$ is upper triangular and so there exists a basis of rational solutions of $Ax=0$. (ii) By Gaussian elimination the system $Ax=b$ admits a rational solution $x\in\mathbb{Q}^{m}$, so by clearing the denominators of $x$ there exists at least one such $\rho\in\mathbb{Q}_{>0}$. Since $\rho b\in\mathbb{Z}^{n}$ the set of such $\rho$ is well-ordered, so there exists a minimum. ∎
Construction 4.4 (Conformal measure).
Let $\sigma=(|\sigma|,N_{\sigma})$ be a polyhedron and suppose $0\not\in\sigma$. let
$H_{\sigma}=\{\lambda(x-y):x,y\in|\sigma|,\lambda\in\mathbb{R}\}$
denote the $\mathbb{Z}$-affine plane in $N_{\mathbb{R}}$ spanned by $|\sigma|$. By Lemma 4.3 (ii) there exists a minimal positive rational number $\rho=\rho_{\sigma}\in\mathbb{Q}_{>0}$ such that
$H_{\rho\sigma}\cap N\neq\emptyset,$
where $\rho\sigma=\{\rho x:x\in\sigma\}$ is the $\rho$-scaling of $\sigma$. Then $H_{\rho\sigma}\cap N$ is a full-rank lattice in $H_{\rho\sigma}$ by Lemma 4.3 (i). Let $\mu_{H_{\rho\sigma}}$ be the unique Lebesgue measure on $H_{\rho\sigma}$ such that $N\cap H_{\rho\sigma}$ has covolume $1$. We write $\mu_{\rho\sigma}$ for the restriction to $\rho\sigma$. Then we define the conformal measure on $\sigma$ as
(4.5) $\mu_{\sigma}\coloneqq\frac{\mu_{\rho\sigma}}{\rho^{\dim\sigma+1}};$
If $0\in\sigma$ we define $\mu_{\sigma}$ as the restriction of the unique Lebesgue measure on $(N_{\sigma})_{\mathbb{R}}$ such that $N_{\sigma}$ has covolume $1$.
We will call $\mu_{\sigma}$ the conformal measure on $\sigma$. The name is justified by Lemma 4.7 below.
Example 4.6.
For instance let $I\subset\mathbb{R}^{n}$ be the line segment given by the equations $12x+2y=$ 1 and $x,y\geq 0$. Then $I$ is $\mathbb{Z}$-PL and the length of $I$ with respect to $\mu_{I}$ is $1/4$.
For another example, the line segment $J\subset\mathbb{R}^{n}$ given by $x+2y=1$ and $x,y\geq 0$ has length $1/2$. Note that the map $\phi:(x,y)\mapsto(2x,y)$ sends $I$ to $J$, and rescales the length by a dilation factor $2$, which is precisely the determinant of $\phi$. This will be generalised in Lemma 4.7 below and motivates our choice of the exponent $\dim\sigma+1$ in Equation (4.5). Indeed, if instead we had chosen the perhaps more natural-looking exponent $\dim\sigma$, then in the above we obtain something that we do not want: $I\to J$ would be an isometry, but $\phi$ is not unimodular (i.e. of determinant 1), and so not an isomorphism.
Lemma 4.7.
Let $\phi:\sigma^{\prime}\to\sigma$ be a morphism of polyhedra (in the sense of Appendix A.31) with $\dim\sigma=\dim\sigma^{\prime}$, then we have that
$\phi_{*}\mu_{\sigma^{\prime}}=\frac{1}{[M_{\sigma^{\prime}}:M_{\sigma}]}\mu_{\sigma}.$
In particular $\phi$ is a $\mathbb{Z}$-PL isomorphism (i.e. a unimodular map) if and only if it is an isometry with respect to the conformal measure, and it is strictly expanding otherwise.
Proof.
We can assume $0\not\in\sigma$. Let $\rho$ (resp. $\rho^{\prime}$) be the minimal $\rho\in\mathbb{Q}_{>0}$ such that $H_{\rho\sigma}\cap N_{\sigma}\neq\emptyset$ (resp. $H_{\rho^{\prime}\sigma^{\prime}}\cap N_{\sigma^{\prime}}\neq\emptyset$. Then we have
$\left(\frac{\rho^{\prime}}{\rho}\right)^{\dim\sigma+1}\mu_{H_{\rho\sigma}}=\frac{\rho^{\prime}}{\rho}\mu_{H_{\rho^{\prime}\sigma}}\overset{\mathrm{def}}{=}\frac{\rho^{\prime}}{\rho}[N\cap H_{\rho^{\prime}\sigma}:N^{\prime}\cap H_{\rho^{\prime}\sigma^{\prime}}]\phi_{*}\mu_{H_{\rho^{\prime}\sigma^{\prime}}}=[N:N^{\prime}]\phi_{*}\mu_{H_{\rho^{\prime}\sigma^{\prime}}}.$
The result now follows by rescaling both sides and using Equation (3.18). ∎
Remark 4.8.
As a corollary of 4.7 we obtain the following change of variables formula: for $\phi:\sigma^{\prime}\to\sigma$ as in the lemma and every integrable $F:\sigma\to\mathbb{R}$ we have $\int_{\sigma}F\mu=[M_{\sigma^{\prime}}:M_{\sigma}]\int_{\sigma^{\prime}}(\phi^{*}F)\mu^{\prime}.$
4.9The conformal measures $\mu_{\Sigma}$.
Let $\Sigma$ be a $\mathbb{Z}$-PL space. By Lemma 4.7 the construction of 4.4 is compatible with isomorphisms and in particular proper subdivisions. Therefore we can glue the measures $\mu_{\sigma}$ for facets $\sigma$ of $\Sigma$ to the conformal measure $\mu_{\Sigma}$ on $\Sigma$.
4.10.
For the remainder of this section we explicitise the construction of $\mu_{\Sigma}$ in the case where $\Sigma=D({X}^{\dagger})$ is a dual complex. As we will see the conformal measure recovers the integral Lebesgue measures of Jonsson-Nicaise [JN20, §5.1] up to a suitable normalisation by the root index, which plays the role of the scaling constant $\rho$, see Remark 4.26.
4.11Notation.
We recall (and simplify) some notation from Section 3: for $x\in F({X}^{\dagger})$, we write $\mathscr{C}_{x}\coloneqq\mathscr{C}_{F({X}^{\dagger}),x}=\mathscr{M}_{{X},x}^{\#}$ for the characteristic monoid, $M_{x}=\mathscr{C}_{x}^{\mathrm{gp}}$ for the characteristic lattice
$N_{x}\coloneqq M_{x}^{*}=(\mathscr{M}_{{X},x}^{\vee})^{\mathrm{gp}}$
for the dual lattice, $r(x)\coloneqq\mathrm{rank}(x)=\mathrm{rank}_{\mathbb{Z}}M_{x}$ for the rank,
$C(x)\coloneqq\sigma_{\mathscr{C}_{x}}=\mathbb{R}_{\geq 0}\mathscr{C}_{x}^{\vee}$
for the dual cone and
$\sigma_{x}=\{\alpha\in C(x):\alpha([\varpi])=1\}$
for the polyhedron of $\mathcal{D}({X}^{\dagger})$ associated to $x$. Note $\dim\sigma_{x}=r(x)-1$ and $M_{\sigma_{x}}=M_{x}$. For any $\rho\in\mathbb{R}$ we write
$H_{\rho,x}:[\varpi]\mapsto\rho$
for the affine hyperplane in $(N_{x})_{\mathbb{R}}$, so by definition $\sigma_{x}=\sigma_{\mathscr{C}_{x}}\cap H_{1,x}$. We write $N_{\rho,x}=N_{x}\cap H_{\rho,x}$. See Figures 4.17 and 4.18 for pictures. Observe the following properties:
  1. (i)
    The hyperplane $H_{0,x}$ is the orthogonal complement of $[\varpi]$ under the natural pairing
    $\langle\cdot,\cdot\rangle:\left(M_{x}\right)_{\mathbb{R}}\times(N_{x})_{\mathbb{R}}\to\mathbb{R}.$
  2. (ii)
    The hyperplanes $H_{\rho,x}$ are parallel translates for varying $\rho\in\mathbb{R}$.
  3. (iii)
    The hyperplane $H_{\rho,x}$ is $\mathbb{Z}$-affine (with respect to $N_{x}$) if and only if $\rho\in\mathbb{Z}$.
4.12Root index.
The root index $\rho_{x}$ at $x$ is defined as the largest positive integer $\rho$ such that the image $[\varpi]$ of $\varpi$ in $\mathscr{C}_{x}$ (equivalently, in $M_{x}$) is divisible by $\rho$. It was earlier introduced in [BN20, §2].
A computation using Bézout-Bachet shows there is a short exact sequence
(4.13) $0\longrightarrow N_{x,0}\longrightarrow N_{x}\overset{\langle[\varpi],\cdot\rangle}{\longrightarrow}\rho_{x}\mathbb{Z}\longrightarrow 0$
So $N_{\rho,x}\neq\emptyset$ if and only if $\rho\in\rho_{x}\mathbb{Z}$.
If $x,y\in F(X^{\dagger})$ and $y$ generalises $x$ then we have $\rho_{x}\mid\rho_{y}$ because $\mathscr{C}_{x}\to\mathscr{C}_{y}$ preserves $[\varpi]$.
4.14Recession cone.
We denote by $\mathrm{rec}(\sigma_{x})$ the recession cone of $\sigma_{x}$, i.e. the cone of vectors $v\in(N_{x})_{\mathbb{R}}$ for which $\sigma_{x}+\mathbb{R}_{\geq 0}v\subset\sigma_{x}$ (also see A.35) The following properties are immediate to verify.
  1. (i)
    The cone $\mathrm{rec}(\sigma_{x})$ is a rational polyhedral cone with associated lattice $N_{0,x}\cap\mathbb{R}|\mathrm{rec}(\sigma_{x})|$.
  2. (ii)
    The cone $\mathrm{rec}(\sigma_{x})$ is trivial if and only if $\sigma_{x}$ is bounded.
  3. (iii)
    The rays of $\mathrm{rec}(\sigma_{x})$ are in one-to-one-correspondence with the rays of $C(x)$ that lie in $H_{0,x}$.
4.15Rays of $C(x)$.
Let $x\in F(X^{\dagger})$. Let $r\in C(x)(1)$ be a ray of $C(x)$. Then $r$ corresponds to a unique component $E_{r}$ of $D_{{X}}$ containing $E_{x}$ by Lemma 2.9 (i). We have two cases:
  1. (i)
    Either $r$ does not lie in $H_{0,x}$, in which case $E_{r}$ is vertical (i.e. a component of ${X}_{s}$), and $r$ defines a ray in $\mathrm{rec}(\sigma_{x})$
  2. (ii)
    Or $r$ defines a unique vertex $v_{r}\coloneqq H_{0,x}\cap r\in\sigma_{x}(0)$ and $E_{r}$ is a horizontal component (i.e. not a vertical component, equivalently $E_{r}$ is flat over $k^{\circ}$). We also write $E_{v_{r}}=E_{r}$. Note that the multiplicity $m_{v_{r}}\coloneqq\mathrm{mult}_{E_{r}}{X}_{s}$ of $E_{r}$ equals $\langle[\varpi],n_{r}\rangle$, where $n_{r}$ is the primitive ray generator of $r$, because $\mathscr{C}_{E_{r}}\cong\mathbb{N}$ is generated by the class of a local equation of $E_{r}$.
4.16.
We picture the situation in Figure 4.17 and Figure 4.18 below, where we write $s(x)\coloneqq\#C(x)(1)$ for the number of rays of the cone $C(x)$. Note $s(x)\geq r(x)$ with equality if and only if $C(x)$ is simplicial. Let $n_{1},\dots,n_{s(x)}\in C(x)(1)$ denote the primitive ray generators of $C(x)$ and let $b(x)=\#\sigma_{x}(0)$ denote the number of vertices of $\sigma_{x}$. After reordering we can assume $n_{1},\dots,n_{b(x)}\not\in H_{0,x}$ (corresponding to the vertices of $\sigma_{x}$) and $n_{b(x)+1},\dots,n_{s(x)}\in H_{0,x}$ (corresponding to the rays of $\mathrm{rec}(\sigma_{x})$).
[…]
Figure 4.17. Picture of a bounded face $\sigma_{x}$ as the intersection of the dual cone $C(x)$ and the plane $H_{1,x}$.
[…]
Figure 4.18. Picture of an unbounded face $\sigma_{x}$ as the intersection of the dual cone $C(x)$ and the plane $H_{1,x}$. The number of vertices of $\sigma_{x}$ is $b(x)$, for some $1\leq b(x)<s(x)$.
Example 4.19.
It follows from the definitions that $\rho_{x}$ divides $\gcd(m_{1},\dots,m_{b(x)})$. If ${X}$ is moreover regular, then $\sigma_{\mathscr{C}_{x}}$ is a regular cone and therefore $\rho_{x}=\gcd(m_{1},\dots,m_{b(x)})$. In general, $\rho_{x}$ can be a strict divisor of $\gcd(m_{1},\dots,m_{b(x)})$ as the following example [^5] The example given in [JN20, §4.3.7] is unfortunately not entirely correct: the multiplicites are equal to 1 instead of $2$ and the monoid is generated by $(1,0),(1,1)$ and $(1,2)$. So in the example of loc. cit. $\rho_{x}=\gcd(m_{1},\dots,m_{b(x)})$ contrary to what is claimed. shows. Suppose ${X}^{\dagger}$ contains a subscheme isomorphic to $(\mathrm{Spec}\,R)^{\dagger}$ where
$R=k^{\circ}[s,t,u,v]/(s\varpi=t^{2},v\varpi=u^{2},ut=sv=\varpi^{2})$
and the log-structure is the standard one induced by $\varpi$. A computation shows that $\mathrm{Spec}\,R$ is of dimension $2$. The special fiber consists of $2$ components given by $(s,t,\varpi)$ and $(u,v,\varpi)$ and these intersect in the singular point $x=(s,t,u,v,\varpi)$. The monoid $\mathscr{C}_{x}$ at $x$ is isomorphic to the monoid generated by $[s]=(1,2),[t]=(1,1),[\varpi]=(1,0),[u]=(1,-1),[v]=(1,-2)$ in $\mathbb{Z}^{2}$, in particular ${X}^{\dagger}$ is log-regular at $x$. The dual cone $C(x)$ is isomorphic to the cone in $\mathbb{R}^{2}$ generated by $[s]^{\vee}=(2,-1)$ and $[v]^{\vee}=(2,1)$. The affine line $H_{1,x}$ is given by $y=1$. It follows that $\rho_{x}=1$ and $m_{1}=(2,-1)\cdot(1,0)=2,m_{2}=(2,1)\cdot(1,0)=2$, so that $\rho_{x}=1\neq 2=\gcd(m_{1},m_{2})$.
Example 4.20.
In general $C(x)$ need not be simplicial: consider for instance the scheme ${X}=\mathrm{Spec}\,k^{\circ}[t,u,v,w]/(\varpi=tu=vw)$ equipped with the standard log-structure, then a computation shows that for the point $x=(t,u,v,w)$ we have $s(x)=4>r(x)=3$ and $\sigma_{x}$ is a rectangle.
Definition 4.21 (determinant of $\sigma_{x}$).
Suppose $C(x)$ is simplicial. Then we define the determinant of $\sigma_{x}$ as the integer $\det(\sigma_{x})=|\det(n_{1},\dots,n_{r(x)})|$. In the literature on toric geometry [Ful93, KKMSD, Oda88] it is also known as the index or multiplicity of $C(x)$.
Definition 4.22 ($\sigma_{x}^{-}$).
For every component $E_{i}$ of $D_{{X}}$, choose a generator $\gamma_{i}$ of $\mathscr{C}_{{X},E_{i}}\cong\mathbb{N}$ (i.e. the class of a local equation of $E_{i}$). Then we let
$\sigma_{x}^{-}=\{\alpha\in\sigma_{x}:\langle\gamma_{i},\alpha\rangle\leq 1\}.$
In other words
$\sigma_{x}^{-}=\left\{\sum_{1\leq i\leq b(x)}a_{i}n_{i}+\sum_{i<b(x)\leq s(x)}b_{i}n_{i}\ :\ a_{i}\in\mathbb{R}_{\geq 0},\,\sum_{1\leq i\leq b(x)}a_{i}=1,\ b_{i}\in[0,1]\right\}.$
By construction, $\sigma_{x}^{-}$ is a bounded polytope and we have $\sigma_{x}^{-}\subset\sigma_{x}$ with equality if and only if $\sigma_{x}$ is bounded. Informally speaking, $\sigma_{x}^{-}$ is obtained by “cutting” $\sigma_{x}$ with the unique hyperplane parallel to $v_{1},\dots,v_{b(x)}$ which passes through $n_{b(x)+1},\dots,n_{s(x)}$
Proposition 4.23.
Suppose $C(x)$ is simplicial. Then we have
(4.24) $\mathrm{vol}_{\mu_{x}}\sigma_{x}^{-}=\frac{\det(\sigma_{x})}{m_{1}\dots m_{b(x)}}.$
Proof.
Let $\rho=\rho_{x}$ be the root index and equip $H_{\rho,x}$ with the Lebesgue measure $\lambda_{\rho,x}$ induced by $N_{\rho,x}=N_{x}\cap H_{\rho,x}$. The vertices of the polyhedron
$\sigma_{\rho,x}\coloneqq C(x)\cap H_{\rho_{x},x}$
coincide with the vectors $n_{i}^{\prime}\coloneqq\rho n_{i}/m_{i}\in N_{\rho,x}$ for $i=1,\dots,b(x)$. Note that $\sigma_{\rho,x}^{-}$ is obtained by scaling $\sigma_{x}^{-}$ along the cone $C(x)$ with scaling factor $\rho$. Hence we compute that
$\mathrm{vol}_{\mu_{x}}\sigma_{x}^{-}={\rho^{-\dim\sigma_{x}-1}}\mathrm{vol}_{\lambda_{\rho,x}}(\sigma_{\rho,x}^{-})={\rho^{-r(x)}}|\det(n_{1}^{\prime},\dots,n_{b(x)}^{\prime},\rho n_{b(x)+1},\dots,\rho n_{r(x)})|=\frac{|\det(n_{1},\dots,n_{r(x)})|}{m_{1}\dots m_{b(x)}}.$
Example 4.25.
In case ${X}^{\dagger}$ is regular, each of the cones $C(x)$ is regular, so Proposition 4.23 reads
$\mathrm{vol}(\sigma_{x}^{-})=\left(m_{1}\dots m_{b(x)}\right)^{-1}.$
In the case of relative curves ($\dim{X}=2$) it follows that the conformal measure induces the potential metric introduced in [BN16, §2].
Remark 4.26.
If $\lambda_{x}$ denotes the integral Lebesgue measure of Jonsson-Nicaise [JN20, §5.1] then it is immediate that $\mu_{\sigma_{x}}=\frac{1}{\rho_{x}}\lambda_{x}.$ So the conformal measure is obtained by normalising the Jonsson-Nicaise measure by the root index.
5. Specialisation of Cartier divisors
5.1Overview.
In this section we study Cartier divisors $\mathscr{D}$ on ${X}$ that are supported on $D_{{X}}$, these can be studied via the PL functions on $\mathcal{D}({X}^{\dagger})$ they induce. For later use we compute some intersection numbers (Proposition 5.13) and we introduce the specialisation $\mathrm{sp}_{*}\mathscr{D}$. The remainder of the section is devoted to an analytic view on specialisations, by means of model metrics. This leads to a natural compactification $\overline{\mathcal{D}({X}^{\dagger})}$ of the dual complex, which we first introduce in an ad hoc fashion.
5.2Assumption.
Throughout this section fix a proper toroidal $S$-scheme ${X}^{\dagger}$.
5.3Support functions.
Recall that for each $x\in F({X}^{\dagger})$, the lattice $M_{x}=M_{\sigma_{x}}=\mathscr{C}_{{X},x}^{\mathrm{gp}}$ is naturally identified (see 2.5) with the lattice of Cartier divisors on $\mathrm{Spec}\,\mathscr{O}_{{X},x}$ supported on $D_{{X}}$. On the other hand $M_{\sigma_{x}}$ is the lattice of $\mathbb{Z}$-affine functions on the dual cone $C(x)$. As in A.33 let us write
$M_{\Sigma({X}^{\dagger})}=\lim_{x\in F({X}^{\dagger})}M_{\sigma_{x}},$
this group coincides the group of $\mathbb{Z}$-PL functions $\Sigma({X}^{\dagger})$ that are linear on the cones of $\Sigma({X}^{\dagger})$. If $\mathscr{D}$ is a Cartier divisor supported on $D_{{X}}$, then the associated $\mathbb{Z}$-PL function is denoted by $F_{\mathscr{D}}$ and is called the support function of $\mathscr{D}$. Explicitly it is described as follows: locally near a stratum $x\in F({X}^{\dagger})$ there exists a $m\in M_{x}$ such that $\mathscr{D}=\mathrm{div}\,m$, and so we have
$F_{\mathscr{D}}|_{\sigma_{x}}=\langle[m],\cdot\rangle$
on $\sigma_{\mathscr{C}_{x}}$. Conversely we can reconstruct $\mathscr{D}$ from $[m]$ near $x$ since a choice of section $M_{x}=\mathscr{C}_{x}^{\mathrm{gp}}\to\mathscr{M}_{{X},x}^{\mathrm{gp}}$ is canonical up to a unit in $\mathscr{O}_{{X},x}^{\times}$ and so produces the same divisor $\mathrm{div}\,m$.
Lemma 5.4 (Cartier divisors and support functions).
The map $\mathscr{D}\to F_{\mathscr{D}}$ establishes a group isomorphism between Cartier divisors (resp. $\mathbb{Q}$-Cartier divisors) supported on $D_{{X}}$ and $M_{\Sigma({X}^{\dagger})}$ (resp. $(M_{\Sigma({X}^{\dagger})})_{\mathbb{Q}}$).
Moreover, effective Cartier divisors (resp. effective $\mathbb{Q}$-cartier divisors) correspond to elements of $M_{\Sigma({X}^{\dagger})}^{+}\coloneqq\lim_{x\in F({X}^{\dagger})}\mathscr{C}_{{X},x}$ (resp. $(M_{\Sigma({X}^{\dagger})}^{+})_{\mathbb{Q}_{>0}}=\lim_{x\in F({X}^{\dagger})}\mathscr{C}_{{X},x}^{\mathbb{Q}_{>0}}$).
Proof.
This follows from the discussion in 5.3. ∎
Remark 5.5.
See [Kat94, §11] for a similar correspondence as in Lemma 5.4 via coherent fractional ideals on fans.
5.6Notation.
By definition, $\mathcal{D}({X}^{\dagger})$ and $\Sigma({X}^{\dagger})$ have the same sets of $\mathbb{Z}$-PL functions and so we let $M_{\mathcal{D}({X}^{\dagger})}=M_{\Sigma({X}^{\dagger})}$, this is the group of functions on $\mathcal{D}({X}^{\dagger})$ that are $\mathbb{Z}$-affine on polyhedra.
Given an element $F\in M_{\mathcal{D}({X}^{\dagger})}$ we write $\mathscr{D}_{F}$ for the associated Cartier divisor as in Lemma 5.4. Note that for all $F,F^{\prime}\in M_{\mathcal{D}({X}^{\dagger})}$ we have $\mathscr{D}_{F+F^{\prime}}=\mathscr{D}_{F}+\mathscr{D}_{F^{\prime}}$ by Lemma 5.4.
Lemma 5.7.
Suppose $F\in M_{\mathcal{D}({X}^{\dagger})}$. Write $\mathscr{D}_{F}$ for the associated Cartier divisor as in Lemma 5.4. Then we have an equality of Weil divisors
$\mathrm{cyc}\,\mathscr{D}_{F}=\sum_{r\in\Sigma({X}^{\dagger})(1)}F_{\mathscr{D}}(n_{r})[E_{r}]=\sum_{v\in\mathcal{D}({X}^{\dagger})(0)}m_{v}F_{\mathscr{D}}(v)[E_{v}]+\sum_{r\in\mathrm{rec}(\mathcal{D}({X}^{\dagger}))(1)}\partial_{r}F[E_{r}],$
where we write:
  • $E_{r}$ for the unique component of $D_{{X}}$ associated to a ray $r\in\Sigma({X}^{\dagger})(1)$ (see 4.15)
  • $n_{r}\in N_{r}\cap r$ for the primitive ray generator of a ray $r$
  • $E_{v}$ for the unique component associated to a vertex $v\in\mathcal{D}({X}^{\dagger})(0)$.
  • $m_{v}\coloneqq\mathrm{mult}_{E_{v}}{X}_{s}$ for the multiplicity of a vertex $v\in\mathcal{D}({X}^{\dagger})(0)$.
  • $\partial_{r}F\coloneqq F(n_{r})$ for any ray ${r\in\mathrm{rec}(\mathcal{D}({X}^{\dagger}))(1)}$ of the recession cone complex.
Moreover, the same assertion holds for $(M_{\mathcal{D}({X}^{\dagger})})_{\mathbb{Q}}$ and $\mathbb{Q}$-Weil divisors.
Proof.
Let $r$ be a ray of $\Sigma({X}^{\dagger})$. Choose a local equation $\mathscr{D}_{F}=\mathrm{div}\,m$ near $E_{r}$, for some $m\in M_{r}$. Note that $M_{r}$ is a $1$-dimensional lattice dual to $n_{r}\mathbb{Z}$. By the description of 5.3 it follows $F(n_{r})=\langle m,n_{r}\rangle=\mathrm{mult}_{E_{r}}\mathscr{D}_{F}$. This gives the first equality.
For the second equality: any ray $r$ of $\Sigma({X}^{\dagger})$ is either contained in [ϖ]superscriptdelimited-[]italic-$[\varpi]^{\perp}$ or not, giving rise to a ray of the recession cone complex of $\mathcal{D}({X}^{\dagger})$ or a vertex $v\in\mathcal{D}({X}^{\dagger})$ respectively (see 4.15). For the latter case, note that $[E_{r}]=[E_{v}]$ and $F(n_{r})=F(v)\langle[\varpi],n_{r}\rangle=F(v)\mathrm{mult}_{E_{v}}\mathrm{div}\,[\varpi]=F(v)m_{v}$. ∎
5.8.
It is immediate to verify that the following are equivalent:
  1. (i)
    The cone complex $\Sigma({X}^{\dagger})$ is simplicial.
  2. (ii)
    For every set of rational numbers $\{q_{r}\}_{r\in\Sigma({X}^{\dagger})(1)}$ there exists a function $F\in(M_{\mathcal{D}({X}^{\dagger})})_{\mathbb{Q}}$ such that $F(n_{r})=q_{r}$.
  3. (iii)
    Any $\mathbb{Q}$-Weil divisor supported on $D_{{X}}$ is $\mathbb{Q}$-Cartier.
In particular if ${X}$ is $\mathbb{Q}$-factorial then $\Sigma({X}^{\dagger})$ is simplicial.
5.9Reminder on intersection products.
Suppose $V$ is any normal Cohen-Macaulay scheme, $\mathscr{D}$ is a cartier divisor and $C$ is a normal curve in $V$ defined over a field $\kappa$. Then $V$ contains no associated points [Liu02, §8.2.15] and so the restriction of Cartier divisors $D|_{E_{\tau}}$ is well-defined. We can therefore define the intersection [^6] One can develop a general refined intersection product of ($\mathbb{Q}$-)Cartier divisors and cycles [Ful84, §20] (lacking intersection products of two arbitrary cycles unless ${X}$ is smooth), but won’t need this. product $\mathscr{D}\cdot C\coloneqq\deg_{\kappa}\mathscr{D}|_{C}\in\mathbb{Z}$, and if $\mathscr{D}$ is only assumed $\mathbb{Q}$-Cartier instead of Cartier we let $\mathscr{D}\cdot C=\frac{1}{e}(e\mathscr{D})\cdot C$ where $e\in\mathbb{Z}_{\geq 1}$ is chosen such that $e\mathscr{D}$ is Cartier. We have the usual rules $(\mathscr{D}+\mathscr{D}^{\prime})\cdot C=\mathscr{D}\cdot C+\mathscr{D}^{\prime}\cdot C$, and $\mathrm{div}\,(f)\cdot C=0$, so that $\mathscr{D}\cdot C$ only depends on the linear equivalence class of $\mathscr{D}$.
We will mainly be interested in the case where $V={X}$ is a toroidal $S$-scheme (which is Cohen-Macaulay by [Kat94, §4.1]), $\kappa=\tilde{k}$ and $C=E_{\tau}$ for some vertical ridge $\tau$ of $\mathcal{D}({X}^{\dagger})$. Note that $\mathscr{D}$ intersects $E_{\tau}$ properly if and only if $F_{\mathscr{D}}|_{\tau}$ is the zero function. Since $\tau$ is vertical, $E_{\tau}$ is a proper $\tilde{k}$-curve (since ${X}$ is assumed proper) and $\mathscr{D}\cdot E_{\tau}$ only depends on the linear equivalence class of $\mathscr{D}$.
We will later need the following outcome of the projection formula (see [Ful84, §20]): let $g:V^{\prime}\to V$ be a proper flat dominant morphism of integral normal $S$-schemes, let $\mathscr{D}$ be a $\mathbb{Q}$-cartier divisor. Suppose $c^{\prime}\in Z_{1}(V^{\prime})$ is a vertical $1$-cycle on $V^{\prime}$ (i.e. a linear combination of curves contained in $V^{\prime}_{s}$). Then (using that $\tilde{k}$ is algebraically closed) we have
$g^{*}{\mathscr{D}}\cdot c^{\prime}=\mathscr{D}\cdot g_{*}c^{\prime}.$
Example 5.10.
For instance if $\mathscr{D}=\mathrm{div}\,(\varpi)$ we have $\mathrm{cyc}\mathscr{D}=[X_{s}]=\sum_{v\in\mathcal{D}({X}^{\dagger})(0)}m_{v}[E_{v}]$ by 5.7. If additionally ${X}$ is $\mathbb{Q}$-factorial then the divisors $[E_{v}]$ for $v\in\mathcal{D}({X}^{\dagger})(0)$ are $\mathbb{Q}$-Cartier, so for every vertical 1-cycle $c\in Z_{1}({X})$ we obtain by the projection formula 5.9 the useful relation
$0=\sum_{v\in\mathcal{D}({X}^{\dagger})(0)}m_{v}[E_{v}]\cdot c.$
5.11Normal vector.
Let $(\sigma,\tau)$ be a facet-ridge pair of a $\mathbb{Z}$-PL complex. Write $m_{\sigma/\tau}$ for the unique generator of $\ker(M_{\sigma}\to M_{\tau})$ such that $\langle m_{\sigma/\tau},\cdot\rangle$ is positive on $\sigma\setminus\tau$. Since $m_{\sigma/\tau}$ kills $M_{\tau}$, the map $\langle m_{\sigma/\tau},\cdot\rangle$ descends to a map $N_{\sigma}/N_{\tau}\to\mathbb{Z}$. Let $n_{\sigma/\tau}$ denote the unique generator of $N_{\sigma}/N_{\tau}$ such that $\langle m_{\sigma/\tau},n_{\sigma/\tau}\rangle>0$. We will also view $n_{\sigma/\tau}$ as a vector of $N_{\tau}$, well-defined up to an element of $N_{\tau}$, and call it the primitive normal vector of $\sigma$ relative to $\tau$.
5.12Notation.
Recall the notion of stars and links from A.37. Suppose $\tau$ is a ridge of $\mathcal{D}({X}^{\dagger})$ and $r\in\overline{\mathrm{star}}(C(\tau))(1)$ is a ray of $\Sigma({X}^{\dagger})$ adjacent to $C(\tau)$, then we write $\sigma\coloneqq\tau+r$ for the unique facet adjacent to $\tau$ such that $r\in C(\sigma)(1)$.
Proposition 5.13.
Suppose $\tau$ is a ridge of $\mathcal{D}({X}^{\dagger})$. Let $r\in\mathrm{link}(C(\tau))(1)$ and $\sigma=\tau+r$. Suppose that the component $E_{r}$ associated to $r$ defines a $\mathbb{Q}$-Cartier divisor (eg. $\mathrm{star}(r)$ is simplicial, see Remark 5.8). Then
(5.14) $[E_{r}]\cdot E_{\tau}=\frac{1}{[N_{\sigma}:N_{\tau}+N_{r}]}.$
Additionally, if $\sigma$ is simplicial, this also equals $\det\tau/\det\sigma$.
Proof.
This resembles a classical fact of toric geometry, see for instance [Ful93, §5.1, p. 98,(2)].
Let $n_{\sigma/\tau}$ be the normal vector as in 5.11. Then $N_{\sigma}=n_{\sigma/\tau}\mathbb{Z}+N_{\tau}$, so
(5.15) $[N_{\sigma}:N_{\tau}+N_{r}]n_{\sigma/\tau}\equiv n_{r}\bmod{N_{\tau}}.$
Suppose $\mathscr{D}=l[E_{r}]$ is Cartier for some $l\in\mathbb{Z}_{>0}$. Note that $\mathrm{supp}\mathscr{D}\cap E_{\tau}=E_{\sigma}$ set-theoretically. So $l[E_{r}]\cdot E_{\tau}=\mathscr{D}\cdot E_{\tau}=\mathrm{ord}_{E_{\sigma}}\mathscr{O}(\mathscr{D})|_{E_{\tau}}$. The support function $F_{\mathscr{D}}$ is given on $\sigma$ by $\langle m,\cdot\rangle$ for some $m\in\ker((M_{\sigma})_{\mathbb{Q}}\to(M_{\tau})_{\mathbb{Q}})$, and we have $\langle m,n_{r}\rangle=l$. It follows from Lemma 2.9 that $E_{\tau}^{\dagger}$ is log-regular with $\mathscr{C}_{E_{\tau},E_{\sigma}}^{\mathrm{gp}}\cong N_{\sigma}/N_{\tau}=n_{\sigma/\tau}\mathbb{Z}$. So by restriction of local equations we obtain that
$\mathrm{ord}_{E_{\sigma}}\mathscr{O}(\mathscr{D})|_{E_{\tau}}=\langle m,n_{\sigma/\tau}\rangle.$
Hence by the previous points we can compute
$[E_{r}]\cdot E_{\tau}=\frac{\langle m,n_{\sigma/\tau}\rangle}{l}=\frac{\langle m,n_{r}\rangle}{l[M_{\sigma}:M_{\tau}+M_{r}]}=\frac{1}{[M_{\sigma}:M_{\tau}+M_{r}]}.$
The final assertion follows immediately from Equation (5.15) and Definition 4.21.
5.16pullback of support functions.
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a finite morphism of proper toroidal $S$-schemes and write $\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})$ for the associated morphism of dual complexes. Let $\mathscr{D}$ a $\mathbb{Q}$-Cartier divisor supported on $D_{{X}}$ with support function $F_{\mathscr{D}}:\mathscr{D}({X}^{\dagger})\to\mathbb{R}$. Then $f^{*}\mathscr{D}$ is supported on $D_{X^{\prime}}$ and has support function
$\phi^{*}F_{\mathscr{D}}\coloneqq F_{f^{*}\mathscr{D}}.$
Lemma 5.17.
Keep notations and assumptions of 5.16, then $\phi^{*}F_{\mathscr{D}}=F_{\mathscr{D}}\circ\phi$.
Proof.
Using Lemma 5.7 we have
$\mathrm{cyc}(f^{*}\mathscr{D})=f^{*}\mathrm{cyc}(\mathscr{D})=\sum_{r\in\Sigma({X}^{\dagger})(1)}F(n_{r})f^{*}[E_{r^{\prime}}]=\sum_{r^{\prime}\in\Sigma(X^{\prime\dagger})(1)}F(n_{r})e(E_{r^{\prime}}/E_{r})[E_{r^{\prime}}].$
By Lemma 3.15 shows $n_{r^{\prime}}=e(E_{r^{\prime}}/E_{r})n_{r}$, and so it follows that $f^{*}\mathscr{D}$ is the $\mathbb{Q}$-Cartier divisor with support function $F_{\mathscr{D}}\circ\phi$. ∎
Definition 5.18 (harmonic and convex functions).
Let $\mathscr{D}$ be a $\mathbb{Q}$-cartier divisor on ${X}^{\dagger}$. We call the support function $F_{\mathscr{D}}$ harmonic (resp. convex (or subharmonic), resp. strongly convex) if for all ridges $\tau$ of $\mathcal{D}({X}^{\dagger})$ we have $\mathscr{D}\cdot E_{\tau}=0$ (resp. $\mathscr{D}\cdot E_{\tau}\geq 0$, resp. $\mathscr{D}\cdot E_{\tau}>0$). This is for instance the case if $\mathscr{D}$ is numerically trivial (resp. nef, resp. ample). The terminology is justified in Section 6.
5.19.
With assumptions as in 5.16, suppose $F_{\mathscr{D}}$ is harmonic (resp. convex, resp. strongly convex), then $\phi^{*}F_{\mathscr{D}}$ is harmonic (resp. convex, resp. strongly convex): this follows from the projection formula of 5.9.
5.20Cycles on $\mathbb{Z}$-PL complexes.
Let $\Sigma$ be a $\mathbb{Z}$-PL complex. We define the group of codimension 1 cycles $Z^{1}(\Sigma)$ as the group of formal sums of ridges $[\tau]$, and effective cycles are those with positive coefficients. Now suppose $\phi:\Sigma^{\prime}\to\Sigma$ is a balanced cover of $\mathbb{Z}$-PL complexes. Then we define pullbacks of cycles via the rule $f^{*}[\tau]=\sum_{\tau^{\prime}}\deg_{\tau^{\prime}}\phi[\tau^{\prime}]$ and extending linearly. Note that for any cycle $W\in Z^{1}(\Sigma)$ we have $\deg f^{*}W=\deg f\deg W$.
Definition 5.21 (Specialisation).
Let $\mathscr{D}$ be a $\mathbb{Q}$-Cartier divisor supported on $D_{{X}}$. Then we define the specialisation of $\mathscr{D}$ as the cycle
(5.22) $\mathrm{sp}_{*}\mathscr{D}\coloneqq\sum_{\tau\text{ vertical}}\frac{\mathscr{D}\cdot[E_{\tau}]}{\mathrm{vol}(\tau)}[\tau]\in Z^{1}(\mathcal{D}({X}^{\dagger})),$
summing over bounded ridges $\tau$ of $\mathcal{D}({X}^{\dagger})$.
5.23.
Note that $\mathrm{sp}_{*}\mathscr{D}$ only depends on the linear equivalence class of $\mathscr{D}$. Also, $\mathrm{sp}_{*}\mathscr{D}$ is trivial (resp. effective) if and only if $F_{\mathscr{D}}$ is harmonic (resp. convex).
Proposition 5.24.
Let $\mathscr{D}\in M_{\mathcal{D}({X}^{\dagger})}$ and $\phi:\mathcal{D}(X^{\prime\dagger})\to\mathcal{D}({X}^{\dagger})$ as in 5.16. Then pullback commutes with specialisation:
(5.25) $\phi^{*}\mathrm{sp}_{*}\mathscr{D}=\mathrm{sp}_{*}f^{*}\mathscr{D}$
Proof.
Let $\tau^{\prime}$ be a ridge of $\mathcal{D}(X^{\prime\prime\dagger})$ and let $\tau=\phi(\tau^{\prime})$. By Lemma 4.7 we have $\mathrm{vol}(\tau)/\mathrm{vol}(\tau^{\prime})=[M_{\tau^{\prime}}:M_{\tau^{\prime}}]$. For any facet $\sigma$ adjacent to $\tau$ we have
$\mathrm{mult}_{E_{\sigma^{\prime}}}f^{*}\mathscr{D}|_{E_{\tau^{\prime}}}=e(E_{\sigma^{\prime}}/E_{\sigma})\mathrm{mult}_{E_{\sigma}}\mathscr{D}|_{E_{\tau}}.$
Therefore by (3.14) we have
$\frac{\mathrm{mult}_{E_{\sigma^{\prime}}}f^{*}\mathscr{D}|_{E_{\tau^{\prime}}}}{\mathrm{vol}\tau^{\prime}}=[M_{\sigma^{\prime}}:M_{\sigma}]\frac{\mathrm{mult}_{E_{\sigma}}\mathscr{D}|_{E_{\tau}}}{\mathrm{vol}\tau}.$
Theorem 3.24 then implies
$\phi^{*}\mathrm{sp}_{*}\mathscr{D}=\sum_{(\sigma^{\prime},\tau^{\prime})}[M_{\sigma^{\prime}}:M_{\sigma}]\frac{\mathrm{mult}_{E_{\sigma}}\mathscr{D}|_{E_{\tau}}}{\mathrm{vol}(\tau)}[\tau^{\prime}]=\sum_{(\sigma^{\prime},\tau^{\prime})}\frac{\mathrm{mult}_{E_{\sigma^{\prime}}}f^{*}\mathscr{D}|_{E_{\tau^{\prime}}}}{\mathrm{vol}(\tau^{\prime})}[\tau^{\prime}]=\mathrm{sp}_{*}f^{*}\mathscr{D}$
where we sum over all facet-ridge pairs $(\sigma^{\prime},\tau^{\prime})$ of $\mathcal{D}(X^{\prime\dagger})$ with $\tau^{\prime}$ bounded. ∎
5.26Horizontal strata.
Suppose that $x\in F({X}^{\dagger})$. Recall from 3.9 that $\sigma_{x}$ is empty if and only if $x$ lies in the special fiber ${X}_{s}$. Suppose $x\not\in{X}_{s}$, we also say $x$ is horizontal and write $F({X}^{\dagger})^{\mathrm{hor}}$ for the horizontal points. Note that the generic point of ${X}$ is also considered to be horizontal. By 2.9 it follows that for all $x\in F({X}^{\dagger})^{\mathrm{hor}}$ we have a strict closed immersion $E^{\dagger}_{x}=E_{x}(\mathrm{log}D_{x})\to X^{\dagger}$ where $D_{x}=E_{x}\setminus E_{x}^{\circ}$ and $E^{\dagger}_{x}/S^{\dagger}$ is again a proper toroidal $S^{\dagger}$-scheme 777if $x$ is vertical this is not the case, because if we equip $E_{x}$ with the pullback log-structure then we obtain a so-called hollow log-scheme over the log-point $\tilde{k}$.. In particular we have an associated dual complex $\mathcal{D}(E^{\dagger}_{x})$.
5.27compactified dual complex.
We write
$\overline{\mathcal{D}}(X^{\dagger})\coloneqq\sqcup_{x\in F(X^{\dagger})^{\mathrm{hor}}}\mathcal{D}(E_{x}^{\dagger}).$
We will see below in 5.31 that $\overline{\mathcal{D}}(X^{\dagger})$ admits a natural compact topology, such that each $\mathcal{D}(E_{x}^{\dagger})\subset\overline{\mathcal{D}}(X^{\dagger})$ is a subspace, so we call $\overline{\mathcal{D}}(X^{\dagger})$ the compactified dual complex – beware it is not a $\mathbb{Z}$-PL complex but rather a $\mathbb{Z}$-PM complex (see Remark 5.34 below).
The specialisation map extends to $\overline{\mathcal{D}}(\mathfrak{X}^{\dagger})$ by amalgamating the various specialisation maps for $x\in F(X^{\dagger})^{\mathrm{hor}}$:
(5.28) $\overline{\mathrm{sp}}_{*}\mathscr{D}\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\overline{\mathrm{sp}}_{*}(\mathscr{D}|_{E_{x}})\in Z^{1}(\overline{\mathcal{D}}(X^{\dagger}))\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}Z^{1}(\mathcal{D}(E_{x}^{\dagger}).$
5.29Notation for analytifications.
Let $Z$ be a finite type $k$-scheme. Then recall the Berkovich analytification $Z^{\mathrm{an}}$ is canonically identified with the set of bounded $k$-valuations $|\cdot(z)|:k(Z)\to\mathbb{R}_{\geq 0}$ of rank one (written multiplicatively) with the topology of pointwise convergence [Ber90], and we have a canonical map $\ker:Z^{\mathrm{an}}\to Z$ that admits a continuous section that we view as an inclusion $Z\subset Z^{\mathrm{an}}$.
5.30Skeleta.
Let ${X}^{\dagger}$ be a proper toroidal $S$-scheme. Write $X={X}_{k}$ for the generic fiber, note it is a proper $k$-variety that is regular away from $(D_{{X}})_{k}$.
By the theory of admissible expansions [BM19, §3], [MN15, §2.4] (not reviewed here) there exists a canonical topological embedding
$\iota_{{X}}:\mathcal{D}({X}^{\dagger})\to({X}\setminus D_{{X}})^{\mathrm{an}}.$
We call the image of $\iota_{{X}}$ the skeleton $\mathrm{sk}({X}^{\dagger})$ of ${X}^{\dagger}$. By loc. cit. the embedding $\iota_{{X}}$ admits a continuous retraction [^8] It is plausible that $\rho_{{X}}$ is a strong deformation retract, but we do not study this.
$\rho_{{X}}:({X}\setminus D_{{X}})^{\mathrm{an}}\to\mathrm{sk}({X}^{\dagger}),$
which is described as follows: let $\mathrm{sp}_{{X}}:X^{\mathrm{an}}\to{X}_{s}$ be the specialisation map, and let
$\xi:({X}\setminus D_{{X}})^{\mathrm{an}}\to F({X}^{\dagger})$
be the map which sends $z$ to the unique minimal stratum containing $\mathrm{sp}_{{X}}(z)$, one verifies that $\xi$ is continuous. Then an element $z=|\cdot(z)|\in({X}\setminus D_{{X}})^{\mathrm{an}}$ gives an additive monoid $k$-valuation
$\rho(z):\mathscr{C}_{{X},\xi(z)}\to\mathbb{R}_{\geq 0}:m\mapsto\mathrm{log}_{|\varpi|}|m(z)|,$
that is, an element
$\rho(z)\in\sigma_{\xi(z)}\subset\mathcal{D}({X}^{\dagger})\overset{i_{X}}{\cong}\mathrm{sk}({X}^{\dagger}),$
also see Remark 3.10. Note that since we have removed the divisor $D_{{X}}$ it follows that $m(z)$ is never zero and so $\rho(z)$ is well-defined.
Recall that by Section 3.20 the dual complex $\mathcal{D}({X}^{\dagger})$, and thus $\mathrm{sk}({X}^{\dagger})$ carries a $\mathbb{Z}$-PL structure, and thus by Section 4 also a canonical conformal measure $\mu_{\mathrm{sk}({X}^{\dagger})}$. Volumes will always be taken with respect to this measure. If $x\in F({X}^{\dagger})$ is a log-stratum of ${X}^{\dagger}$, by a slight abuse we also denote by $\sigma_{x}$ the associated polyhedron of $\mathrm{sk}({X}^{\dagger})$. Recall that the lattice of $\mathbb{Z}$-PL functions $M_{\sigma_{x}}$ coincides with the group of Cartier divisors on $\mathscr{O}_{{X},x}$ supported on $D_{{X},x}$.
5.31Compactified skeleta.
The open strata $E^{\circ}_{x}$ for $x\in F({X}^{\dagger})^{\mathrm{hor}}$ are disjoint and cover the generic fiber $X$, and by 5.30 we have continuous skeleton retraction maps
$\rho_{E_{x}^{\dagger}}:(E_{x}^{\circ})^{\mathrm{an}}\to\mathrm{sk}(E^{\dagger}_{x})\subset(E_{x}^{\circ})^{\mathrm{an}}.$
For all horizontal $x$ we have closed embeddings $E_{x}^{\mathrm{an}}\to X^{\mathrm{an}}$, and so we obtain a retraction
(5.32) $\overline{\rho}=\overline{\rho}_{{X}^{\dagger}}:X^{\mathrm{an}}\longrightarrow\overline{{\mathrm{sk}}}({X}^{\dagger})\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\mathrm{sk}(E^{\dagger}_{x}).$
We call $\overline{\rho}$ the compactified retraction to the compactified skeleton $\overline{\mathrm{sk}}({X}^{\dagger})$. Since the topological boundary of $\mathrm{sk}({X}^{\dagger})$ is contained in $D_{{X}}^{\mathrm{an}}$, it follows by induction that $\overline{\mathrm{sk}}({X}^{\dagger})$ is the closure of $\mathrm{sk}({X}^{\dagger})$ and $\overline{\rho}$ is continuous.
In what follows we give a more intrinsic description of $\overline{\rho}$ when restricted to $X^{\mathrm{an}}$. Let $z\in X^{\mathrm{an}}$. Write
$\mathfrak{p}_{z}=\{m\in\mathscr{C}_{{X},\xi(z)}:|m(z)|=0\}.$
Then either $\mathfrak{p}_{z}=\mathscr{C}_{{X},\xi(z)}$ or $\mathfrak{p}_{z}$ is a prime ideal of the monoid $\mathscr{C}_{{X},\xi(z)}$. Recall that prime ideals of $\mathscr{C}_{\xi(z)}$ correspond to strata containing $\xi(z)$. In fact, $\mathfrak{p}_{z}$ corresponds to the minimal stratum containing $\ker(z)$, which is necessarily horizontal, we denote this stratum by $E_{z}$. By 2.9 (v) (also see Remark 2.12) we have
$\mathscr{C}_{{X},\xi(z)}/\mathfrak{p}_{z}\cong\mathscr{C}_{E_{z},\xi(z)}^{0}.$
Therefore, the multiplicative valuation $|\cdot(z)|$ induces a well-defined additive valuation
(5.33) $\mathscr{C}_{E_{z},\xi(z)}\to\mathbb{R}_{\geq 0}:m\mapsto\mathrm{log}_{|\varpi|}|m(z)|;$
which defines the element $\overline{\rho}(z)$ of $\mathcal{D}(E_{z}^{\dagger})$. The point is that $|m(z)|\neq 0$ for all $m\in\mathscr{C}_{E_{z},\xi(z)}$ by construction of $E_{z}$.
It follows from the constructions that $\rho$ and $\overline{\rho}$ are functorial with respect to toroidal covers.
Remark 5.34 ($\mathbb{Z}$-PM spaces).
The compactified skeleton $\overline{sk}({X}^{\dagger})$ is an example of an integral piecewise monomial space (short: $\mathbb{Z}$-PM space), these will be studied systematically in upcoming work of Ducros-Thuillier (in preparation). In short, $\mathbb{Z}$-PM spaces are multiplicative extensions of $\mathbb{Z}$-PL spaces: they are locally cut out in $\mathbb{R}_{\geq 0}$ by equations of the form $cx_{1}^{a_{1}}\dots x_{n}^{a_{n}}\leq 1$ where $c\in|k^{\times}|$ and $a_{i}\in\mathbb{Z}$, where the coordinates are possibly 00 – informally speaking this allows to study the boundary of $\mathbb{Z}$-PL spaces “at infinity”.
5.35Support functions as model metrics.
Let $\mathscr{D}$ be a Cartier divisor on ${X}$ supported on $D_{{X}}$. Then we can extend the definition of the support function $F_{\mathscr{D}}$ to all of ${X}^{\mathrm{an}}$ as follows:
$F_{\mathscr{D}}:{X}^{\mathrm{an}}\to\mathbb{R}_{\geq 0}\cup\{\infty\}:z\mapsto\mathrm{log}_{|\varpi|}\|1\|_{\mathscr{D},z},$
here $\|\cdot\|_{\mathscr{D}}$ denotes the model metric associated to the line bundle $\mathscr{O}(\mathscr{D})$ – it is characterised as follows: if $f$ is local equation of $\mathscr{D}$ near $\xi(z)\in F({X}^{\dagger})$, then we have
$F_{\mathscr{D}}(z)=\mathrm{log}_{|\varpi|}|f(z)|.$
Note that the previous expression only depends on the image of $f$ in $\mathscr{C}_{{X},\xi(z)}$, and so it follows that the support function $F_{\mathscr{D}}$ factors through the retraction map $\overline{\rho}$.
6. Harmonicity
6.1Overview.
In this section we introduce $\mathbb{Z}$-PL tropical complexes, these are, as the name suggests, a $\mathbb{Z}$-PL enhancement of the weak [^9] The tropical complexes of [Car20] are required to satisfy a positivity condition for the local Hodge index, this condition will be of no importance for the current work, so we will omit the adjective “weak” in the sequel. tropical complexes of [Car20]. In short, these are simplicial $\mathbb{Z}$-PL complexes additionally decorated with some structure constants called $\alpha$-numbers, modeled after the intersection numbers $[E_{r}]\cdot E_{\tau}$ for a ridge $\tau$ and a ray $r\in C(\tau)(1)$. Any dual complex $\mathcal{D}({X}^{\dagger})$ of a $\mathbb{Q}$-factorial toroidal $S$-scheme is a $\mathbb{Z}$-PL tropical complex.
Given a PL function $F$ on a $\mathbb{Z}$-PL tropical complex $\Sigma$, we give a definition of the slopes $\partial_{\sigma/\tau}$ for any facet-ridge pair $(\sigma,\tau)$, and the Laplacian of $F$ is defined as $\Delta(F)=\sum_{(\sigma,\tau)}\partial_{\sigma/\tau}[\tau]$, summing over facet-ridge pairs with $\tau$ vertical. We prove that balanced covers are harmonic morphisms in the sense that harmonic functions (those satisfying $\Delta(F)=0$) are preserved via pullback (Theorem 6.22).
6.2Assumptions for this section.
In this section let ${X}^{\dagger}\to X^{\prime\dagger}$ denote a finite cover of proper $\mathbb{Q}$-factorial toroidal $S$-schemes.
6.3Computations.
As before let $\Sigma({X}^{\dagger})$ for the cone complex associated to ${X}^{\dagger}$, note that $\Sigma({X}^{\dagger})$ is simplicial by the hypothesis that ${X}$ is $\mathbb{Q}$-factorial, see Remark 5.8. Let $\tau$ be a ridge of $\mathcal{D}({X}^{\dagger})$ and let $C(\tau)$ be the cone generated by $\tau$ in $\Sigma({X}^{\dagger})$. For any ray $r\in C(\tau)(1)$ write $\alpha_{r,t}=-[E_{r}]\cdot E_{\tau}\in\mathbb{Q}$. For any vertex $v\in\tau(0)$ write $\alpha_{v,\tau}=\alpha_{r_{v},\tau}$ where $r_{v}$ is the ray in $\Sigma({X}^{\dagger})$ generated by $v$. We define the bounded multiplicity of $\tau$ as
$\mathrm{mult}_{b}\tau\coloneqq\sum_{v\in\tau(0)}m_{v}\alpha_{v,\tau}.$
By Example 5.10 and Proposition 5.13 we then have
(6.4) $\mathrm{mult}_{b}\tau=\sum_{v\in\mathrm{link}(\tau)(0)}\frac{m_{v}\det(\tau^{-})}{\det(\sigma^{-})}$
For every $F\in M_{\Sigma({X}^{\dagger})}$ we compute, using Proposition 5.13 again, that
(6.5) $\mathscr{D}_{F}\cdot E_{\tau}$ $=\sum_{r\in\Sigma(1)}F(n_{r})[E_{r}]\cdot E_{\tau}=\sum_{r\in(\mathrm{link}_{\Sigma}C(\tau))(1)}F(n_{r})[E_{r}]\cdot E_{\tau}+\sum_{r\in C(\tau)(1)}F(n_{r})[E_{r}]\cdot E_{\tau}$
We can split the last sum in two parts $(1)+(2)$, the first part being
(6.6) $(1)$ $=\sum_{r\in(\mathrm{link}_{\Sigma}C(\tau))(1)}F(n_{r})[E_{r}]\cdot E_{\tau}$
(6.7) $=\sum_{v\in(\mathrm{link}_{\mathcal{D}}\tau)(0)}m_{v}F(v)[E_{v}]\cdot E_{\tau}-\sum_{v\in\tau(0)}m_{v}\alpha_{v,\tau}[E_{v}]$
(6.8) $=\sum_{v\in(\mathrm{link}_{\mathcal{D}}\tau)(0)}m_{v}[E_{v}]\cdot E_{\tau}\left(F(v)-\frac{\sum_{v\in\tau(0)}m_{v}\alpha_{v,\tau}[E_{v}]}{\mathrm{mult}_{b}\tau}\right)$
And the second part is given by
(6.9) $(2)=$ $\sum_{r\in\mathrm{link}_{\mathrm{rec}(\mathcal{D})}\mathrm{rec}(\tau)(1)}\partial_{r}F[E_{r}]\cdot E_{\tau}-\sum_{r\in\mathrm{rec}(\tau)(1)}\partial_{r}F\alpha_{r,\tau}$
(6.10) $=\sum_{r\in\mathrm{link}_{\mathrm{rec}(\mathcal{D})}\mathrm{rec}(\tau)(1)}[E_{r}]\cdot E_{\tau}\left(\partial_{r}F-\frac{\sum_{r\in\mathrm{rec}(\tau)(1)}\partial_{r}F\alpha_{r,\tau}}{\mathrm{mult}_{u}\tau}\right),$
where we define the unbounded multiplicity of $\tau$ as
$\mathrm{mult}_{u}\tau\coloneqq\sum_{r\in\mathrm{link}_{\mathrm{rec}(\mathcal{D})}\mathrm{rec}(\tau)(1)}[E_{r}]\cdot E_{\tau}=\sum_{r\in\mathrm{link}_{\mathrm{rec}(\mathcal{D})}\mathrm{rec}(\tau)(1)}\det((\tau+r)^{-})/\det(\tau^{-}).$
Remark 6.11.
Our definition of $\mathrm{mult}_{b}(\tau)$ corresponds to $\deg_{b}(\Delta_{\tau})$ in [GRW16, 6.4] - we use $\mathrm{mult}$ rather than $\deg$ to avoid confusion with the local degree of a finite cover.
Definition 6.12 ($\mathbb{Z}$-PL tropical complex).
A $\mathbb{Z}$-PL tropical complex $(\Sigma,\alpha)$ is a simplicial $\mathbb{Z}$-PL complex $\Sigma$ endowed with structure constants $\alpha_{r,\tau}\in\mathbb{Q}$ for any ray-ridge pair $(r,\tau)$ with $r\in C(\tau)(1)$ such that 6.4 holds for any ridge $\tau$.
A balanced finite cover of $\mathbb{Z}$-PL tropical complexes $\phi:(\Sigma^{\prime},\alpha^{\prime})\to(\Sigma,\alpha)$ is a balanced finite cover of $\mathbb{Z}$-PL complexes such that additionally for each ridge $\tau^{\prime}$ of $\Sigma^{\prime}$ the conclusion of the Lemma 6.13 below holds (where we let $[E_{\tau^{\prime}}:E_{\tau}]=\deg_{\tau^{\prime}}\phi/[M_{\tau^{\prime}}:M_{\tau}]$).
Lemma 6.13.
Let $\tau^{\prime}$ be a ridge of $X^{\prime\dagger}$. Suppose $r^{\prime}\in C(\tau^{\prime})(1)$, and let $r=\phi(r^{\prime})$ and $\tau=\phi(\tau^{\prime})$, then
(6.14) $\alpha_{r^{\prime},\tau^{\prime}}=[E_{\tau^{\prime}}:E_{\tau}]\alpha_{r,\tau}.$
Proof.
This follows from the projection formula (5.9) and the fact that $\{r^{\prime}\}=f^{-1}(r)\cap C(\tau^{\prime})(1)$. ∎
Definition 6.15 (Slopes).
Let $(\Sigma,\alpha)$ be a $\mathbb{Z}$-PL tropical complex and let $F\in(M_{\Sigma})_{\mathbb{Q}}$ be a PL function affine on the polyhedra of $\Sigma$. For any facet-ridge pair $(\sigma,\tau)$, let $r\in C(\sigma)(1)\setminus C(\tau)(1)$ be the unique ray of $C(\sigma)$ not in $C(\tau)$ (such an $r$ is unique since $\Sigma$ is assumed simplicial). Then we have two cases:
  • $E_{r}$ is vertical: then we define
    $\partial_{\sigma/\tau}F\coloneqq\frac{1}{\mathrm{vol}(\sigma^{-})}\left(F(v_{r})-\frac{\sum_{w\in\tau(0)}m_{w}\alpha_{w,\tau}F(w)}{\deg_{b}\tau}\right).$
  • $E_{r}$ is horizontal: then we define
    $\partial_{\sigma/\tau}F\coloneqq\frac{1}{\mathrm{vol}(\sigma^{-})}\left(\partial_{n_{r}}F-\frac{\sum_{s\in C(\tau)(1)}\alpha_{s,\tau}\partial_{n_{s}}F}{\deg_{u}\tau}\right).$
Definition 6.16 (Discrete Laplacian).
Let $(\Sigma,\alpha)$ be a $\mathbb{Z}$-PL tropical complex. Let $F:\Sigma\to\mathbb{R}$ be linear on the faces of $\Sigma$. Then the Laplacian of $F$ is defined as the cycle
(6.17) $\Delta(F)\coloneqq\sum_{(\sigma,\tau);\ \tau\text{ bounded }}\partial_{\sigma/\tau}F[\tau],$
summing over facet-ridge pairs $(\sigma,\tau)$ of $\mathcal{D}(X^{\dagger})$ such that $\tau$ is bounded (equivalently $E_{\tau}$ is vertical). We call $F$ harmonic near $x\in\Sigma$ in case for all ridges $\tau$ containing $x$, we have
(6.18) $\sum_{\sigma}\partial_{\sigma/\tau}F=0$
Theorem 6.19 (Poincaré-Lelong Slope formula).
Let $F\in M_{\Sigma}$ where $\Sigma=\mathcal{D}({X}^{\dagger})$ is a dual complex. Then
(6.20) $\mathrm{sp}_{*}\mathscr{D}_{F}=\Delta(F)$
Proof.
The proof is carried out in the computations of  6.3: there we compared the $[\tau]$-coefficients of both sides, and the definitions of the slopes are chosen such that equality holds. ∎
Remark 6.21.
For comparison, note that the slope formula of [GRW16, §6.8] deals with algebraically closed ground fields and principal divisors.
Theorem 6.22 (Harmonic morphisms).
Let $\phi:(\Sigma^{\prime},\alpha^{\prime})\to(\Sigma,\alpha)$ be a balanced finite cover of $\mathbb{Z}$-PL tropical complexes. Then
(6.23) $\phi^{*}\Delta(F)=\Delta(F\circ\phi)$
In particular, $\phi$ is a harmonic morphism in the sense that harmonic functions pull back to harmonic functions.
Proof.
We have
$\phi^{*}\Delta_{*}F_{\mathscr{D}}=\sum_{(\sigma^{\prime},\tau^{\prime})}[M_{\sigma^{\prime}}:M_{\sigma}]\partial_{\sigma^{\prime}/\tau^{\prime}}[\tau^{\prime}],$
summing over facet-ridge pairs $(\sigma^{\prime},\tau^{\prime})$ of $\mathcal{D}(X^{\prime\dagger})$ with $\tau^{\prime}$ bounded. So it suffices to prove that $[M_{\sigma^{\prime}}:M_{\sigma}]\partial_{\sigma^{\prime}/\tau^{\prime}}=\partial_{\sigma/\tau}$. By 4.7 it follows that $[M_{\sigma^{\prime}}:M_{\sigma}]=\mathrm{vol}(\sigma^{-})/\mathrm{vol}(\sigma^{\prime-})$, and so by Definition of $\sigma_{\sigma/\tau}$ (see 6.15) it suffices to prove that $a_{v^{\prime},\tau^{\prime}}/a_{v,\tau}$ is independent of the choice of vertex $v^{\prime}\in\tau^{\prime}(0)$: this is implied by the definition of a balanced morphism of $\mathbb{Z}$-PL tropical complexes, see 6.12. ∎
Remark 6.24.
The case of dual complexes follows more directly from the slope formula (Theorem 6.20), Proposition 5.24 and Lemma 5.17, since these give
$\phi^{*}\Delta(F)=\phi^{*}\mathrm{sp}_{*}\mathscr{D}_{F}=\mathrm{sp}_{*}f^{*}\mathscr{D}_{F}=\Delta(F_{f^{*}\mathscr{D}_{F}})=\Delta(F\circ\phi).$
Remark 6.25.
Both Theorems 6.20 and 6.22 extend to the compactified skeleta (see 5.31) in the obvious way: we have $\overline{\mathrm{sp}}_{*}\mathscr{D}_{F}=\overline{\Delta}(F)$ where we let $\overline{\Delta}(F)\coloneqq\bigsqcup_{x\in F({X}^{\dagger})^{\mathrm{hor}}}\Delta(F|_{\mathcal{D}(E_{x}^{\dagger})})\in Z^{1}(\overline{\mathcal{D}}({X}^{\dagger}))$, and we have $\phi^{*}\overline{\Delta}(F)=\overline{\Delta}(F\circ\phi)$. We leave the details to the interested reader.
7. Complements on log-canonical sheaves
7.1Overview.
In this section we recall the calculus of logarithmic differential forms for use in Section 8. Theorem 7.28 gives an elementary proof of a formula comparing canonical sheaves to log-canonical sheaves.
7.2Definition of logarithmic differentials.
Suppose $f^{\dagger}:X^{\dagger}=(X,\mathscr{M}_{X})\to Y^{\dagger}=(Y,\mathscr{M}_{Y})$ is a morphism of log-schemes. The $\mathscr{O}_{X}$-sheaf $\Omega_{X^{\dagger}/Y^{\dagger}}$ is defined as the universal log-derivation $(\mathrm{d},\mathrm{dlog})$ satisfying the rules $\mathrm{d}\,m=m\,\mathrm{dlog}(m)$ and $\mathrm{dlog}(n)=0$ for all sections $m\in\mathscr{M}_{X}$, $n\in f^{-1}\mathscr{M}_{y}$. In fact $\Omega_{X^{\dagger}/Y^{\dagger}}$ is the quotient of $\Omega_{X/Y}\oplus\left(\mathscr{O}_{X}\otimes\mathscr{M}_{X}^{\mathrm{gp}}\right)$ by the aforementioned relations [Ogu18, Proof of IV.1.2.4]. Note that if $f$ is locally of finite type and the log-structures are coherent, then $\Omega_{X^{\dagger}/Y^{\dagger}}$ is a finitely generated $\mathscr{O}_{X}$-module. Similar as for usual schemes, formation of log-Kähler differentials is compatible with arbitrary base change in the category of log-schemes as well in the categories of fine log-schemes or fs log-schemes [Ogu18, IV.1.2.15].
7.3log-cotangent sequences..
We have the following logarithmic analogues of the usual cotangent sequences for Kähler differentials: suppose $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is a morphism of log-schemes over a log-scheme $S^{\dagger}$, then we have an exact sequence of $\mathscr{O}_{X}$-modules [Ogu18, Proposition IV.2.3.1]
(7.4) $f^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\longrightarrow\Omega_{X^{\dagger}/S^{\dagger}}\longrightarrow\Omega_{X^{\dagger}/Y^{\dagger}}\longrightarrow 0.$
If $f^{\dagger}$ is log-smooth (see 2.4), then (7.4) is left-exact and locally split [Ogu18, Theorem IV.3.2.3], and if additionally $X^{\dagger},Y^{\dagger}$ are fine, then $\Omega_{X^{\dagger}/Y^{\dagger}}$ is locally free of finite rank.
7.5.
A morphism of log-schemes $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is called a strict closed immersion (resp. strict regular closed immersion) if $f^{\dagger}$ is strict and $f$ is a closed immersion (resp. regular closed immersion). If $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is a strict closed immersion of log-schemes over $S^{\dagger}$, defined by a coherent ideal sheaf $\mathscr{I}$, then we denote by $\mathscr{C}_{X/Y}\coloneqq\mathscr{I}/\mathscr{I}^{2}$ the conormal sheaf of $f$, and we have the cotangent exact sequence [Ogu18, Theorem IV.2.3.2]
[…]
where $\delta$ sends a section $s\in\mathscr{C}_{X/Y}$ to the class of $(\mathrm{d}s,1)$.
Lemma 7.7 (Hyodo-Kato).
Let $f:X^{\dagger}\to Y^{\dagger}$ be a morphism locally of finite type of fine (resp. fs) log-schemes. Then étale-locally on $X$ there exists a fine (resp. fs) log-scheme $W^{\dagger}$ and a factorisation $X^{\dagger}\to W^{\dagger}\to Y^{\dagger}$ where $X^{\dagger}\to W^{\dagger}$ is a strict closed immersion and $W^{\dagger}\to Y^{\dagger}$ is log-smooth.
Proof.
This follows from [HK94, §2.9.2,§2.9.3] in the fine case and the proof also works in the fs case as we explain now. If in [HK94, 2.9.2] the charts $P,Q$ are chosen fs, then so are $Z_{1}^{\dagger}$ and $Z_{2}^{\dagger}$ since they admit charts by $P$ and $Q\oplus\mathbb{N}^{r}$ for some $r\in\mathbb{N}$. The proof of [HK94, 2.9.3] relies on [Kat89, 4.10], and if in the proof of [Kat89, 4.10] the chart $h:P\to Q$ is fs then so is the monoid $Q^{\prime}=(h^{\mathrm{gp}})^{-1}(P)\subset Q^{\mathrm{gp}}$: indeed, if $q\in Q^{\mathrm{gp}}$ and $q^{n}\in Q^{\prime}$ for some $n\in\mathbb{N}$ then $h(q)^{n}=h(q^{n})\in P$ implies $q\in Q^{\prime}$.
Definition 7.8 (Log-l.c.i.).
A morphism of fs log-schemes $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is called log-l.c.i. (short for log-locally of complete intersection) if étale locally on $X$, the morphism $f^{\dagger}$ factors as a strict regular closed immersion followed by a log-smooth morphism.
Lemma 7.9.
Suppose $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is a morphism locally of finite type of log-regular log-schemes. Then $f^{\dagger}$ is log-l.c.i.
Proof.
Since the statement is étale-local on $X$, we may assume there exists a factorisation $X^{\dagger}\to W^{\dagger}\to Y^{\dagger}$ as in Lemma 7.7. Since $W^{\dagger}\to Y^{\dagger}$ is log-smooth, it follows that $W^{\dagger}$ is log-regular and $X^{\dagger}\to W^{\dagger}$ is a strict closed immersion of log-regular log-schemes, and from [Kat94, 4.2] it follows that $X^{\dagger}\to W^{\dagger}$ is a strict regular closed immersion. ∎
Proposition 7.10.
Suppose $f^{\dagger}:X^{\dagger}\longrightarrow Y^{\dagger}$ is a log-l.c.i. morphism (Definition 7.8) of fs log-schemes over $S^{\dagger}$, whose underlying morphism $f:X\to Y$ is a dominant and generically smooth morphism of integral schemes. Assume that the log-structures of $X^{\dagger}$ and $Y^{\dagger}$ are generically trivial. Then the cotangent sequence is left-exact:
(7.11) $0\longrightarrow f^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\longrightarrow\Omega_{X^{\dagger}/S^{\dagger}}\longrightarrow\Omega_{X^{\dagger}/Y^{\dagger}}\longrightarrow 0$
Proof.
The statement is étale-local with respect to $X$, so we may assume that $f^{\dagger}$ factors as a strict regular closed immersion $i^{\dagger}:X^{\dagger}\to W^{\dagger}$ followed by a log-smooth morphism $h^{\dagger}:W^{\dagger}\longrightarrow Y^{\dagger}$. We claim that the morphism
$\delta:\mathscr{C}_{X/W}\longrightarrow i^{*}\Omega_{W^{\dagger}/Y^{\dagger}}$
as in (7.6) is injective. To verify the claim, first note that $\mathscr{C}_{X/W}$ is locally free of finite rank since $i$ is a regular closed immersion. As $h^{\dagger}$ is log-smooth, $f^{*}\Omega_{Y^{\dagger}/Z^{\dagger}}$ is locally free of finite rank too. It will therefore suffice to show that $\delta$ is injective at the generic point of $X$. But generically, the log-structures are trivial, and upon choosing bases of $\mathscr{C}_{X/W}$ and $i^{*}\Omega_{W/Y}$ one sees that $\delta$ is given by a Jacobian matrix of $f$, which by the assumption on generic smoothness has maximal rank, and this settles the claim.
Note that since $h^{\dagger}$ is log-smooth the morphism $h^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\to\Omega_{W^{\dagger}/S^{\dagger}}$ is split injective, implying the same is true for $f^{*}\Omega_{Y^{\dagger}/S^{\dagger}}\to i^{*}\Omega_{W^{\dagger}/S^{\dagger}}$. The conclusion now follows from a diagram chase in the following commuting diagram of exact complexes.
00{0}${\mathscr{C}_{X/W}}$${i^{*}\Omega_{W^{\dagger}/Y^{\dagger}}}$${i^{*}\Omega_{W^{\dagger}/S^{\dagger}}}$${f^{*}\Omega_{Y^{\dagger}/S^{\dagger}}}$${\Omega_{X^{\dagger}/S^{\dagger}}}$00{0}00{0}
Remark 7.12.
In the case where $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is an integral morphism, we can also interprete the proof of Proposition 7.10 via the log-cotangent complex: By [Ols05, 6.9] Olsson’s logarithmic cotangent complex $\mathbb{L}_{X^{\dagger}/Y^{\dagger}}$ associated to $f^{\dagger}$ is concentrated in degrees $[-1,0]$, and is quasi-isomorphic to
[…]
Since $\delta$ is injective the homology in degree $-1$ is trivial. Since $f:X^{\dagger}\to Y^{\dagger}$ is an integral morphism, we have by [Ols05, 1.1.6] the distinguished transitivity triangle
$f^{*}\mathbb{L}_{Y^{\dagger}/S^{\dagger}}\to\mathbb{L}_{X^{\dagger}/Y^{\dagger}}\to\mathbb{L}_{X^{\dagger}/S^{\dagger}}\overset{+1}{\to},$
and the result follows by examination of the associated long exact sequence. Perhaps there is also an approach using Gabber’s cotangent complex [Ols05, §8], but we do not pursue the details.
7.13Residue exact sequence..
Let $D$ be a prime Cartier divisor on an integral scheme $X$ and consider the log-scheme $X^{\dagger}=X(\mathrm{log}D)$. Abbreviate $\Omega_{X}=\Omega_{X/\mathbb{Z}}$ and $\Omega_{X^{\dagger}}=\Omega_{X^{\dagger}/\mathbb{Z}^{\dagger}}$, where $\mathbb{Z}$ is given the trivial log-structure, and denote by $j:D\to X$ the closed immersion. Then we claim that there is an exact sequence
(7.14) $0\longrightarrow\Omega_{X}\longrightarrow\Omega_{X^{\dagger}}\overset{\mathrm{res}}{\longrightarrow}j_{*}\mathcal{O}_{D}\longrightarrow 0,$
The first map in (7.14) is the canonical map, to see it is injective we argue as follows. Suppose $g\in\mathscr{O}_{X,x}$ is a local equation for $D$ near $x\in X$. Then we have the following pushout diagram.
${\mathscr{O}_{X,x}}$${\mathscr{O}_{X,x}}$${\Omega_{X,x}}$${\Omega_{X^{\dagger},x}}$$\scriptstyle{\cdot g}$$\scriptstyle{\cdot\mathrm{d}g}$$\scriptstyle{\cdot\mathrm{dlog}g}$
Note that the upper horizontal arrow is injective, and therefore by abstract properties of pushouts the lower horizontal arrow is so too.
The second map of (7.14), also called the Poincaré residue map [Ogu18, IV.1.2.14], is induced by the map
$\Omega_{X,x}\oplus\mathscr{O}_{X,x}(\mathrm{dlog}g)\longrightarrow(j_{*}\mathscr{O}_{D})_{x}=\mathscr{O}_{X,x}/(g):(\omega,h\cdot\mathrm{dlog}(g))\mapsto h\bmod{g},$
which clearly factors through the relation $\mathrm{d}g=g\,\mathrm{dlog}(g)$. Note that by construction the residue map is surjective and that $\mathrm{res}([(\omega,h\mathrm{dlog}(g))])=0$ only if $h=gk$ for some $k\in\mathscr{O}_{X,x}$, and in that case the log-form $[(\omega,h\mathrm{dlog}(g))]=[(\omega+k\mathrm{d}g,0)]$ comes from $\Omega_{X,x}$. This settles the claim made above: (7.14) is an exact sequence.
7.15Reminders on determinants..
Let $X$ be a scheme. Recall that an $\mathcal{O}_{X}$-module $\mathscr{F}$ is called perfect if $\mathscr{F}$ has finite projective dimension, or in other words, $\mathscr{F}$ locally admits a finite length projective resolution. If $X$ is a regular scheme then all locally finitely generated $\mathcal{O}_{X}$-modules are perfect [Sta, 066Z].
Knudsen and Mumford [KM76, Theorem 2] have shown the existence of a determinant functor, unique up to unique isomorphism, which is a pair $(\det,i)$ consisting of a functor
$\det:\mathbf{Perf}\text{-}\mathbf{Mod}_{\mathcal{O}_{X}}\text{-}\mathbf{Iso}\longrightarrow\mathbf{Inv}\text{-}\mathbf{Mod}_{\mathcal{O}_{X}}\text{-}\mathbf{Iso}$
from the category of perfect $\mathcal{O}_{X}$-modules with isomorphisms to the category of invertible $\mathcal{O}_{X}$-modules with isomorphisms; together with “adjunction” isomorphisms
$i(\alpha,\beta):\det\mathscr{F}^{\prime}\otimes\det\mathscr{F}^{\prime\prime}\overset{\sim}{\longrightarrow}\det\mathscr{F}$
for every short exact sequence of perfect $\mathcal{O}_{X}$-modules
[…]
satisfying the following properties (a)-(g).
  1. (a)
    Both $\det$ and $i$ commute with flat base change;
  2. (b)
    We have $\det 0=\mathcal{O}_{X}$;
  3. (c)
    If $\mathscr{F}$ is a finite rank locally free $\mathcal{O}_{X}$-module then $\det\mathscr{F}$ is the top wedge power of $\mathscr{F}$.
  4. (d)
    The adjunctions $i(0,\mathrm{Id}_{\mathscr{F}})$ and $i(\mathrm{Id}_{\mathscr{F}},0)$ are the obvious ones.
  5. (e)
    If $\mathscr{F},\mathscr{F}^{\prime},\mathscr{F}^{\prime\prime}$ are all locally free of finite ranks $r,r^{\prime},r^{\prime\prime}$ respectively, then $i(\alpha,\beta)$ is the usual adjunction map determined by the rule
    $(s_{1}^{\prime}\wedge\cdots\wedge s^{\prime}_{r^{\prime}})\otimes(\beta(s_{1})\wedge\cdots\wedge\beta(s_{r^{\prime\prime}}))\mapsto\alpha(s_{1}^{\prime})\wedge\cdots\alpha(m^{\prime}_{r^{\prime}})\wedge s_{1}\wedge\cdots s_{r^{\prime\prime}}$
    for any choice of sections $s^{\prime}_{1},\dots,s^{\prime}_{r^{\prime}}\in\mathscr{F}^{\prime}$ and $s_{1},\dots,s_{r^{\prime\prime}}\in\mathscr{F}$.
  6. (f)
    The adjunctions $i(\alpha,\beta)$ are “functorial in isomorphisms of short exact sequences”.
  7. (g)
    The adjunctions $i(\alpha,\beta)$ are “functorial in short exact sequences of short exact sequences”.
In general, $\det\mathscr{F}$ is constructed as an alternating product of determinants of the terms in a projective resolution, and carefully bookkeeping isomorphisms, see [KM76] for details.
Definition 7.16 (log-canonical modules).
Let $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ be a morphism of log-schemes. In case the following conditions are met:
  1. (a)
    the log-structures of $X^{\dagger}$ and $Y^{\dagger}$ are coherent;
  2. (b)
    the morphism $f$ is locally of finite type;
  3. (c)
    the $\mathscr{O}_{X}$-module $\Omega_{X^{\dagger}/Y^{\dagger}}$ is perfect;
then we say that the log-canonical module of $f^{\dagger}$ is defined, and it is given by
$\omega_{X^{\dagger}/Y^{\dagger}}\coloneqq\det\Omega_{X^{\dagger}/Y^{\dagger}}.$
Let us note that later in 7.27 we will consider log-canonical modules in a more general setting.
7.17.
As a special case, when the log-structures are trivial, we recover the definition of canonical modules: let $f:X\to Y$ be a locally of finite type morphism of schemes such that $\Omega_{X/Y}$ is a perfect $\mathscr{O}_{X}$-module. Then the canonical module $\omega_{X/Y}=\det\Omega_{X/Y}$ is defined.
7.18.
If $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is a locally of finite type morphism of coherent log-schemes, then $\Omega_{X^{\dagger}/Y^{\dagger}}$ is a perfect $\mathscr{O}_{X}$-module if for instance $X$ is a regular scheme (by 7.15) or if $X^{\dagger}$ and $Y^{\dagger}$ are both log-regular, as we explain now. If $X^{\dagger}$ and $Y^{\dagger}$ are log-regular, then by Lemma 7.9 this implies that $f^{\dagger}$ is log-l.c.i. So by the claim in the proof of Proposition 7.10 the sequence (7.6) is exact and this yields a two-term locally free resolution of $\Omega_{X^{\dagger}/Y^{\dagger}}$.
7.19.
Formation of $\omega_{X^{\dagger}/Y^{\dagger}}$ is compatible with flat base change in the following sense. If $\omega_{X^{\dagger}/Y^{\dagger}}$ is defined and $Y^{\prime\dagger}\to Y^{\dagger}$ is a morphism of log-schemes whose underlying scheme morphism is flat, then $\omega_{X^{\prime\dagger}/Y^{\dagger}}$ is defined and canonically isomorphic to $(X^{\prime\dagger}\to X^{\dagger})^{*}\omega_{X^{\dagger}/Y^{\dagger}}$ where $X^{\prime\dagger}=X^{\dagger}\times_{Y^{\dagger}}Y^{\prime\dagger}$. The previous fact remains true in the categories of fine and fs log-schemes (provided we switch to fine or fs base change).
7.20.
Let $X$ be an integral scheme and let $i:\eta\to X$ denote the inclusion of the generic point of $X$. Let $\mathscr{F}$ be a coherent sheaf on $X$. Then $i^{*}\mathscr{F}$ is a $K(X)$-vector space of dimension $\mathrm{rank}(\mathscr{F})$, and we write $\mathscr{F}_{\eta}\coloneqq i_{*}i^{*}\mathscr{F}$ for the associated constant sheaf on $X$. In the case where $\mathscr{F}$ is torsion-free, we identify $\mathscr{F}$ as a sub-$\mathscr{O}_{X}$-module of $\mathscr{F}_{\eta}$ via the natural map $\mathscr{F}\to i_{*}i^{*}\mathscr{F}$.
In particular, for any perfect $\mathscr{O}_{X}$-module $\mathscr{F}$ we view $\det\mathscr{F}$ as a sub-$\mathscr{O}_{X}$-module of $\wedge^{\mathrm{rk}(\mathscr{F})}\mathscr{F}_{\eta}$. Any adjunction $i(\alpha,\beta)$ as in 7.15 is compatible with flat base change (property (a) of determinants), and so is generically compatible with $i(\alpha_{\eta},\beta_{\eta})$.
Lemma 7.21.
Let $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ be a morphism of log-schemes with generically trivial log-structures such that $\omega_{X^{\dagger}/Y^{\dagger}}$ is defined. Additionally assume that $f$ is a dominant generically étale morphism of integral schemes. Then $\omega_{X^{\dagger}/Y^{\dagger},\eta}=\mathscr{O}_{X,\eta}$ and $\mathscr{O}_{X}\subset\omega_{X^{\dagger}/Y^{\dagger}}\subset\mathscr{O}_{X,\eta}$. So, $\omega_{X^{\dagger}/Y^{\dagger}}^{-1}$ is an ideal of $\mathscr{O}_{X}$.
Proof.
The assumptions imply $\Omega_{X^{\dagger}/Y^{\dagger},\eta}=\Omega_{X/Y,\eta}=0$ so that $\omega_{X^{\dagger}/Y^{\dagger},\eta}=\mathscr{O}_{X,\eta}$. Choose a surjective map $\mathscr{O}_{X}^{n}\longrightarrow\Omega_{X^{\dagger}/Y^{\dagger}}$ for some $n\in\mathbb{N}$ and let $\mathscr{G}$ denote the kernel. Then $\mathscr{G}$ is free of rank $n$ and we have a sequence
$0\longrightarrow\mathscr{G}\longrightarrow\mathscr{O}_{X}^{n}\longrightarrow\Omega_{X^{\dagger}/Y^{\dagger}}\longrightarrow 0$
By 7.15 (a) and (d) the associated adjunction isomorphism can then be viewed as an equality of submodules of $\mathscr{O}_{X,\eta}$ given by $\omega_{X^{\dagger}/Y^{\dagger}}=\left(\det\mathscr{G}\right)^{-1}$, and by 7.15 (c) we have $\det\mathscr{G}\subset\mathscr{O}_{X}$ as submodules of $\mathscr{O}_{X,\eta}$.∎
Proposition 7.22 (log-adjunction).
Suppose $X^{\dagger}\overset{f^{\dagger}}{\longrightarrow}Y^{\dagger}\longrightarrow S^{\dagger}$ are morphisms locally of finite type of log-regular log-schemes with generically trivial log-structures. Further assume that the underlying morphism $f:X\to Y$ is a flat, dominant and generically smooth morphism of integral schemes. Then as $\mathscr{O}_{X}$-submodules of $\omega_{X/Y,\eta}$ we have the following equality, called log-adjunction:
(7.23) $\omega_{X^{\dagger}/Y^{\dagger}}\otimes f^{*}\omega_{Y^{\dagger}/S^{\dagger}}=\omega_{X^{\dagger}/S^{\dagger}}.$
Proof.
This follows from Lemma 7.9, Proposition 7.10 and the existence of the adjunction morphisms as in 7.15. ∎
7.24.
Specialising Proposition 7.22 to the setting of trivial log-structures, we obtain the following. Let $f:X\to Y$ be a flat, dominant and generically smooth morphism of regular integral locally finite type $S$-schemes. Then as $\mathscr{O}_{X}$-submodules of $\omega_{X/Y,\eta}$ we have the following equality, called adjunction:
(7.25) $\omega_{X/Y}\otimes f^{*}\omega_{Y/S}=\omega_{X/S},$
7.26Assumptions.
For later reference consider the following assumptions. Let $f:X\to Y$ be a flat and generically smooth morphism of normal integral schemes, and $D_{X}$ and $D_{Y}$ are reduced effective divisors on $X$ and $Y$ respectively such that $D_{X}\subset f^{*}D_{Y}$ and suppose that $f|_{D_{X}}:D_{X}\to D_{Y}$ maps generic points to generic points.
Denote by $f^{\dagger}:X^{\dagger}=X(\mathrm{log}\,D_{X})\longrightarrow Y^{\dagger}=Y(\mathrm{log}\,D_{Y})$ the induced morphism of divisorial log-schemes.
7.27Construction.
Assumptions as in 7.26. In this section we construct the canonical sheaf $\omega_{X/Y}$ of $f$ and the log-canonical sheaf $\omega_{X^{\dagger}/Y^{\dagger}}$ of $f^{\dagger}$, in a fashion compatible with Definition 7.16. Since $Y$ is assumed normal there exists a regular open subscheme $V\subset Y$ such that $Y\setminus V$ has codimension at least $2$. We may suppose that $D_{V}\coloneqq(D_{Y})|_{V}$ is a normal crossings divisor on $V$, because we can remove the singular locus of $D_{V}$ from $V$ and $\mathrm{codim}(Y\setminus V)\geq 2$ will still hold.
Similarly, we can find a regular open subscheme $U\subset f^{-1}(V)$ so that $\mathrm{codim}(f^{-1}(V)\setminus U)\geq 2$ and $D_{U}\coloneqq(D_{X})|_{U}$ is normal crossings on $U$. By our assumption on the generic points of $D_{X}$ and $D_{Y}$, it follows that $\mathrm{codim}(X\setminus U)\geq 2$.
Now denote by $\iota:U\to X$ the open immersion and $U^{\dagger}\coloneqq U(\mathrm{log}(D_{U}))\to V^{\dagger}\coloneqq V(\mathrm{log}(D_{V}))$ for the induced morphism of log-schemes. Then $\omega_{U/V}$ and $\omega_{U^{\dagger}/V^{\dagger}}$ are well-defined and we define
$\omega_{X/Y}\coloneqq\iota^{*}\omega_{U/V}\text{ and }\omega_{X^{\dagger}/Y^{\dagger}}\coloneqq\iota^{*}\omega_{U^{\dagger}/V^{\dagger}}.$
To see that this construction is independent of the choice of $V$ and $U$: it follows from [EHN15, 3.3.1] that the sheaf $\omega_{X/Y}$ (resp. the sheaf $\omega_{X^{\dagger}/Y^{\dagger}}$) is characterised as the unique reflexive sheaf whose restriction to $U$ coincides with $\omega_{U/V}$ (resp. $\omega_{U^{\dagger}/V^{\dagger}}$). This characterisation moreover implies that the construction is compatible with the earlier definitions of (log-)canonical sheaves, if these are applicable. Beware that a priori $\omega_{X/Y}$ and $\omega_{X^{\dagger}/Y^{\dagger}}$ are only defined as reflexive rank one sheaves, and we do not study the question in what situations these are invertible sheaves.
Theorem 7.28 (Comparison of canonical and log-canonical modules).
Assumptions as 7.26. Then as $\mathcal{O}_{X}$-submodules of $\omega_{X/Y,\eta}$, we have
(7.29) $\omega_{X^{\dagger}/Y^{\dagger}}=\omega_{X/S}(D_{X}-f^{*}D_{Y}).$
Proof.
By Construction 7.27 it suffices to prove the equality in the case were $X$ and $Y$ are regular and $X^{\dagger}$ and $Y^{\dagger}$ are log-regular because the sheaves $\omega_{X/Y}$ and $\omega_{X^{\dagger}/Y^{\dagger}}$ are determined as the unique reflexive extensions of the sheaves $\omega_{U/V}$ and $\omega_{U^{\dagger}/V^{\dagger}}$, with $U,V$ as in 7.27.
By Proposition 7.10 we have a commutative diagram
00{0}${f^{*}\Omega_{Y}}$${\Omega_{X}}$${\Omega_{X/Y}}$00{0}00{0}${f^{*}\Omega_{Y^{\dagger}}}$${\Omega_{X^{\dagger}}}$${\Omega_{X^{\dagger}/Y^{\dagger}}}$00{0}$\scriptstyle{\lambda_{X/Y}}$
with exact rows. By the residue exact sequences (7.14) for $X^{\dagger}$ and $Y^{\dagger}$, the long exact sequence provided by the snake lemma has the form
$0\longrightarrow\mathrm{ker}\lambda_{X/Y}\longrightarrow f^{*}(j_{Y})_{*}\mathscr{O}_{D_{Y}}\longrightarrow(j_{X})_{*}\mathscr{O}_{D_{X}}\longrightarrow\mathrm{coker}\lambda_{X/Y}\longrightarrow 0,$
where $j_{X}:D_{X}\to X$ and $j_{Y}:D_{Y}\to Y$ denote the closed embeddings of $D_{X}$ and $D_{Y}$ in $X$ and $Y$ respectively. All terms in the previous exact sequence are finitely generated $\mathscr{O}_{X}$-modules, therefore these are perfect as $X$ is assumed regular. The adjunction isomorphisms then imply the equalities
$\det(j_{X})_{*}\mathscr{O}_{D_{X}}\otimes\left(f^{*}\det(j_{Y})_{*}\mathscr{O}_{D_{Y}}\right)^{-1}$ $=\det\ker{\lambda_{X/Y}}\otimes\left(\det\mathrm{coker}{\lambda_{X/Y}}\right)^{-1}$
$=\omega_{X^{\dagger}/Y^{\dagger}}\otimes\omega_{X/Y}^{-1}$
as submodules of $\omega_{X/Y,\eta}$.
To finish the proof we need to check that if $j:D\to X$ is the closed embedding of a reduced effective divisor on a regular scheme $X$, then $\det j_{*}\mathscr{O}_{D}=\mathscr{O}_{X}(D)$. This follows from the adjunction isomorphism associated to the short exact sequence
$0\longrightarrow\mathscr{O}_{X}(-D)\longrightarrow\mathscr{O}_{X}\longrightarrow j_{*}\mathscr{O}_{D}\longrightarrow 0$
Lemma 7.30.
Let $X^{\dagger}$ be a log-regular scheme over $S$ (equipped with the trivial log-structure) and let $E_{x}=\overline{\{x\}}$ be the log-stratum associated to a point $x\in F(X^{\dagger})$. Let $D_{x}$ be as in 2.9. Then
$\left(\omega_{X/S}\otimes\mathscr{O}_{X}(D_{X})\right)|_{E_{x}}\cong\omega_{E_{x}/S}\otimes\mathscr{O}_{E_{x}}(D_{x}).$
Proof.
By Lemma 2.9, $E_{x}^{\dagger}=E_{x}(\mathrm{log}D_{x})$ is log-regular. By induction we can assume $x$ has height $1$. Then $E_{x}$ is a regular codimension $1$ subscheme in $X$. After possibly removing a codimension $2$ subscheme of $X$, we may assume that $X$ is regular (see Construction 7.27) and hence $E_{x}\to X$ is l.c.i., it then follows as in the proof of 7.10 that the (non-logarithmic) cotangent sequence 7.6 is left-exact:
(7.31) $0\longrightarrow\mathscr{C}_{E_{x}/X}{\longrightarrow}\Omega_{X/S}|_{E_{x}}\longrightarrow\Omega_{E_{x}/S}\longrightarrow 0,$
By taking determinants and using [Liu02, §9.1.36] we obtain the desired adjunction formula:
$\omega_{{X}/S}|_{E_{x}}=\omega_{E_{x}/S}\otimes\mathscr{C}_{E_{x}/X}\cong\omega_{E_{x}/S}\otimes\mathscr{O}(E_{x}|_{E_{x}})=\omega_{E_{x}/S}\otimes\mathscr{O}_{E_{x}}(D_{X}|_{E_{x}}-D_{x}).$
Corollary 7.32.
Suppose $X^{\dagger}$ is a generically smooth toroidal scheme over $S^{\dagger}$ and $x\in F(X^{\dagger})$.
  1. (i)
    If $x$ is vertical, then $\omega_{X^{\dagger}/S^{\dagger}}|_{E_{x}}\cong\omega_{E_{x}/\tilde{k}}\otimes\mathscr{O}_{E_{x}}(D_{x})$. If $\tilde{k}$ is perfect this is $\omega_{E_{x}^{\dagger}/\tilde{k}}$.
  2. (ii)
    If $x$ is horizontal, then $E_{x}^{\dagger}$ is a toroidal scheme over $S^{\dagger}$ and $\omega_{X^{\dagger}/S^{\dagger}}|_{E_{x}}\cong\omega_{E_{x}/S}$. If $E_{x}$ is generically smooth (eg. $k$ is perfect) then this is $\omega_{E_{x}^{\dagger}/S^{\dagger}}$.
Proof.
By Theorem 7.28 we have the comparison formula $\omega_{{X}^{\dagger}/S^{\dagger}}\cong\omega_{{X}/S}\otimes\mathscr{O}_{{X}}(D_{X}-\pi^{*}D_{S})$ where $\pi:X\to S$ is the structure map. If $x$ is vertical then $\pi^{*}D_{S}|_{E_{x}}=0$, and if $x$ is horizontal then $\pi^{*}D_{S}|_{E_{x}}=(E_{x}\to S)^{*}D_{S}$. The conclusions then follow from Lemma 7.30 and 7.28. ∎
7.33Trace section and relative canonical divisor.
Suppose $f^{\dagger}:X^{\dagger}\to Y^{\dagger}$ is a morphism of log-schemes such that $\omega_{X^{\dagger}/Y^{\dagger}}$ is defined, and additionally assume that $f$ is of relative dimension zero. Then by Lemma 7.21 the sheaf $\omega_{X^{\dagger}/Y^{\dagger}}$, viewed as an $\mathscr{O}_{X}$-submodule of $\omega_{X/Y,\eta}$, contains $\mathscr{O}_{X}$. We call the image of the unit section $1\in\Gamma(X,\mathscr{O}_{X})\to\Gamma(X,\omega_{X^{\dagger}/Y^{\dagger}})$ the trace section $\tau_{X/Y}$. By construction, trace sections are compatible with localisation.
The relative canonical divisor is defined as
$K_{X^{\dagger}/Y^{\dagger}}\coloneqq\mathrm{div}\,(\tau_{X^{\dagger}/Y^{\dagger}}).$
Note that $K_{X^{\dagger}/Y^{\dagger}}$ is an effective Cartier divisor.
8. Riemann-Hurwitz for skeleta
8.1.
In this section we apply the Poincaré-Lelong slope formula (Theorem 6.20) and our study of the relative log-canonical divisor of Section 7 to recover a Riemann-Hurwitz formula for covers of skeleta. We also study the potential theory of differential forms on skeleta.
8.2Notation.
In this section, assume $k$ is perfect and let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a finite cover of proper generically smooth toroidal $S$-schemes and suppose $f$ is étale above ${X}\setminus D_{{X}}$. Write $\phi:\Sigma^{\prime}\to\Sigma$ for the associated cover of dual complexes. We also view $\phi$ as a simultaneous skeleton of $f^{\mathrm{an}}:(X^{\prime}_{k})^{\mathrm{an}}\to({X}_{k})^{\mathrm{an}}$.
Definition 8.3 (Tropical canonical divisors).
For a vertical ridge $\tau$ of $\Sigma$, we write
$\chi_{\Sigma}(\tau)\coloneqq 2-2g(E_{\tau})-\mathrm{val}_{\Sigma}(\tau)=\deg\omega_{E_{\tau}^{\dagger}/\tilde{k}},$
and we define the tropical canonical divisor as
$K_{\Sigma}\coloneqq\sum_{\tau\text{ vertical}}-\frac{\chi_{\Sigma}(\tau)}{\mathrm{vol}(\tau)}[\tau].$
We define the tropical relative canonical divisor as
$K_{\Sigma^{\prime}/\Sigma}=\sum_{\tau^{\prime}\text{ vertical}}-\frac{\chi_{\Sigma^{\prime}/\Sigma}(\tau^{\prime})}{\mathrm{vol}(\tau^{\prime})}[\tau^{\prime}],$
where for brevity we write $\chi_{\Sigma/\Sigma^{\prime}}(\tau^{\prime})=\chi_{\Sigma^{\prime}}(\tau)-[E_{\tau^{\prime}}:E_{\tau}]\chi_{\Sigma}(\tau)$.
8.4The different.
The relative canonical divisor $K_{X^{\prime\dagger}/{X}^{\dagger}}$ is effective Cartier (7.33) and supported on $D_{{X}}$ (by the hypothesis that $f$ is étale above ${X}\setminus D_{{X}}$), and so $K_{X^{\prime\dagger}/{X}^{\dagger}}$ is determined (see 5.4) by a unique non-negative $\mathbb{Z}$-PL function
(8.5) $\delta_{X^{\prime\dagger}/{X}^{\dagger}}:\mathcal{D}(X^{\prime\dagger})\to\mathbb{R}_{\geq 0}$
that we call the different associated to $f:{X^{\prime\dagger}\to{X}^{\dagger}}$.
Lemma 8.6.
For any $x^{\prime}\in\mathrm{sk}(X^{\prime\dagger})^{\mathrm{div}\,}$ we have $\delta_{X^{\prime\dagger}/{X}^{\dagger}}(x^{\prime})=\frac{1}{e(\mathscr{H}(x^{\prime})/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}\,\Omega^{\mathrm{log}}_{\mathscr{H}(x^{\prime})^{\circ}/\mathscr{H}(f(x^{\prime}))^{\circ}}$.
Proof.
If $x^{\prime}$ is the divisorial point associated to a component of $X^{\prime}_{s}$, then this follows from the fact that the trace section of 7.33 is compatible with localisation and the fact that the different ideal $\mathrm{Fitt}^{0}\Omega^{\mathrm{log}}_{\mathscr{H}(x)^{\circ}/\mathscr{H}(f(x))^{\circ}}$ is equal to $(\omega_{\mathscr{H}(x)^{\circ}/\mathscr{H}(f(x))^{\circ}}^{\mathrm{log}})^{-1}$.
In general, we reduce to the previous setting using toroidal subdivisions as follows. The divisorial points of $\mathrm{sk}({X}^{\dagger})$ coincide with the $\mathbb{Q}$-PL points of $\mathrm{sk}({X}^{\dagger})$, see for instance [MN15, §3.2]. So there exists a proper barycentric subdivision ${X}_{0}\to{X}$ of such that $f(x^{\prime})$ is the divisorial point associated to the exceptional component. By Lemma 2.16 the pullback subdivision $X^{\prime}_{0}\to X^{\prime}$ such that $x^{\prime}$ is the divisorial point associated to a component of $(X^{\prime}_{0})_{s}$. Since $g:X^{\prime}_{0}\to X^{\prime}$ is log-étale we have $g^{*}\omega_{X^{\prime\dagger}/{X}^{\dagger}}=\omega_{(X^{\prime}_{0})^{\dagger}/({X}_{0})^{\dagger}}$, which implies $\delta_{X^{\prime\dagger}/{X}^{\dagger}}(x^{\prime})=\delta_{(X^{\prime}_{0})^{\dagger}/({X}_{0})^{\dagger}}(x^{\prime})$. ∎
Theorem 8.7.
There exists a $\mathbb{Z}$-PL function $\delta:\mathrm{sk}(X^{\prime\dagger})\to\mathbb{R}_{\geq 0}$, called the different, such that for all $x^{\prime}\in\mathrm{sk}(X^{\prime\dagger})^{\mathrm{div}\,}$ we have
$\delta(x^{\prime})=\frac{1}{e(\mathscr{H}(x^{\prime})/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}\,\Omega^{\mathrm{log}}_{\mathscr{H}(x^{\prime})^{\circ}/\mathscr{H}(f(x^{\prime}))^{\circ}}.$
Moreover if $X^{\prime\dagger},{X}^{\dagger}$ are $\mathbb{Q}$-factorial we have the Riemann-Hurwitz formula
$\Delta(\delta)=K_{\Sigma^{\prime}/\Sigma}.$
Proof.
The first statement follows from Lemma 8.6. For the second statement, note that by corollary 7.32 we have $\mathrm{sp}_{*}\omega_{{X}^{\dagger}/S^{\dagger}}=K_{\Sigma}$. By the adjunction formula 7.22 and 5.24 it then follows that $\mathrm{sp}_{*}K_{X^{\prime\dagger}/{X}^{\dagger}}=K_{\Sigma^{\prime}}-\phi^{*}K_{\Sigma}$. A computation, using 4.7 and the pullback of cycles defined in 5.20, shows that the tropical adjunction formula holds:
$K_{\Sigma^{\prime}/\Sigma}=K_{\Sigma^{\prime}}-\phi^{*}K_{\Sigma}.$
The conclusion now follows from the Poincaré-Lelong formula (Theorem 6.20):
$\Delta(\delta)=\mathrm{sp}_{*}K_{X^{\prime\dagger}/{X}^{\dagger}}=K_{\Sigma^{\prime}/\Sigma}.$
Remark 8.8 (Analytic different).
In a series of works Temkin and collaborators [CTT16, Tem16, Tem17, BT20] have defined an analytic different function
$\delta:(X^{\prime})_{k}^{\mathrm{an}}\to\mathbb{R}_{\geq 0}:x^{\prime}\mapsto\mathrm{log}_{|\varpi|}\mathrm{cont}\,\Omega^{\mathrm{log}}_{\mathscr{H}(x^{\prime})^{\circ}/\mathscr{H}(x)^{\circ}},$
where for any $k^{\circ}$-module $Q$ we write $\mathrm{cont}\,Q=\inf_{\alpha}\mathrm{cont}\,Q_{\alpha}$ for the exponential length or content [Tem16, §3], defined as infimum of contents of finitely presented subquotients. If $x^{\prime}\in X^{\prime\mathrm{an}}$ is divisorial then $\delta_{f}(x^{\prime})=\frac{1}{e(\mathscr{H}(x)/k)}\mathrm{length}_{\mathscr{H}(x)^{\circ}}\,\Omega^{\mathrm{log}}_{\mathscr{H}(x)^{\circ}/\mathscr{H}(f(x))^{\circ}}$. Note that we apply a logarithm to comply with our earlier notation, whereas Temkin e.a. usually work with multiplicative notation which is better suited for the non-discretely valued setting.
In the upcoming work of Hübner-Temkin [HT25] it will be shown that over any nonarchimedean field $k$ (not necessarily discretely valued or algebraically closed) the analytic different function $\delta$ associated to an étale morphism of $k$-analytic spaces defines a continuous function which is additionally PL on PL-charts when $k$ is stable. If we admit continuity of $\delta$, defined via the aforementioned formula, then by Lemma 8.6 and the fact that $\mathrm{sk}({X}^{\dagger})^{\mathrm{div}\,}$ is dense in $\mathrm{sk}({X}^{\dagger})$ (see [MN15, §2.4.12]) we see that the constructions agree: $\delta|_{\mathrm{sk}({X}^{\dagger})}=\delta_{X^{\prime\dagger}/{X}^{\dagger}}$
8.9.
Along the same lines of Theorem 8.7 one can obtain the following result on the piecewise linearity of the “absolute” different $\lambda$.
Proposition 8.10.
There exists a $\mathbb{Z}$-affine function $\lambda:\mathrm{sk}({X}^{\dagger})\to\mathbb{R}_{\geq 0}$ such that for all $x\in\mathrm{sk}({X}^{\dagger})^{\mathrm{div}\,}$ we have $\lambda(x)=\frac{1}{e(\mathscr{H}(x)/k)}\mathrm{length}_{\mathscr{H}(x^{\prime})^{\circ}}\left(\Omega^{\mathrm{log}}_{\mathscr{H}(x)^{\circ}/k^{\circ}}\right)_{\mathrm{tor}}$.
Proof.
Similar as 8.6, using the line bundle $\det\left[\left(\Omega_{{X}^{\dagger}/S^{\dagger}}\right)_{\mathrm{tor}}\right]$ instead. ∎
Remark 8.11.
Proposition 8.10 this gives some evidence for an expectation of Temkin on the piecewise linearity of Kähler norms, see [Tem16, §8.3.4]. However, we do not know if $\lambda(x)=\mathrm{log}_{|\varpi|}\mathrm{cont}\,\left(\Omega_{\mathscr{H}(x)^{\circ}/k^{\circ}}^{\mathrm{log}}\right)_{\mathrm{tor}}$ for all monomial points $x$. We also do not know how to describe $\Delta(\lambda)$.
8.12Potential theory of weight functions.
We finish the paper by clarifying the relation between the different and the weight functions introduced by Mustaţă-Nicaise [MN15, §4.3].
By 7.32 we have $\mathrm{sp}_{*}\omega_{{X}^{\dagger}/S^{\dagger}}^{\otimes m}=mK_{\Sigma}$. Suppose ${X}$ is $\mathbb{Q}$-factorial, then by 6.20 it follows that if $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}^{\otimes m}/k})\setminus\{0\}$ is a nonzero $m$-canonical form with $\mathrm{div}\,\omega$ supported on $D_{X}$ then
$\Delta(F_{\mathrm{div}\,\omega})=mK_{\Sigma}.$
By definition, the support function $F_{\mathrm{div}\,\omega}$ is equal to the weight function on $\Sigma$
$\mathrm{wt}_{\omega}:\Sigma\to\mathbb{R}.$
Now equip ${X}$ with the standard log-structure and let $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}/k})\setminus\{0\}$. Then we obtain the formula
$\Delta(\mathrm{wt}_{\omega}|_{\Sigma})=mK_{\Sigma}-\rho_{*}\mathrm{div}\,\omega$
where we define
$\rho_{*}\mathrm{div}\,\omega:=\sum_{(\sigma,\tau):\ \sigma\text{ unbounded},\ \tau\text{ bounded}}\partial_{\sigma/\tau}[\tau]$
summing over facet-ridge pairs with $\sigma$ unbounded and $\tau$ bounded. This generalises the results of [BN16, §3.2] to any dimension.
Proposition 8.13 (different as discrepancy of weight functions).
Let $\omega\in\Gamma({X}_{k},\omega_{{X}_{k}^{\otimes m}/k})\setminus\{0\}$ be supported on $D_{{X}}$, then as functions on $\mathrm{sk}(X^{\prime\dagger})$ we have
$\frac{1}{m}\delta_{X^{\prime\dagger}/{X}^{\dagger}}=\mathrm{wt}_{f^{*}\omega}-\mathrm{wt}_{\omega}\circ\phi$
Proof.
By the adjunction formula 7.22 we have
$\omega_{{X}^{\dagger}/X^{\prime\dagger}}\cong\underline{\mathrm{Hom}}(f^{*}\omega_{X^{\prime\dagger}/S^{\dagger}},\omega_{{X}^{\dagger}/S^{\dagger}}),$
and via this isomorphism the trace section $\tau_{X^{\prime\dagger}/{X}^{\dagger}}$ (see 7.33) corresponds to the pull-back of $m$-canonical forms. The previous then implies the desired equality is true at all divisorial points of $X^{\prime}$ attached to components of $X^{\prime}_{s}$. By 8.12 and Theorem 8.7 both left- and right-handside define functions which are $\mathbb{Z}$-affine on $\mathrm{sk}(X^{\prime\dagger})$, therefore they coincide. ∎
Appendix A Conventions on PL geometry
A.1.
To fix notation, we collect here in this Appendix some well-known material on piecewise linear geometry. We first discuss cones and fans and their “logarithmic” description via monoids and monoidal spaces. Then we discuss $\mathbb{Z}$-PL spaces.
A.2Cones.
A convex rational polyhedral cone, henceforth simply cone, is defined as a pair $\sigma=(|\sigma|,N_{\sigma})$ consisting of a topological space $|\sigma|$ and a lattice $N_{\sigma}$ such that $|\sigma|$ is the intersection of finitely many half-spaces of the form
$\{x\in(N_{\sigma})_{\mathbb{R}}:l(x)\geq 0\}\quad\text{ for some }\quad l\in M_{\sigma}\coloneqq N_{\sigma}^{*}=\mathrm{Hom}(N_{\sigma},\mathbb{Z})$
and require that
$|\sigma|^{\mathrm{gp}}=(N_{\sigma})_{\mathbb{R}}.$
Note that if $\sigma$ is a cone, then $|\sigma|$ defines a convex sub-$\mathbb{R}_{\geq 0}$-monoid of $(N_{\sigma})_{\mathbb{R}}$ which, as an $\mathbb{R}_{\geq 0}$-monoid, is generated by finitely many elements of $N_{\sigma}$. We write $\langle\cdot,\cdot\rangle:M_{\sigma}\times N_{\sigma}\to\mathbb{Z}$ for the natural bilinear pairing; we call $M_{\sigma}$ the lattice of $\mathbb{Z}$-linear functions.
A morphisms of cones $\sigma\to\sigma^{\prime}$ is a group morphism $h:N_{\sigma}\to N_{\sigma^{\prime}}$ with $h_{\mathbb{R}}(|\sigma|)\subset|\sigma^{\prime}|$; note then $h_{\mathbb{R}}|_{|\sigma|}:|\sigma|\to|\sigma^{\prime}|$ is a morphism of $\mathbb{R}_{\geq 0}$-monoids. This defines the category of cones $\mathbf{Cone}$.
If $(\sigma,N_{\sigma})$ is a cone, then it is immediate that the monoid of integral points $\sigma\cap N_{\sigma}$ is integral, saturated and torsion-free. By Gordan’s lemma $\sigma\cap N_{\sigma}$ is also finitely generated101010proof: Write $N=N_{\sigma}$ and $\sigma=\sum_{i=1}^{k}\mathbb{R}_{\geq 0}m_{i}$ for some $m_{1},\dots,m_{k}\in\sigma$. Denote by $\|\cdot\|_{N}$ the Euclidean norm on $N_{\mathbb{R}}$ induced by a $\mathbb{Z}$-basis of $N$. Then every element in $\sigma\cap N$ is at most a distance $\epsilon=\sum_{i=1}^{k}\|m_{i}\|_{N}$ removed from an element in the submonoid $\sum_{i=1}^{k}\mathbb{N}m_{i}$ of $\sigma\cap N$. It follows that $\{m_{1},\dots,m_{k}\}$ and the collection of lattice points in the ball of radius $\epsilon$ around 00 jointly generate $M_{\sigma}$, as desired..
To simplify the notation, we will write $\sigma$ instead of $|\sigma|$ if the lattice $N_{\sigma}$ is clear from context.
A.3Associated cone.
Let $M$ be a fine torsion-free monoid. Then $M^{\mathrm{gp}}$ is a lattice and we view $M$ as an additive submonoid of $M^{\mathrm{gp}}$. The $\mathbb{R}_{\geq 0}$-submonoid of $M^{\mathrm{gp}}_{\mathbb{R}}$ generated by $M$ is denoted by $\mathbb{R}_{\geq 0}M$. We call the cone $(\mathbb{R}_{\geq 0}M,M^{\mathrm{gp}})$ the cone associated to $M$.
Note that $\mathbb{R}_{\geq 0}M=\mathbb{R}_{\geq 0}M^{\mathrm{sat}}$, where $M^{\mathrm{sat}}$ denotes the saturation of $M$. Also, $M$ is sharp if and only if $\mathbb{R}_{\geq 0}M$ is strictly convex, that is $\mathbb{R}_{\geq 0}M$ contains no lines. In general there exists a set of elements of $M$, part of a system of generators, which generate $M^{\mathrm{gp}}_{\mathbb{Q}}$ as a $\mathbb{Q}$-vector space. This implies that
$(\mathbb{R}_{\geq 0}M)\cap M^{\mathrm{gp}}=M^{\mathrm{sat}}.$
Proposition A.4 (fs monoids correspond to cones).
There exists an equivalence of categories
$\mathbf{Mon}^{\mathrm{fs}}\overset{\sim}{\longrightarrow}\mathbf{Cone}$
between the categories of fs monoids and cones, given by $M\mapsto(\mathbb{R}_{\geq 0}M,M^{\mathrm{gp}})$, with inverse $(\sigma,N_{\sigma})\mapsto\sigma\cap N_{\sigma}$. An fs monoid $M$ is sharp if and only if its associated cone is strictly convex (i.e. contains no linear subspaces).
Proof.
This is immediate from A.2 and A.3. ∎
A.5Subcones and faces.
A subcone of $\sigma=(|\sigma|,N_{\sigma})$ is a cone of the form $\sigma^{\prime}=(|\sigma^{\prime}|,N\cap\mathbb{R}|\sigma^{\prime}|)$ for some $|\sigma^{\prime}|\subset|\sigma|$. A subcone $\sigma^{\prime}\subset\sigma$ is called a face, also written
$\sigma^{\prime}\leq\sigma,$
if additionally for each $x,y\in\sigma$ we have that $x+y\in\sigma^{\prime}$ implies $x\in\sigma^{\prime}$ and $y\in\sigma^{\prime}$. The $n$-dimensional faces of $\sigma$ are denoted $\sigma(n)$. A ray is a $1$-dimensional face, i.e. an element of $\sigma(1)$. A cone is called simplicial if the integral ray generators are linearly independent.
We call a morphism of cones $h:\sigma^{\prime}\to\sigma$ a immersion if the cokernel of $h:N^{\prime}\to N$ is torsion-free, equivalently $h(\sigma^{\prime})$ defines a subcone of $\sigma$. A face embedding is an injective immersion with $h(\sigma^{\prime})\leq\sigma$.
A.6Dual monoids.
If $M$ is an abelian group, we denote by $M^{*}=\mathrm{Hom}_{\mathbb{Z}}(M,\mathbb{Z})$ the dual group. Recall that if $M$ is a lattice then the natural map $M\to M^{**}$ is an isomorphism. If $M$ is a monoid, then we denote by $M^{\vee}=\mathrm{Hom}(M,\mathbb{N})$ the dual monoid of $M$. It is immediate that $M^{\vee}$ is sharp. If $M$ is finitely generated, then $M^{\vee}$ is fs by [Ogu18, I.2.1.17(8)]. Moreover, by [Ogu18, I.2.2.3] if $M$ is fine and sharp then the natural map $(M^{\vee})^{\mathrm{gp}}\to(M^{\mathrm{gp}})^{*}$ is an isomorphism and the natural map $M\to M^{\vee\vee}$ factors through an isomorphism $M^{\mathrm{sat}}\cong M^{\vee\vee}$. So in particular for sharp fs monoids the functor $M\to M^{\vee}$ is an involution.
A.7Dual cone.
Suppose that $M$ is a fine torsion-free monoid and view $M$ as an additive submonoid of the lattice $M^{\mathrm{gp}}$. Then we denote by
$\sigma_{M}\coloneqq\mathbb{R}_{\geq 0}\left(M^{\vee}\right)$
the dual cone associated to $M$; its associated lattice is $(M^{\vee})^{\mathrm{gp}}=(M^{\mathrm{gp}})^{*}$. Suppose $M$ is sharp; by A.6 we have $(M^{\vee})^{\mathrm{gp}}_{\mathbb{R}}=\mathrm{Hom}_{\mathbb{Z}}(M^{\mathrm{gp}},\mathbb{R})=\mathrm{Hom}_{\mathbb{R}}(M^{\mathrm{gp}}_{\mathbb{R}},\mathbb{R})$ and it follows $\sigma_{M}=\mathrm{Hom}(M,\mathbb{R}_{\geq 0})=\mathrm{Hom}_{\mathbb{R}_{\geq 0}}(\mathbb{R}_{\geq 0}M,\mathbb{R}_{\geq 0}),$ that is $\sigma_{M}$ is the cone dual to the cone $\mathbb{R}_{\geq 0}M$ associated to $M$ (as $\mathbb{R}_{\geq 0}$-monoids).
Note that $\sigma_{M^{\vee}}=\mathbb{R}_{\geq 0}M^{\vee\vee}\cong\mathbb{R}_{\geq 0}M^{\mathrm{sat}}=\mathbb{R}_{\geq 0}M$ is the cone associated to $M$. An application of Gordan’s lemma shows that $M^{\mathrm{sat}}\cong\sigma_{M^{\vee}}\cap M^{\mathrm{gp}}$ is finitely generated. So, if $M$ is fine and sharp then the saturation $M^{\mathrm{sat}}$ is so too.
A.8.
For each strictly convex cone $(\sigma,N_{\sigma})$, the monoid $\sigma\cap N_{\sigma}$ of integral points is fs by A.4. We write
$\mathscr{C}_{\sigma}\coloneqq(\sigma\cap N_{\sigma})^{\vee}$
for the monoid of non-negative $\mathbb{Z}$-affine functions on $\sigma$, it is an fs monoid by A.6, and we have $\mathscr{C}_{\sigma}^{\mathrm{gp}}=M_{\sigma}$ (the lattice of $\mathbb{Z}$-affine functions) and $\mathscr{C}_{\sigma}=\{\alpha\in M_{\sigma}:\langle\alpha,\sigma\cap N_{\sigma}\rangle\subset\mathbb{N}\}$ justifying terminology.
A.9.
To sum up A.7 and A.8: for every sharp fs monoid $M$ there exists a unique cone $\sigma$ such that $M=\mathscr{C}_{\sigma}$ is the monoid of non-negative $\mathbb{Z}$-affine functions on $\sigma$, namely the dual cone $\sigma_{M}$.
Definition A.10 (Cone complex).
A convex rational polyhedral cone complex, henceforth simply cone complex is a pair $\Sigma=(|\Sigma|,\tau)$ consisting of a locally compact topological space $|\Sigma|$ equipped with a cone atlas $\tau$, which is a set of cones satisfying the following conditions.
  1. (i)
    The set $\{|\sigma|\}_{\sigma\in\tau}$ defines a covering of closed subsets of $|\Sigma|$.
  2. (ii)
    Each face of a cone in $\tau$ is again a cone in $\tau$.
  3. (iii)
    For any two cones $\sigma_{1},\sigma_{2}\in\tau$, there is a unique cone $\sigma_{1}\cap\sigma_{2}\in\tau$ such that $|\sigma_{1}\cap\sigma_{2}|=|\sigma_{1}|\cap|\sigma_{2}|$ and $\sigma_{1}\cap\sigma_{2}$ admits a face embedding in both $\sigma_{1}$ and $\sigma_{2}$.
  4. (iv)
    Each $x\in\Sigma$ admits a neighbourhood of the form $|\sigma_{1}|\cup\dots\cup|\sigma_{n}|$ for some $\sigma_{1},\dots,\sigma_{n}\in\tau$.
To ease notation we will identify a cone complex $\Sigma$ with its geometric realisation $|\Sigma|$. The lattice points of $\Sigma$ are denoted by $\Sigma(\mathbb{N})=\mathrm{colim}N_{\sigma}$. We write $M_{\Sigma}=\lim M_{\sigma}$ for the group of $\mathbb{Z}$-affine functions.
We define a morphism of cone complexes $\phi:\Sigma\to\Sigma^{\prime}$ as a map $h:\tau\to\tau^{\prime}$ together with a family $(h_{\sigma}:\sigma\to h(\sigma))_{\sigma\in\tau}$ of compatible cone morphisms, where compatibility means that whenever $\rho$ is a face of $\sigma$, the restriction $(h_{\sigma})|_{\rho}$ is the composition of $h_{\rho}$ with the face embedding of $h(\rho)$ in $h(\sigma)$. Clearly $h$ induces a continuous map $|h|:|\Sigma|\to|\Sigma^{\prime}|$, which we will often identify with $h$ for simplicity.
[…]
Remark A.11.
A definition equivalent to A.10 is given in Definition A.21, see Remark A.26.
A.12Constructions with monoids..
Recall that $\mathbf{Mon}$ admits all limits and colimits, where limits and coproducts commute with the forgetful functor to $\mathbf{Set}$ and coequalisers are described via congruence relations, i.e. equivalence relations that are submonoids [Ogu18, §1].
A.13Ideals and faces..
Let $M$ be monoid. A subset $I$ of $M$ is called an ideal of $M$ if $IM\subset I$, that is $I$ is a sub-$M$-set of $M$. By convention $I=\emptyset$ is also an ideal, and we call $I$ proper if $I\neq M$. Note that every monoid $M$ admits a maximal proper ideal, namely $M\setminus M^{\times}$. A proper ideal $I$ of $M$ is called a prime ideal if for all $m,m^{\prime}\in M$ we have $mm^{\prime}\in I$ only if $m\in I$ or $m^{\prime}\in I$. Note that if $\mathfrak{p}$ is a prime ideal then the complement $M\setminus\mathfrak{p}$ is a submonoid. The submonoids which arise in this fashion are called faces of $M$, and they are equivalently characterised as those submonoid $F\subset M$ for which $mm^{\prime}\in F$ implies $m\in F$ and $m^{\prime}\in F$. If $M$ is a fine monoid then every face of $M$ is also fine by [Ogu18, I.2.1.17(3)]. If $M$ is fine torsion-free then the map $F\mapsto\mathbb{R}_{\geq 0}F$ yields a one-to-one correspondence between faces of $M$ and faces of $\mathbb{R}_{\geq 0}M$.
Arbitrary unions of ideals (resp. prime ideals) are again ideals (resp. prime ideals), and finite intersections of ideals are again ideals. If $\phi:M\to M^{\prime}$ is a morphism of monoids then the inverse image of a prime ideal is a prime ideal.
A.14Quotients by ideals.
A quotient of a monoid $M$ by an ideal $I$ introduces an absorbing element or zero, i.e. an element 00 such that $m\cdot 0=0$ for all $m\in M$. For instance, $(\mathbb{N}\cup\{+\infty\},+)$ has zero element +$\infty$+ ∞. Any $M$-set $I$ (i.e. a set with action by $M$) can be given a 00 by constructing the coproduct $I^{0}\coloneqq I\oplus 0$ in the category of $M$-sets. If $I$ is an ideal of $M$ then $M/I$ is defined as the cokernel of $I^{0}\to M^{0}$ in the category of $M$-sets. As a set $M^{0}/I^{0}=M\setminus I\oplus 0$, and $M^{0}/I^{0}$ is canonically a monoid with zero (products that lie in $I$ are declared equal to 00). If $I$ is prime then $M/I$ is the face $M\setminus I$ with an additional zero.
Remark A.15.
The construction in A.14 suggests to work systematically with monoids with zero, this is possible after some small changes to make sure not to divide by zero (for instance one changes the notion of an integral monoid to require the cancellation law “$ab=ac$ implies $b=c$ for all nonzero $a$”). We do not pursue this any further.
A.16Spectrum of a monoid.
The spectrum $\mathrm{Spec}\,M$ of a monoid $M$ is the topological space obtained by equipping the collection of prime ideals of $M$ with the Zariski topology: the closed sets are given by $Z(I)=\{\mathfrak{p}\in\mathrm{Spec}\,M:\ \mathfrak{p}\supset I\}$ where $I$ is an ideal of $M$ – equivalently the sets $D(f)=\{\mathfrak{p}:\ f\not\in\mathfrak{p}\}$ for $f\in M$, form a basis for the topology of $\mathrm{Spec}\,M$. The spectrum yields a functor
$\mathrm{Spec}\,:\mathbf{Mon}^{\mathrm{op}}\to\mathbf{Top}.$
Let $M$ be a monoid and let $S$ be a submonoid. Then the localisation $S^{-1}M$ is the quotient of $S\times M$ by the congruence relation generated by the relations $(s,m)\sim(s^{\prime},m^{\prime})$ for all $s,s^{\prime}\in S$ and $m,m^{\prime}\in M$ such that $sm^{\prime}=sm^{\prime}$. For $f\in M$ we write $M_{f}=\{f^{n}:n\in\mathbb{N}\}^{-1}M$. We have a canonical homeomorphism $D(f)=\mathrm{Spec}\,M_{f}$. For $\mathfrak{p}\in\mathrm{Spec}\,M$ we write $M_{\mathfrak{p}}=(M\setminus\mathfrak{p})^{-1}M$, and as a subset of $\mathrm{Spec}\,M$, we have that $\mathrm{Spec}\,M_{\mathfrak{p}})$ (resp. $\mathrm{Spec}\,(M/\mathfrak{p}\setminus\{0\})=\mathrm{Spec}\,M\setminus\mathfrak{p}$ (resp. are those primes $\mathfrak{q}$ contained in $\mathfrak{p}$ (resp. containing $\mathfrak{p}$).
Let $f\in M$. Then there exists a smallest face containing $f$, namely
$\langle f\rangle\coloneqq\{m\in M:\exists m^{\prime}\in M,\exists n\in\mathbb{N}:mm^{\prime}=f^{n}\}.$
Therefore $\mathrm{Spec}\,M_{f}$ contains a unique closed point, namely $M_{f}\setminus\langle f\rangle$. The localisation $\langle f\rangle^{-1}M$ coincides with $M_{f}$.
Let $\dim M$ denote the Krull dimension of $\mathrm{Spec}\,M$, then for each $\mathfrak{p}\in\mathrm{Spec}\,M$ we have that $\dim M_{\mathfrak{p}}$ coincides with the height $h(\mathfrak{p})$ of $\mathfrak{p}$, defined as the supremum of the length of a strictly descending chain of prime ideals
$\mathfrak{p}=\mathfrak{p}_{0}\supset\mathfrak{p}_{1}...\supset\mathfrak{p}_{h}=\emptyset.$
If $M$ is finitely generated then $h(\mathfrak{p})$ is finite and $h(\mathfrak{p})=\dim M_{\mathfrak{p}}\leq\mathrm{rank}_{\mathbb{Z}}(M^{\sharp})^{\mathrm{gp}}$ with equality if $M$ is fine, by [Ogu18, I.1.4.7]. So if $M$ is fine then $\dim M=\dim(M^{\sharp})^{\mathrm{gp}}$ is finite and so is $\mathrm{Spec}\,M$.
If $M$ is fine and sharp then $\dim M=\dim\mathbb{R}_{\geq 0}M$, and so if $F$ is a face of $M$ then the height $h(\mathfrak{p})$ of $\mathfrak{p}=M\setminus F$ equals $\dim\mathbb{R}_{\geq 0}M-\dim\mathbb{R}_{\geq 0}F$.
A.17.
The category $\mathbf{MonSp}$ of monoidal spaces is defined as follows. Objects are pairs $(T,\mathscr{M}_{T})$ consisting of a topological space $T$ endowed with a sheaf of monoids $\mathscr{M}_{T}$. A morphism of monoidal spaces $(T,M_{T})\to(T^{\prime},M_{T^{\prime}})$ is a pair $(f,f^{\flat})$ consisting of a continuous map $f:T\to T^{\prime}$ and a homomorphism $f^{\flat}:f^{-1}\mathscr{M}_{T^{\prime}}\to M_{T}$ such that that induced stalk maps
$f^{\flat}_{t}:\mathscr{M}_{T^{\prime},f(t)}\to\mathscr{M}_{T,t}$
are local, that is $\mathscr{M}_{T^{\prime},f(t)}^{\times}\subset(f^{\flat}_{t})^{-1}\mathscr{M}_{T,t}^{\times}$.
A.18.
Many of the constructions in $\mathbf{Mon}$ extend to sheaves of monoids on a topological space, including the construction of limits and colimits and the functors $(\cdot)^{\mathrm{gp}},(\cdot)^{\mathrm{int}},(\cdot)^{\mathrm{sat}},(\cdot)^{\sharp}$, by applying these constructions locally and then sheafifying the outcome. All these constructions commute with passing to stalks.
For instance, if $(T,\mathscr{M}_{T})$ is a monoidal space then we write $(T,\mathscr{M}_{T})^{\sharp}=(T,\mathscr{M}_{T}^{\sharp})$ for the sharpification. We say $(T,\mathscr{M}_{T})$ is sharp if $\mathscr{M}_{T}=\mathscr{M}_{T}^{\sharp}$, equivalently all stalks $\mathscr{M}_{T,t}$ are sharp. Note that if $(f,f^{\flat}):(T,M_{T})\to(T^{\prime},M_{T^{\prime}})$ is a morphism of sharp monoidal spaces then the requirement that $f^{\flat}_{t}$ is local means that $(f^{\flat}_{t})^{-1}\{1\}=\{1\}$.
A.19.
Let $M$ be a monoid and let $T=\mathrm{Spec}\,M$. We define a sheaf of monoids $\mathscr{M}_{T}$ on $T$ via the rule $D(f)\mapsto M_{f}$ for all $f\in M$ and sheafifying the result. Since $D(f)=\mathrm{Spec}\,M_{f}$ admits a unique closed point $\mathfrak{p}=M\setminus\langle f\rangle$, it follows that $\mathscr{M}_{T}(D(f))=\mathscr{M}_{T,\mathfrak{p}}={\langle f\rangle}^{-1}M=M_{f}$, so sheafifying was in fact not needed. In particular if $f=1$ is trivial we find that $\mathscr{M}_{T}(T)=M$.
A computation similar as the one for schemes, see for instance [Ogu18, II.1.2.2] shows that the functor
$\mathbf{MonSp}\to\mathbf{Set}:(T,\mathscr{M}_{T})\mapsto\mathrm{Hom}(M,\Gamma(T,\mathscr{M}_{T}))$
is representable by the monoidal space $(\mathrm{Spec}\,M,\mathscr{M}_{\mathrm{Spec}\,M})$.
A.20Affine fans.
An affine fan is defined as a monoidal space of the form $(\mathrm{Spec}\,M,\mathscr{M}_{\mathrm{Spec}\,M})^{\sharp}$, in particular these are always sharp. We have $\mathscr{M}_{\mathrm{Spec}\,M}^{\sharp}=\mathscr{M}_{\mathrm{Spec}\,M^{\sharp}}$ and by A.19 it follows that the category of affine fans is anti-equivalent to the category of sharp monoids. In particular, affine fans admit fibered products. The affine fan $(\mathrm{Spec}\,M,\mathscr{M}_{\mathrm{Spec}\,M})^{\sharp}$ is called fs if $M$ is fs.
Definition A.21.
A fan is defined as a sharp monoidal space which is locally isomorphic to an affine fan. A fan is called locally fs (resp. fs) if it is a union (resp. a finite union) of fs affine fans. The category of fans (resp. locally fs fans) is written $\mathbf{Fan}$ (resp. $\mathbf{Fan}^{fs}$). The fibered product for affine fans glues to fibered products for fans [Ogu18, II.1.3.5], and similarly for locally fs fans.
A fan $F$ is called regular at a point $x$ if $F$ is locally fs at $x$ and
$\mathrm{rank}_{\mathbb{Z}}\mathscr{M}_{F,x}^{\mathrm{gp}}=\dim\mathscr{M}_{F,x}.$
We write $F_{\mathrm{reg}}$ for the regular points of $F$. Given a fan $F$ and a monoid $M$, we will write
$F(M)=\mathrm{Hom}(\mathrm{Spec}\,M,F).$
If $F$ is a fan we call $F(\mathbb{N})$ the set of integral points on $F$, and similarly we have the sets of rational points $F(\mathbb{Q}_{\geq 0})$ and real points $F(\mathbb{R}_{\geq 0})$.
A.22.
Let $M$ be a sharp monoid, then the set $(\mathrm{Spec}\,M)(\mathbb{N})$ of integral points of $\mathrm{Spec}\,M$ coincides with $M^{\vee}$. In particular, if $M$ is also fine then $(\mathrm{Spec}\,M)(\mathbb{N})$ is the integral points of the dual cone $\sigma_{M}$ (see A.7) and $(\mathrm{Spec}\,M)(\mathbb{R}_{\geq 0})=\sigma_{M}$.
Construction A.23 (Cone complex associated to a fan $F$).
To every locally fs fan $F$ we associate a strictly convex cone complex $\Sigma(F)$ as follows.
For every $x,y\in F$ such that $x$ is a specialisation of $y$, that is $x\in\overline{\{y\}}$, we have a local morphism $\mathscr{M}_{F,y}\longrightarrow\mathscr{M}_{F,x}$, and so we get an immersion of dual cones
(A.24) $\sigma_{\mathscr{M}_{F,x}}\longrightarrow\sigma_{\mathscr{M}_{F,y}}.$
The maps (A.24) define face embeddings and hence glue to a well-defined cone complex $\Sigma(F)=(|\sigma(F)|,\{\sigma_{\mathscr{M}_{F,x}}\}_{x\in F})$ such that
(A.25) $|\Sigma(F)|=\cup_{x\in F}\sigma_{\mathscr{M}_{F,x}}.$
In other words $\Sigma(F)=\mathrm{colim}_{x\in F}\sigma_{\mathscr{M}_{F,x}}$.
A.26.
Conversely, from every strictly convex cone complex one can reconstruct a locally fs fan and this gives an equivalence of categories – we omit the details.
Under this equivalence regular points of a locally fs fan correspond to regular cones in the corresponding cone complex, where a regular cone is defined as a simplicial cone (A.5) for which the integral points are generated by ray vectors.
The parallel definitions of subdivisions (A.10) are follows.
Definition A.27 (Subdivision).
A subdivision of locally fs fans is defined as a morphism $\phi:F^{\prime}\to F$ such that
  1. (a)
    For all $x^{\prime}\in F^{\prime}$, the map $\mathscr{M}^{\mathrm{gp}}_{F,\phi(x)}\to\mathscr{M}^{\mathrm{gp}}_{F^{\prime},x}$ surjects.
  2. (b)
    The map $F^{\prime}(\mathbb{N})\to F(\mathbb{N})$ is injective
A subdivision $\phi$ is proper if additionally $\phi$ has finite fibers and $F^{\prime}(\mathbb{N})\to F(\mathbb{N})$ is bijective.
Proposition A.28 (Existence regular subdivisions).
Let $F$ be a locally fs fan. Then there exists a proper subdivision $\phi:F^{\prime}\to F$ such that $F^{\prime}$ is regular and moreover we can require that the induced map $\phi^{-1}(F_{\mathrm{reg}})\longrightarrow F_{\mathrm{reg}}$ is an isomorphism. We will also call $\phi$ a regular subdivision.
Proof.
See either [GR18, §6.6.31] or [Wło22, §4.6.1].∎
A.29.
Let $F$ be a $2$-dimensional fs fan. Then we claim there exists a minimal regular subdivision $F^{\prime}\to F$. Indeed, if $\sigma$ is a $2$-dimensional cone then consider the convex hull $H$ of $|\sigma|\cap N_{\sigma}$. Consider the set $S$ of points in $N_{\sigma}\cap H$ which are integral ray generators. A computation shows that $S$ is the minimal generating set of $|\sigma|\cap N_{\sigma}$. The rays of these generators define a subdivision that contains any regular subdivision of $\sigma$. The set $S$ can be computed explicitly using Hirzebruch-Jung continued fractions, see for instance [Oda88, §1].
A.30.
For the remainder of this Appendix we recall the definition and some elementary properties of $\mathbb{Z}$-PL spaces. In the literature, these are also known as $\mathbb{Z}_{|k^{\times}|}-$ or $(\mathbb{Z},|k^{\times}|)-$PL spaces. Since $k$ is discretely valued, we can fix an isomorphism $|k^{\times}|\cong\mathbb{Z}$ and we will omit $|k^{\times}|$ to lighten the notation.
These spaces generalise the cone complexes defined in A.10, in fact we recover these by letting $b=0$ throughout Definitions A.31 and A.33 below.
Definition A.31 (Polyhedra).
If $N$ is a lattice a $\mathbb{Z}$-affine function on $N_{\mathbb{R}}$ as any function of the form
$\varphi:x\mapsto\langle a,x\rangle+b$
for some $a\in N^{*}$ and $b\in\mathbb{Z}$. A $\mathbb{Z}$-PL polyhedron, or short polyhedron is a pair $\sigma=(|\sigma|,N_{\sigma})$ of a topological space $|\sigma|$ and a lattice $N_{\sigma}$ such that $|\sigma|$ is the intersection of finitely many half-spaces
$\{x\in(N_{\sigma})_{\mathbb{R}}:\varphi(x)\geq 0\}\subset(N_{\sigma})_{\mathbb{R}}^{n}$
for some $\mathbb{Z}$-affine functions $\varphi$ on $\mathbb{N}_{\sigma}$ and $(N_{\sigma})_{\mathbb{R}}=\mathbb{R}|\sigma|$. The last condition ensures
$C(\sigma)\coloneqq(\overline{\mathbb{R}_{\geq 0}|\sigma|},N_{\sigma})$
is a cone with the same lattice $N_{\sigma}$ – here $\overline{\cdot}$ denotes the closure in $(N_{\sigma})_{\mathbb{R}}$. We call $C(\sigma)$ the cone associated to $\sigma$, note $N_{\sigma}=N_{C(\sigma)}$, and $M_{\sigma}\coloneqq N_{\sigma}^{*}=M_{C(\sigma)}$ is the lattice of $\mathbb{Z}$-affine functions on $\sigma$. The monoid of non-negative $\mathbb{Z}$-affine functions on $\sigma$ is $\mathscr{C}_{\sigma}\coloneqq\mathscr{C}_{C(\sigma)}$.
A subpolyhedron of $\sigma=(|\sigma|,N_{\sigma})$ is a polyhedron of the form $\sigma^{\prime}=(|\sigma^{\prime}|,N_{\sigma}\cap\mathbb{R}|\sigma^{\prime}|)$ with $|\sigma^{\prime}|\subset|\sigma|$. A face of a polyhedron $\sigma$ is a subpolyhedron obtained as an intersection of $|\sigma|$ with the boundary of a half-space $H_{l}$ containing $\sigma$, and is possibly $\emptyset$ or $\sigma$. Given a polyhedron $\sigma$ we write $\sigma(n)$ for the set of $n$-dimensional faces of $\sigma$.
A morphism of polyhedra $\sigma^{\prime}\to\sigma$ is a group morphism $h:N_{\sigma^{\prime}}\to N_{\sigma}$ such that $h_{\mathbb{R}}(|\sigma^{\prime}|)\subset|\sigma|$. We call $h$ an immersion if $\mathrm{coker}(h)$ is torsion-free, equivalently $h(\sigma^{\prime})$ defines a subpolyhedron. A face embedding is an injective immersion with $h(\sigma^{\prime})\leq\sigma$.
A polytope is a polyhedron which is bounded (equivalently, compact).
Lemma A.32.
Let $\sigma$ be a polyhedron. Then $\mathscr{C}_{\sigma}^{\mathrm{gp}}\subset M_{\sigma}$ with equality if and only if $C(\sigma)$ is strictly convex.
Proof.
It follows from the definitions $\mathscr{C}_{\sigma}^{\mathrm{gp}}\subset M_{\sigma}$. If $C(\sigma)$ is strictly convex, then we can choose a $\mathbb{Z}$-affine function $\varphi^{\prime}\in M_{\sigma}$ such that $C(\sigma)\setminus\{0\}\subset\{\varphi^{\prime}>0\}$. Let $n_{1},\dots,n_{k}$ be primitive ray generators of $C(\sigma)$. Let $\varphi\in M_{\sigma}$, after rescaling we may assume $\varphi^{\prime}(n_{i})>\max\{-\varphi(n_{i}),0\}$ for all $i=1,\dots,k$. Since the function $\max\{-\varphi,0\}$ is convex, it follows $\varphi+\varphi^{\prime}\in\mathscr{C}_{\sigma}$ and so $\varphi=(\varphi+\varphi^{\prime})-\varphi^{\prime}\in\mathscr{C}_{\sigma}^{\mathrm{gp}}$. Conversely, any element of $\mathscr{C}_{\sigma}$ vanishes on linear spaces contained in $C(\sigma)$, so the same is true for elements of $\mathscr{C}_{\sigma}^{\mathrm{gp}}$. So any element of $M_{\sigma}$ which is non-trivial on a non-trivial linear space contained in $C(\sigma)$ does not lie in $\mathscr{C}_{\sigma}^{\mathrm{gp}}$. ∎
Definition A.33 ($\mathbb{Z}$-PL space).
A $\mathbb{Z}$-PL polyhedral complex, or simply $\mathbb{Z}$-PL complex is a pair $\Sigma=(|\Sigma|,\tau)$ consisting of a locally compact topological space $|\Sigma|$ and a polyhedral atlas $\tau$, which is a set of polyhedrons such that (i)-(iv) holds:
  1. (i)
    The set $\{|\sigma|\}_{\sigma\in\tau}$ defines a covering of closed subsets of $|\Sigma|$.
  2. (ii)
    Each face of a polyhedron in $\tau$ is again a polyhedron in $\tau$.
  3. (iii)
    For any two polyhedrons $\sigma_{1},\sigma_{2}\in\tau$, there is a unique polyhedron $\sigma_{1}\cap\sigma_{2}\in\tau$ such that $|\sigma_{1}\cap\sigma_{2}|=|\sigma_{1}|\cap|\sigma_{2}|$ and $\sigma_{1}\cap\sigma_{2}$ admits a face embedding in both $\sigma_{1}$ and $\sigma_{2}$.
  4. (iv)
    Each $x\in|\Sigma|$ admits a neighbourhood of the form $|\sigma_{1}|\cup\dots\cup|\sigma_{n}|$ for some $\sigma_{1},\dots,\sigma_{n}\in\tau$.
We write
$M_{\Sigma}=\lim_{\sigma\in\Sigma}M_{\sigma}$
for the group of functions on $\Sigma$ that are $\mathbb{Z}$-affine on the faces. Similarly define $\mathscr{C}_{\Sigma}=\lim_{\sigma\in\Sigma}\mathscr{C}_{\sigma}$ for the non-negative $\mathbb{Z}$-affine functions.
[…]
Consider the equivalence relation on polyhedral complexes generated by proper subdivisions. We then define a $\mathbb{Z}$-PL space as an equivalence class of polyhedral complexes, and a morphism of $\mathbb{Z}$-PL spaces is induced from a morphism of representative polyhedral complexes. By a polyhedron of a $\mathbb{Z}$-PL space $\Sigma$ we mean any polyhedron of a representative. To ease notation, we will often identify a $\mathbb{Z}$-PL space $\Sigma=(|\Sigma|,[\tau])$ with its support $|\Sigma|$, and suppress the equivalence class of polyhedral atlasses $[\tau]$. The group of $\mathbb{Z}$-PL-functions on $\sigma$ is defined as
$\mathrm{PL}_{\Sigma}\coloneqq\mathrm{colim}_{\Sigma^{\prime}\to\Sigma}M_{\sigma^{\prime}}$
where the colimit is taken over all (not necessarily proper) subdivisions $\Sigma^{\prime}\to\Sigma$; similarly define $\mathrm{PL}_{\Sigma}^{+}$. A morphism $\Sigma^{\prime}\to\Sigma$ then induces natural homomorphisms $\mathrm{PL}_{\Sigma}\to\mathrm{PL}_{\Sigma}$ and $\mathrm{PL}_{\Sigma}^{+}\to\mathrm{PL}_{\Sigma}^{+}$. It follows from the definitions that a continuous function $\Sigma\to\mathbb{R}$ is $\mathbb{Z}$-PL (resp. $\mathbb{N}$-PL) if and only if it is $\mathbb{Z}$-PL (resp. $\mathbb{N}$-PL) on the polyhedra of some polyhedral atlas.
One may similarly define $\mathbb{Q}$-PL and $\mathbb{R}$-PL-spaces. By a PL-space we mean a $\mathbb{R}$-PL space.
A.34Associated cone complex.
To any $\mathbb{Z}$-PL space $\Sigma$ we can functorially associate a cone complex $C(\Sigma)$ as follows: The functor which sends a polyhedron $(\sigma,N_{\sigma})$ to the cone $C(\sigma)=(\overline{\mathbb{R}_{\geq 0}|\sigma|},N_{\sigma})$ assocatied to $\sigma$, preserves immersions and face embeddings, and so globalises to define a functor $C(\cdot)$ from the category of $\mathbb{Z}$-PL spaces to the category of cone complexes.
A.35Recession cone complex.
Let $\Sigma$ be a $\mathbb{Z}$-PL space, then we construct the recession cone complex $\mathrm{rec}(\Sigma)$ as follows. If $\sigma$ is a polyhedron, then
$|\mathrm{rec}(\sigma)|=\{v\in(N_{\sigma})_{\mathbb{R}}:\sigma+\mathbb{R}_{\geq 0}v\subset\sigma\},$
and the induced lattice is the restriction of $N_{\sigma}$. This construction is functorial and preserves face embeddings, hence glues to any $\mathbb{Z}$-PL space. The cone complex $\mathrm{rec}(\Sigma)$ is trivial if and only if $\Sigma$ is bounded, i.e. all its polyhedra are bounded, and $\mathrm{rec}(\Sigma)$ is strictly convex if and only if $\Sigma$ is.
Lemma A.36.
Let $\Sigma$ be a polyhedral complex. Then the recession cone complex $\mathrm{rec}(\Sigma)$ is a subcomplex of $C(\Sigma)$.
Proof.
It suffices to consider the case of a polyhedron $\sigma$. Let $v\in\mathrm{rec}(\sigma)$. Then for all $w\in\sigma$ and $\lambda\in\mathbb{R}_{>0}$ we have $w+\lambda v\in\sigma$ and so $\lambda^{-1}w+v\in C(\sigma)$. Take the limit as $\lambda\to\infty$ and it follows $v\in C(\sigma)$. ∎
A.37Stars and links.
Let $\Sigma$ be a polyhedral complex. Given a polyhedron $\tau$ of $\Sigma$ we write $\mathrm{star}(\tau)=\cup_{\sigma\geq\tau}\sigma$ for the star of $\tau$, and
$\mathrm{link}_{\Sigma}(\tau)=\overline{\mathrm{star}}_{\Sigma}(\tau)\setminus\mathrm{star}_{\Sigma}{\tau}$
for the link of $\tau$, which is again a PL space. For any polyhedron $\rho$ of $\mathrm{link}_{\Sigma}(\tau)$ we write $\tau+\rho$ for the unique minimal polyhedron of $\overline{\mathrm{star}}_{\Sigma}(\tau)$ containing $\tau$ and $\sigma$.
It is immediate to verify that $\mathrm{star}(\mathrm{rec}(\sigma))=\mathrm{rec}(\mathrm{star}(\sigma))$ and $\mathrm{link}(\mathrm{rec}(\sigma))=\mathrm{rec}(\mathrm{link}(\sigma))$.